### v1

# Junyu Cao and Mohsen Bayati

[…]
Junyu Cao and Mohsen Bayati
[…]
A Probabilistic Approach for Alignment with Human Comparisons
[…]
A Probabilistic Approach for Alignment with Human Comparisons
[…]
Junyu Cao
[…]
A growing trend involves integrating human knowledge into learning frameworks, leveraging subtle human feedback to refine AI models. Despite these advances, no comprehensive theoretical framework describing the specific conditions under which human comparisons improve the traditional supervised fine-tuning process has been developed. To bridge this gap, this paper studies the effective use of human comparisons to address limitations arising from noisy data and high-dimensional models. We propose a two-stage “Supervised Fine Tuning+Human Comparison” (SFT+HC) framework connecting machine learning with human feedback through a probabilistic bisection approach. The two-stage framework first learns low-dimensional representations from noisy-labeled data via an SFT procedure, and then uses human comparisons to improve the model alignment. To examine the efficacy of the alignment phase, we introduce a novel concept termed the “label-noise-to-comparison-accuracy” (LNCA) ratio. This paper theoretically identifies the conditions under which the “SFT+HC” framework outperforms pure SFT approach, leveraging this ratio to highlight the advantage of incorporating human evaluators in reducing sample complexity. We validate that the proposed conditions for the LNCA ratio are met in a case study conducted via an Amazon Mechanical Turk experiment.
[…]
model alignment, supervised fine-tuning, probabilistic bisection, human-AI interaction.
1 Introduction
In light of the increasing availability of data, the demand for data-driven algorithmic solutions has surged. Machine learning has emerged as the standard technique for numerous tasks, including natural language processing and computer vision, among others. Its great importance has also been demonstrated across various domains of operations research.
Traditional methodologies have predominantly relied on supervised fine-tuning (SFT) [^1] Supervised fine-tuning is usually used in Large Language Models (LLM). We use supervised fine-tuning and supervised learning interchangeably in this paper. techniques by training on noisy-labeled data. A common practice in this context for decision-makers is to directly adopt the output of the best predictor identified by SFT, where the best predictor can be defined by the minimizer of the empirical risk, for example. However, some issues may arise in this approach. One significant challenge emerges from the prevalence of large noise, which can substantially hinder model performance. On the other hand, to mitigate the model misspecification issue, the growing trend toward increasingly complex models may lead to over-parameterization. Over-parameterization presents additional obstacles to the design of efficient and accurate AI systems. In instances where the model exhibits high complexity, such as when the feature dimension is comparable to or exceeds the number of data points, multiple distinct estimators may exhibit seemingly equivalent empirical performance on the training dataset but perform poorly on unseen data, a scenario exacerbated by the inherent under-specification of models (D’Amour et al. 2022). Thus, decision-makers must consider strategic data utilization and refine models to achieve a better performance under different environments.
Recently, an ever-growing number of papers have incorporated human-knowledge into the learning framework (Christiano et al. 2017, Li et al. 2016). Fine-tuning with human feedback has shown a promising direction. For instance, in the design of chatbots, the performance can be adaptively and subsequently improved through the interactions with humans (Ziegler et al. 2019, Ouyang et al. 2022). Ouyang et al. (2022) focuses on fine-tuning approaches to aligning language models, where InstructGPT is fine-tuned with human feedback. In many cases, human evaluators possess domain expertise that goes beyond what can be encoded into a dataset or estimation method. Their deep understanding of the task domain enables them to make informed comparisons that consider context, relevance, and real-world applicability (Deng et al. 2020).
“Thinking is difficult, that’s why most people judge” – Carl Jung. Coming up with an answer is much more difficult than selecting a better answer from pair-wise comparisons. Zhou and Xu (2020) use pair-wise comparisons for story generation; Xu et al. (2023) collect human comparisons over two sampled LLM for the algorithm fine-tuning process. While prior research has showcased remarkable results when integrating human feedback across various applications, it has yet to establish a comprehensive theoretical framework detailing the specific conditions under which human comparison significantly augments the conventional SFT framework. To build on this research trajectory and fill this gap, our paper provides a theoretical framework for the strategic utilization of human comparisons. The primary focus is on addressing the inherent limitations posed by noisy data and the complexities of high-dimensional models. By actively and effectively leveraging human comparisons, we aim to navigate the intricacies of model alignment with the minimal labeling effort.
To connect the machine learning framework with human feedback, we propose a two-stage tractable framework through a probabilistic method. In the first stage, noisy labeled data is fed into the SFT oracle to obtain low-dimensional representations. More specifically, the algorithm can use lasso to select important features from high dimensions or use matrix completion to learn low-rank representations. The second stage employs human comparisons to improve model alignment within this low-dimensional space. Our framework, which utilizes the random utility model to describe human’s comparison behavior, can incorporate the possibility of false selection in human comparisons. Human may provide incorrect responses, though this noise tends to be substantially lower than that present in labeled data. To accommodate this aspect, we develop a probabilistic bisection approach for the alignment phase. We call this two-stage process as “SFT+HC” framework.
The goal of our paper is to refine the model with a certain precision level and a confidence level using the minimum possible data. The rationale behind this framework design is rooted in the observation that some representation learning techniques, such as autoencoders or deep learning architectures (Zhou and Paffenroth 2017, Akrami et al. 2019), are inherently robust to noisy labels (Li et al. 2021, Taghanaki et al. 2021). These methods aim to capture underlying patterns and structures in the data. Recently, Tsou et al. (2023) show that linear probes trained on the final embedding of pre-trained GPT-2 are surprisingly robust, offering competitive performance and, in some cases, even surpassing the results of full-model fine-tuning, where all weights are updated. In the second stage, we adopt human comparisons instead of human estimations because comparison is a much easier task. Our approach involves sequential design of comparisons and active data acquisition, with the objective of maximizing information and consequently minimizing sample complexity.
To describe the effectiveness of the refinement phase through human comparison, we introduce a new concept called “label-noise-to-comparison-accuracy” (LNCA) ratio. We theoretically characterize the condition for the “SFT+HC” framework performing better than the pure SFT algorithm based on this LNCA ratio. In the large-noise scenario, human evaluators with high selection accuracy can significantly reduce the sample complexity.
1.1 Our Contributions
Our contributions are summarized as follows:
  1. (I)
    Modeling: In a large-noise environment, machine learning models may exhibit fragility, due to the sensitivity to variations in input data and the dynamic nature of their surroundings. To address this challenge, we introduce a novel approach that leverages human comparisons to refine these models. To the best of our knowledge, our study represents the first theoretical attempt aimed at explaining the substantial enhancement of model performance under specific conditions through human comparisons. We present a novel two-stage framework based on a utility model for human comparisons. The model captures a natural phenomena that making distinctions between two options becomes increasingly challenging as they are more similar. Consequently, the accuracy of these comparisons may decrease as the model refines.
  2. (II)
    Theoretical Contributions:
    1. (a)
      Probabilistic bisection method: In the second phase, the decision-maker (DM) sequentially selects two options for human evaluators to compare. The fundamental question at this stage pertains to the efficient acquisition of valuable information from these human comparisons. To address this challenge, we introduce a probabilistic bisection method, which factors in the uncertainty arising from the accuracy of human comparisons. This method involves two types of actions: the ‘vertical move’ involves human evaluations of the same comparison. Once the algorithm identifies the superior choice with high probability, it proceeds with the ‘horizontal move,’ where the system selects the next pair for comparison. Our objective is to determine identify correct model parameters with a precision of $\epsilon$ and a confidence level of $1-\delta$. We establish the sample complexity required to achieve these precision and confidence levels.
    2. (b)
      Label-noise-to-comparison-accuracy ratio: We introduce a novel concept known as the ‘noise-to-human-knowledge ratio,’ which quantifies the relative scale between label noise and human-made comparison accuracy. In essence, when label noise is substantial, but humans can make selections with high accuracy, our two-stage framework outperforms pure Sequential Fine-Tuning (SFT). Specifically, when human selection accuracy deviates from 50%, the sample complexity for the refinement stage is $O(s\log(1/(\delta\varepsilon)))$. In cases where selection accuracy approaches 50% as the two options become indistinguishable from the true model, the sample complexity increases. We characterize the LNCA ratio conditions under which our SFT+HC framework outperforms pure SFT.
  3. (III)
    Numerical Performance: We evaluate the LNCA ratio through an empirical study on Amazon Mechanical Turk with a stylized task. The experiment result validates the condition under which “SFT+HC” performs better than pure SFT. Subsequently, we evaluated the performance of our proposed SFT+HC framework and compared it with the pure SFT approach across varying LNCA ratios. Our findings align with our theoretical results, revealing that the superiority of the SFT+HC framework becomes more pronounced under conditions of higher observational noise, higher accuracy of human comparisons, and lower dimensions of important features.
1.2 Literature Review
Alignment and Learning From Human Feedback. The alignment challenge seeks to synchronize human values with machine learning systems and direct learning algorithms towards the objectives and interests of humans. Machine learning models frequently demonstrate unforeseen deficiencies in their performance when they are implemented in practical contexts, which was identified as underspecification (D’Amour et al. 2022). Predictors that perform similarly in the training dataset can exhibit significant variations when deployed in various domains. The presence of ambiguity in a model might result in instability and suboptimal performance in real-world scenarios. Recent papers have focused on aligning models with human intentions, including training simple robots in simulated environments and Atari games (Christiano et al. 2017, Ibarz et al. 2018), fine-tuning language models to summarize text fine-tuning (Stiennon et al. 2020), optimize dialogue (Jaques et al. 2019). It has been demonstrated that using reinforcement learning from human feedback (RLHF) significantly boosts performance (Bai et al. 2022, Glaese et al. 2022, Stiennon et al. 2020, Dwaracherla et al. 2024). Ouyang et al. (2022) use reinforcement learning from human feedback (Stiennon et al. 2020) to fine-tune GPT-3 to follow a broad class of written instructions. In particular, Zhou and Xu (2020) use pair-wise comparisons for story generation. They find that when comparing two natural language generation (NLG) models, instead of asking human annotator to assign scores separately for samples generated by different models, which resembles the case in the ADEM model (Lowe et al. 2017), it is much easier for human annotators to directly compare one sample generated by the first model against another sample from the second model pairwisely and compute the win/loss rate. Zhu et al. (2023) provide reinforcement learning framework with human feedback, specifically, for function approximation for $K$-wise comparisons with policy learning as the target. Xu et al. (2023) propose a fine-tuning algorithm by inducing a complementary distribution over the two sampled LLM responses to model human comparison processes, in accordance with the Bradley-Terry model for pairwise comparisons. Different than those above-mentioned work, we provide a theoretical two-stage framework for integrating supervised learning with actively acquired human comparison labels, and characterize the conditions where the two-stage framework outperforms pure SFT algorithms.
Active Learning. The core concept behind active learning in machine learning theory is the ability of a model to interactively query a human or an oracle to obtain labels for new data points. Instead of passively learning from a fixed dataset, an active learning algorithm intelligently selects the most informative instances to query for labels. We refer readers to Settles (2009) for a comprehensive review of the active learning algorithm. The literature has concentrated on label complexity to demonstrate the benefit of active learning, which refers to the number of labels requested to attain a specific accuracy. For example, Cohn et al. (1994) demonstrates that active learning can have a substantially lower label complexity than supervised learning in the noiseless binary classification issue. In the presence of noise, other papers focus on the classification problems (Hanneke 2007, Balcan et al. 2006, Hanneke 2011) and regression problems (Sugiyama and Nakajima 2009, Beygelzimer et al. 2009), etc. Though our study has a similar goal of using the fewest possible samples to achieve the target accuracy, we concentrate on the human-AI collaboration framework and try to reduce the sample complexity of human comparisons.
Data-driven Decision Making and Human in the Loop. A large body of recent literature on operations research or operations management studied how to effectively utilize data to improve their decisions, including pricing, assortment, recommendation, etc. Motivated by the observation that human may sometimes have “private” information which an algorithm does not have access to that can improve performance, Balakrishnan et al. (2022) examined new theory that describes algorithm overriding behavior. They designed and experimentally tested feature transparency as an implementable approach that can help people better identify and use their private information. Ibrahim et al. (2021) studied how to address the shared information problem between humans and algorithms in a setting where the algorithm makes the final decision using the human’s prediction as input. Dietvorst et al. (2018, 2015) study algorithm aversion phenomenon where people often fail to use algorithms after learning that they are imperfect, though evidence-based algorithms consistently outperform human forecasters. Other papers study Human-AI collaboration on a large variety of topics (Bastani et al. 2021, Wu et al. 2022, Deng et al. 2020), including ride-hailing (Benjaafar et al. 2022) and healthcare (Reverberi et al. 2022), etc.
Probabilistic Bisection. The other related literature, which is independent of the above human-AI ML topics, is probabilistic bisection algorithm (Horstein 1963). Algorithms utilizing the PBA framework have been employed in several contexts, notably including the task of target localization (Tsiligkaridis et al. 2014), scanning electron microscopy (Sznitman et al. 2013), topology estimation (Castro and Nowak 2008), edge detection (Golubev and Levit 2003), and value function approximation for optimal stopping problems (Rodriguez and Ludkovski 2015). From the computational perspective, Frazier et al. (2019) has shown that probabilistic bisection converges almost as quickly as stochastic approximation. In this work, we design the model alignment algorithm by utilizing the probabilistic bisection approach, where the randomness is involved due to the inherent variability in the human judgement process.
1.3 Organization of the Paper
The remainder of this paper is organized as follows. Section 2 introduces the model setup. In Section 3, we introduce the model of human comparisons and characterize the sample complexity in a one-dimensional space. In Section 4, we present a two-stage framework which organically integrate SFT and human comparison. We specifically discuss the sample complexity when applying lasso in the first stage. Section 6 shows numerical experiments and managerial insights. Section 7 concludes the paper. Additional proofs and parameter calibrations are provided in the appendices.
2 Model
In this section, we introduce our model setup. We first introduce the supervised learning oracle that will be potentially used during the initial learning stage, and then present the fine-tuning procedure through human comparisons during the refinement stage.
2.1 Supervised Learning and Underspecifications
Given data $\{(X_{i},Y_{i})\}_{i\leq n}$ where $X_{i}\in\mathcal{X}$ and $Y_{i}\in\mathcal{Y}\subseteq\mathbb{R}$, the goal is to learn the true function that is parameterized by $\boldsymbol{\vartheta}^{*}\subseteq\mathbb{R}^{d}$, $h_{\boldsymbol{\vartheta}^{*}}\in\mathcal{H}:\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathcal{H}$ is the function class. Each data point is generated from a stochastic response:
$Y_{i}=h_{\boldsymbol{\vartheta}^{*}}(X_{i})+\epsilon_{i}\,,$
where $\epsilon_{i}$ is the noise with mean 0 and variance $\sigma^{2}$, where $\sigma$ could potentially be very large.
We specifically consider a high-dimensional problem that can be represented by low-dimensional manifolds; that is,
$h_{\boldsymbol{\vartheta}^{*}}(x)=f_{\boldsymbol{\theta}^{*},\varphi(\cdot)}(x)\,,$
where $f(\cdot)$ is assumed to be a generalized linear function in $\boldsymbol{\theta}^{*}$ and $\varphi(x)$. Here $\boldsymbol{\theta}^{*}\in\Theta\in\mathbb{R}^{s}$ is the unknown parameter [^2] We denote the parameter vector using the bold symbol., $\varphi(x)\in\mathbb{R}^{s}$ is the low-dimensional representation of feature $x$. It is assumed that $s\ll d$. For simplicity, we will remove the subscript $\varphi(\cdot)$ from $f$ when it is clear from the context.
We first provide the following two examples as an illustration.
Example 2.1 (Sparse Linear Models)
The parameter $\boldsymbol{\vartheta}^{*}$ has nonzero entries in the first $s$ dimensions and has zero entries in the rest of dimensions. In the linear model, $f_{\boldsymbol{\vartheta}^{*}}(x)=\sum_{i=1}^{s}\vartheta^{*}_{i}x_{i}$. In this case, $\boldsymbol{\theta}^{*}=(\vartheta^{*}_{1},\cdots,\vartheta^{*}_{s})$, $\varphi(x)=(x_{1},\cdots,x_{s})$, and $f(\cdot)$ is a linear function.
Example 2.2 (Generalized Low-rank Models)
In the generalized low-rank models, it is assumed that
$f_{\boldsymbol{\Theta}^{*}}(X)=\mu(\left\langle\boldsymbol{\Theta}^{*},xz^{\top}\right\rangle)=\mu(x^{\top}\boldsymbol{\Theta}^{*}z)$
where $\mu(\cdot)$ is the link function, $\boldsymbol{\Theta}^{*}\in\mathbb{R}^{d\times d}$ is a low-rank matrix and $X=xz^{\top}$, i.e., $\boldsymbol{\Theta}^{*}=\sum_{i=1}^{s}\sigma_{i}u_{i}v_{i}^{\top}$ where $u_{i}$ and $v_{i}\in\mathbb{R}^{d}$. Thus, $f_{\boldsymbol{\Theta}^{*}}(X)=\sum_{i=1}^{s}\sigma_{i}(x^{\top}u_{i})(v_{i}^{\top}z)$. In this case, $\varphi(X)$ is the projection of $X$ onto the low-rank space and $\boldsymbol{\theta}^{*}$ is the $s$-dimensional vector of singular values.
Suppose $h_{\boldsymbol{\vartheta}}$ can be estimated through a statistical loss function $\ell:\mathcal{Y}\times\mathbb{R}\rightarrow\mathbb{R}$, and the total loss is
$L_{n}(h_{\boldsymbol{\vartheta}})=\sum_{i=1}^{n}\ell(Y_{i},h_{\boldsymbol{\vartheta}}(X_{i})).$
For example, $\ell$ can represent the squared loss, the regularized squared loss, among others. To deal with the high-dimensional issue, the algorithm learns the low-dimensional representation $\varphi(x)$ as well as the parameter $\boldsymbol{\theta}^{*}$. To learn the sparse linear models for example, one can apply lasso to select important features and learn the model parameters. To learn the low-rank models, one can use trace regression to learn the low-dimensional representations.
Define $h_{\hat{\boldsymbol{\vartheta}}_{n}}$ as the function contained in $\mathcal{H}$ that minimizes empirical risk $L_{n}$. In the existing literature, it is very common that the decision-maker directly employs the predictor $h_{\hat{\boldsymbol{\vartheta}}_{n}}$. However, some issues may arise. In high-dimensional problems where significant noise is present, approaches striving to balance the variance-bias trade-off may occasionally diverge from intuitive expectations. We use the following example to illustrate this potential issue.
Example 2.3
Consider a linear pricing model with high-dimensional contexts, i.e., $Y_{i}=\vartheta_{1}\cdot p_{i}+\vartheta_{2}^{\top}X_{i}+\epsilon_{i}$ where $Y_{i}$ is the demand, $p_{i}$ is the price, and the vector $X_{i}$ represents features of product $i$, and also assume that $\vartheta_{1}<0$. However, in a high-noise regime we may incorrectly estimate a positive value for $\vartheta_{1}$. Please see Appendix 9.1 for numerical illustrations.
In Example 2.3, the positive correlation between price and demand is misinterpreted due to model underspecification. The requirement of $\vartheta_{1}<0$ should be specified based on human knowledge. In the face of the challenge posed by underspecification within a potentially very noisy environment, we take a critical step towards refinement, which we term “fine-tuning through human comparisons.” Human can do comparisons with a much higher precision. For example, based on the human knowledge, it can be immediately corrected that the demand and the price is negatively correlated. This particular stage comes into play after the output generated by our machine learning pipelines, represented as $\hat{\varphi}(\cdot)$ and $\hat{\boldsymbol{\theta}}_{0}$ (which denotes the estimator of $\varphi$ and $\boldsymbol{\theta}$), has been produced. During this stage, we focus on enhancing the model within a lower-dimensional space. We move on to the refinement step after the first stage where the kernel function, which provides a low-dimensional representation, has been effectively trained on the noisy-labeled data.
2.2 Utility Model of Human Comparisons
In this section, we introduce the human comparison model. Given two potential models $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$, the system asks human to compare and select the better prediction model. If $f_{\boldsymbol{\theta}_{1}}$ is more preferable than $f_{\boldsymbol{\theta}_{2}}$, we denote it as $f_{\boldsymbol{\theta}_{1}}\succeq f_{\boldsymbol{\theta}_{2}}$.
The preference over the model $f_{\boldsymbol{\theta}}$ is captured by a utility function $u(\boldsymbol{\theta})$, where $u(\boldsymbol{\theta})$ depicts the distance between the model parameterized by $\boldsymbol{\theta}$ and the true model. That is, we assume $u(\boldsymbol{\theta})=-d(\boldsymbol{\theta},\boldsymbol{\theta}^{*})$ where $d(\cdot,\cdot)$ is a distance function. The larger value of utility indicates a better model. The true model parameter $\boldsymbol{\theta}^{*}$ achieves the highest utility. For example, it can be the negative two-norm distance, i.e., $u(\boldsymbol{\theta})=-\|\boldsymbol{\theta}-\boldsymbol{\theta}^{*}\|_{2}$. We formally state the assumption about the utility function as follows:
{assumption}
[Utility function] Assume $u(\boldsymbol{\theta})=-d(\boldsymbol{\theta},\boldsymbol{\theta}^{*})$, where $d(\cdot,\cdot)$ is a distance function. The utility function is unimodal in each coordinate. That is, for each coordinate $\theta_{i}$, when fixing all other coordinates, $u(\boldsymbol{\theta})$ is monotonically increasing when $\theta_{i}<\theta_{i}^{*}$ and monotonically decreasing when $\theta_{i}>\theta_{i}^{*}$. Moreover, the utility gets lower when $\boldsymbol{\theta}$ gets far away from the true parameter. The utility function can be either bounded or unbounded (e.g., the Euclidean distance). We give another example of the bounded utility function as an illustration.
Example 2.4
[…]
The comparison between $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$ provides information on which side is closer to the true parameter $\boldsymbol{\theta}^{*}$. Define the hyperplane $H(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})=\{\boldsymbol{\theta}:d(\boldsymbol{\theta},\boldsymbol{\theta}_{1})=d(\boldsymbol{\theta},\boldsymbol{\theta}_{2})\}$, and half-space $H^{-}(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})=\{\boldsymbol{\theta}:d(\boldsymbol{\theta},\boldsymbol{\theta}_{1})<d(\boldsymbol{\theta},\boldsymbol{\theta}_{2})\}$ and $H^{+}(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})=\{\boldsymbol{\theta}:d(\boldsymbol{\theta},\boldsymbol{\theta}_{1})>d(\boldsymbol{\theta},\boldsymbol{\theta}_{2})\}$. The condition $u(\boldsymbol{\theta}_{1})\geq u(\boldsymbol{\theta}_{2})$ implies that $\boldsymbol{\theta}^{*}\in H^{-}(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})$. Assumption 2.2 implies that $H(\theta_{1},\theta_{2})$ contains a unique point when $\theta_{1}$ and $\theta_{2}$ are in a one-dimensional space.
Remark 2.5
In practice, the comparison between two models can be conducted in various ways. For example, it can be evaluated on a single point with high confidence (either from human’s past experience or their observation), denoted as $\{x,y\}$. When comparing two models, the one with smaller square loss is more preferred, i.e.,
$\mathbb{P}(\boldsymbol{\theta}_{1}\succeq\boldsymbol{\theta}_{2})$ $=\mathbb{P}\left((y-f_{\boldsymbol{\theta}_{1}}(x))^{2}\leq(y-f_{\boldsymbol{\theta}_{2}}(x))^{2}\right).$
We will elaborate this process in Section 5.
The overall goal is to find $\boldsymbol{\theta}$ that achieves near the highest utility, or equivalently, gets close to the true model parameter $\boldsymbol{\theta}^{*}$ within a certain precision. Suppose that human’s selection process follows a random utility model (RUM). In the process of querying, human’s opinion regarding $f_{\boldsymbol{\theta}}$ is
$U(\boldsymbol{\theta})=u(\boldsymbol{\theta})+\beta_{h}+\xi,$
where $\beta_{h}$ represents the human bias, and $\xi$ is the unobserved noise. The human bias $\beta_{h}$ can vary between individuals. We assume the noise $\xi$ follows Gumbel distribution with parameter $\gamma$, which represents the expert level of the human. We say the human makes the correct choice if he selects $\boldsymbol{\theta}_{1}$ ($\boldsymbol{\theta}_{2}$) when $u(\boldsymbol{\theta}_{1})$ ($u(\boldsymbol{\theta}_{2})$) is higher than $u(\boldsymbol{\theta}_{2})$ ($u(\boldsymbol{\theta}_{1})$). A smaller value of $\gamma$ corresponds to a higher certainty of the selection correctness. When $\gamma$ approaches 0, human can always select the better one with almost 100% accuracy.
Lemma 2.6 (Precision)
For any two models $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$, the probability that human will make the right selection is
$\frac{1}{1+\exp(-|u(\boldsymbol{\theta}_{2})-u(\boldsymbol{\theta}_{1})|/\gamma)}\,,$
which is strictly larger than 1/2.
Lemma 2.6 implies that under RUM, the selection accuracy is always higher than 50%, even with the nonzero human bias. Intuitively, the differentiation would become harder when $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$ both get closer to the true parameter. However, how fast it converges to 1/2 when $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$ approach $\boldsymbol{\theta}^{*}$ depend on the distance function. Under this probabilistic selection model, we will design the algorithm and investigate the sample complexity. The parameter $\boldsymbol{\theta}$ may lie in a multi-dimensional space. To illustrate our algorithm, we first describe the one-dimensional human comparison procedure, and then extend to the multi-dimensional scenario.
3 One-dimensional Human Comparison
Once the low-dimensional representation $\hat{\varphi}$ has been established in the initial stage, as detailed in Section 4, we introduce a framework for model alignment through human comparisons. The parameter $\boldsymbol{\theta}$ may lie in a multi-dimensional space. To illustrate our algorithm, we first describe the one-dimensional human comparison procedure, and then extend to the multi-dimensional scenario.
The primary challenge in designing the framework lies in strategically selecting pairwise comparisons to minimize sample complexity, which bears resemblance to the principles of active learning. In this section, we first discuss the deterministic selection (with 100% accuracy) and then study the probabilistic selection. For the probabilistic selection, we propose an algorithm named Alignment through Probabilistic Bisection (APB) and characterize its sample complexity.
3.1 Deterministic Bisection
We first illustrate the algorithm in a one-dimensional space with 100% selection accuracy. This scenario corresponds to $\gamma\to 0$ in RUM. In the one-dimensional case, suppose $\theta^{*}\in\Theta\subseteq\mathbb{R}$ and $|\theta|\leq\beta_{\Theta}$ for all $\theta\in\Theta$.
The overall goal is to use the least samples to find some $\theta$ such that $d(\theta,\theta^{*})\leq\varepsilon$. When human is able to select a better model with 100% accuracy, the deterministic bisection algorithm can be utilized to approach the true parameter. The algorithm operates by maintaining a search interval interval $[\theta^{-},\theta^{+}]$ that contains the true parameter $\theta^{*}$. The interval is initialized by $[-\beta_{\Theta},\beta_{\Theta}]$, and at round $k$ the center of this interval with respect to the distance $d(\cdot,\cdot)$ is used as query point. Specifically, defining the query point $\theta_{k}$ to be the single point in the set $H(\theta^{-},\theta^{+})$, we ask the human to compare utility of two points, denoted by $c_{\Delta}^{-}(\theta_{k})$ and $c_{\Delta}^{+}(\theta_{k})$, that have equal distance $\Delta/2$ from $\theta_{k}$ in opposite directions, where $\Delta$ is called comparison granularity and is a tuning parameter. For example, for the Euclidean distance, at each round $\theta_{k}=(\theta^{-}+\theta^{+})/2$, and $c_{\Delta}^{-}(\theta_{k})=\theta_{k}-\Delta/2$ and $c_{\Delta}^{+}(\theta_{k})=\theta_{k}+\Delta/2$. If the human selects $c_{\Delta}^{-}(\theta_{k})$ then we eliminate the upper half of the search interval by updating $\theta^{+}$ to be equal to $\theta_{k}$; otherwise, we eliminate the bottom half of the search interval by setting $\theta^{-}$ to be equal to $\theta_{k}$. After this bisection step, we continue the process by selecting $\theta_{k+1}$ to be $H(\theta^{-},\theta^{+})$ again, and repeat the above till the length of the search interval is below the precision $\varepsilon$. The pseudo code for this is presented in Algorithm 1.
$\theta_{k}$$\theta^{*}$$\theta^{-}$$\theta^{+}$$c_{\Delta}^{+}(\theta_{k})$$c_{\Delta}^{-}(\theta_{k})$keepeliminate
Figure 1: An illustrative example. Among two choices $c_{\Delta}^{-}(\theta_{k})$ and $c_{\Delta}^{+}(\theta_{k})$, $c_{\Delta}^{-}(\theta_{k})$ would be selected since it is closer to the true parameter. In this case, the interval $(\theta_{k},\theta^{+}]$ is eliminated.
Algorithm 1 DB (Deterministic Bisection)
1:input: precision $\varepsilon$; comparison granularity $\Delta$; $k=0$; $\theta_{k}=0$; $\theta^{+}=\beta_{\Theta}$; $\theta^{-}=-\beta_{\Theta}$;
2:repeat
3:     Ask human to evaluate $c_{\Delta}^{-}(\theta_{k})$ and $c_{\Delta}^{+}(\theta_{k})$
4:     if $c_{\Delta}^{+}(\theta_{k})$ is chosen then
5:         $\theta^{-}=\theta_{k}$;
6:     else $\theta^{+}=\theta_{k}$;
7:     end if
8:     $k=k+1$; $\theta_{k}=H(\theta^{-},\theta^{+})$;
9:until $d(\theta^{+},\theta^{-})\leq\varepsilon$.
Proposition 3.1
Suppose human makes the selection with 100% accuracy $(\gamma\to 0)$. Fix $\varepsilon>0$. After $O(\log_{2}\left(\frac{\beta_{\Theta}}{\varepsilon}\right))$ rounds, Algorithm 1 reaches a point $\theta_{k}$ such that $d(\theta_{k},\theta^{*})\leq\varepsilon$.
However, the deterministic model of human selection described above exhibits two significant limitations: firstly, in practical scenarios, humans may not always be able to identify the correct choice, implying that $\gamma>0$; secondly, the accuracy of human selections is influenced by the comparison granularity $\Delta$. To address these concerns, we introduce a more general probabilistic model in the following subsection.
3.2 Probabilistic Bisection
First, we enrich the model to capture the probabilistic nature of human comparisons by noting that, in the RUM, when $\gamma$ is strictly positive, human’s selection is random with certain distribution that favors the correct answer, parameterized by $\gamma$. Precisely, Lemma 2.6 indicates that human will select $c_{\Delta}^{-}(\theta)$ with probability
$p^{-}_{\gamma}(\theta):=\left[1+\exp\left\{\frac{u(c_{\Delta}^{+}(\theta))-u(c_{\Delta}^{-}(\theta))}{\gamma}\right\}\right]^{-1}\,,$
and selects $c_{\Delta}^{+}(\theta)$ with probability $p^{+}_{\gamma}(\theta)$ equal to $1-p^{-}_{\gamma}(\theta)$.
We note that, in the above equation, the selection probability depends on the comparison granularity. This dependence, naturally captures difficulty of the selection for the human. Specifically, when $\Delta$ is too small, $c_{\Delta}^{-}(\theta)$ and $c_{\Delta}^{+}(\theta)$ converge to each other, hence $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$ converge to $1/2$, representing the difficulty of distinguishing between too highly similar choices. On the other hand, when $\Delta$ is too large, one expects the selection to become difficult as well as both options are far away from $\theta^{*}$. This can be captured for example when the distance function converges to a finite value for far away points from $\theta^{*}$ as in Example 2.4. In this case, $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$ could also converge to 1/2 (see Example 3.2). In the remaining of the paper, for brevity, we do not include $\Delta$ in the notations $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$.
Example 3.2
Consider $u(\theta)=-\frac{1}{1+\exp(-|\theta|)}$ with the highest utility achieved at $\theta=0$. When $\Delta$ is too large, both $u(c_{\Delta}^{+}(\theta))$ and $u(c_{\Delta}^{-}(\theta))$ are close to -1 even $\theta$ is far away from 0, and thus $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$ are both close to 1/2.
Definition 3.3 ($(\varepsilon,\delta)$-alignment problem)
Let $\varepsilon>0$ denote the precision level and $\delta>0$ represent the confidence level. The $(\varepsilon,\delta)$-alignment problem aims to identify a $\theta$ such that $\mathbb{P}(d(\theta,\theta^{*})\leq\varepsilon)\geq 1-\delta$.
Our objective is to solve the $(\varepsilon,\delta)$-alignment problem with small sample complexity. The algorithm design is inspired by the probabilistic bisection developed in Frazier et al. (2019). While their work primarily focuses on the analysis of convergence rates, our emphasis is on conducting a sample complexity analysis for an algorithm designed to terminate when both the precision level and confidence level criteria are met.
If $\theta<\theta^{*}$, then $p_{\gamma}^{-}(\theta)<1/2$ and $p_{\gamma}^{+}(\theta)>1/2$; vice versa. We introduce some notation for simplicity. Let $Y(\theta)$ denote the selection: $Y(\theta)=c_{\Delta}^{-}(\theta)$ with probability $p_{\gamma}^{-}(\theta)$ and $Y(\theta)=c_{\Delta}^{+}(\theta)$ with probability $p_{\gamma}^{+}(\theta)$. Define $\tilde{Z}(\theta)=2\cdot\boldsymbol{1}(Y(\theta)=c_{\Delta}^{+}(\theta))-1$, so that $\tilde{Z}(\theta)=1$ if $Y(\theta)=c_{\Delta}^{+}(\theta)$ (so that we believe $\theta^{*}$ more likely to be to the right of $\theta$) and $\tilde{Z}(\theta)=-1$ if $Y(\theta)=c_{\Delta}^{-}(\theta)$ (so that we believe $\theta^{*}$ more likely to be to the left of $\theta$). Note that we have $\tilde{Z}(\theta)=1$ with probability $p_{\gamma}^{+}(\theta)$. Also, define
$\tilde{p}(\theta)=\begin{cases}\mathbb{P}(\tilde{Z}(\theta)=1),&\mbox{if }\theta\leq\theta^{*}\\ \mathbb{P}(\tilde{Z}(\theta)=-1),&\mbox{if }\theta>\theta^{*}\end{cases}\,,$
so that $\tilde{p}(\theta)$ is the probability that $\tilde{Z}(\theta)$ correctly indicates the direction of the root $\theta^{*}$ from $\theta$.
Figure 2 shows the structure of introducing our main algorithm. The general idea of the Algorithm APB is described as follows. There are two types of moves: the horizontal moves and the vertical moves. For any fixed $\theta$, we ask human to repeatedly evaluate $c_{\Delta}^{-}(\theta)$ and $c_{\Delta}^{+}(\theta)$. We refer this as the vertical moves. Once the vertical stopping criteria is satisfied, the algorithm decides the next $\theta$ to evaluate according to the updated distribution. We call this step as the horizontal moves. The horizontal moves continue until the horizontal stopping criteria is satisfied. In the following two sections, we discuss horizontal moves and vertical steps, as well as the stopping criteria, in sequence. Then we introduce the two-stage Human-AI collaboration framework for multiple dimensions in Section 4 and practical implementations in Section 5.
One-dimensional APB (Section 3.2) Horizontal moves (Section 3.2.2); Complexity characterization (Section 3.3) Sample complexity (Theorem 3.5) Two-stage framework for Human-AI collaboration (Section 4.1); complexity characterization (Section 4.3) Practical implementations (Section 5) Vertical moves (Section 3.2.1); Complexity characterization (Section 3.4)
Figure 2: Roadmap for introducing 2.
3.2.1 Vertical moves
Fix some $\theta$. Repeatedly, we ask human evaluations of the query point $\theta$, achieved through comparisons of $c^{+}_{\Delta}(\theta)$ and $c^{-}_{\Delta}(\theta)$. We observe a sequence of signs $\{\tilde{Z}_{i}(\theta)\}_{i}$. If $\theta<\theta^{*}$, then the expectation $\mathbb{E}[\tilde{Z}_{i}(\theta)]=2\tilde{p}(\theta)-1>0$; if $\theta>\theta^{*}$, then the expectation $\mathbb{E}[\tilde{Z}_{i}(\theta)]=1-2\tilde{p}(\theta)<0$. From the response $\tilde{Z}_{i}(\theta)$, we define the simple random walk $S(\theta)=\{S_{m}(\theta):m\geq 0\}$ with $S_{0}(\theta)=0$ and $S_{m}(\theta)=\sum_{i=1}^{m}\tilde{Z}_{i}(\theta)$ for $m\geq 1$. Sequential test of power one indicates whether the drift $\omega$ of a random walk satisfies the hypothesis $\omega<0$ versus $\omega>0$.
Define the stopping time $\tau^{\uparrow}_{1}(\theta)=\inf\{m\in\mathbb{N}:|S_{m}(\theta)|\geq\hbar_{m}\}$ where $\{\hbar_{m}\}_{m}$ is a positive sequence. Here $\uparrow$” ($\rightarrow$) represents “vertical” (horizontal, respectively). The test decides that $\theta<\theta^{*}$ if $S_{\tau_{1}^{\uparrow}(\theta)}(\theta)\geq\hbar_{\tau^{\uparrow}_{1}(\theta)}$, that $\theta>\theta^{*}$ if $S_{\tau_{1}^{\uparrow}(\theta)}(\theta)\leq-\hbar_{\tau^{\uparrow}_{1}(\theta)}$ and the test does not make a decision if $\tau^{\uparrow}_{1}(\theta)=\infty$.
Assume now that the algorithm queries at some point $\theta_{k}\neq\theta^{*}$ at the $(k+1)^{th}$ iteration. We then observe a random walk with $m^{th}$ term $S_{k,m}=S_{k,m}(\theta_{k})=\sum_{i=1}^{m}Z_{k,i}(\theta_{k})$ until we reach the end of the power-one test. Define the new signal
$Z_{k}(\theta_{k})=\begin{cases}+1,&\mbox{if }S_{k,\tau_{1}^{\uparrow}(\theta_{k})}>0\,,\\ -1,&\mbox{if }S_{k,\tau_{1}^{\uparrow}(\theta_{k})}<0\,.\end{cases}$
Lemma 3.4
Define $\hbar_{m}=(2m(\ln(m+1)-\ln\eta))^{1/2}$. On the event $\theta_{k}>\theta^{*}$, $\mathbb{P}(Z_{k}(\theta_{k})=+1|\theta_{k})\leq\eta/2$. On the event $\theta_{k}<\theta^{*}$, $\mathbb{P}(Z_{k}(\theta_{k})=-1|\theta_{k})\leq\eta/2$. That is, the detection accuracy $p(\theta)$ when the vertical move stops is bounded by $p(\theta)\geq 1-\eta/2$.
Lemma 3.4 states that the detection accuracy is lower bounded by $p_{c}:=1-\eta/2$ where the accuracy can be chosen through the selection of $\eta$. As $\theta$ approaches $\theta^{*}$, the differentiation between choices becomes increasingly challenging because the drift tends to diminish, potentially requiring more steps in the testing process. However, it is crucial to remember that our objective is to locate a $\theta$ such that $|\theta-\theta^{*}|\leq\varepsilon$ with a confidence level of $1-\delta$. In the vicinity of $\theta^{*}$, waiting until $|S_{m}(\theta)|\geq\hbar_{m}$ to terminate the test may not be necessary. A high volume of queries at a particular point $\theta_{k}$ suggests close proximity between $\theta_{k}$ and $\theta$, making it advisable to consider terminating the test sooner.
Based on this insight, we introduce the stopping time $\tau^{\uparrow}(\delta,\varepsilon)$ that is determined by the precision and confidence level. This stopping time regulates the maximum number of queries carried out at a single point as $\theta$ approaches the local neighborhood of the true parameter (a concept to be elaborated on in Section 3.4). Specifically, as $\theta$ converges towards $\theta^{*}$ within a defined local neighborhood, denoted as $\{\theta^{\prime}:d(\theta,\theta^{\prime})\leq a\}$ for some positive constant $a$, surpassing the threshold $\tau^{\uparrow}(\delta,\varepsilon)$ signifies the difficulty in distinguishing between the parameters due to their proximity. The algorithm terminates.
3.2.2 Horizontal moves
Algorithm 2 begins with a prior density $\rho_{0}$ on $[-\beta_{\Theta},\beta_{\Theta}]$ that is positive everywhere. For example, $\rho_{0}$ can be chosen as a uniform distribution. Choose some $p\in(1/2,p_{c})$ and let $q=1-p$. When the algorithm queries at the point $\theta_{k}$ to obtain $Z_{k}$, we update the density $\rho_{k}$:
  1. (I)
    If $Z_{k}(\theta_{k})=+1$, then
    $\rho_{k+1}(\theta^{\prime})=\begin{cases}2p\rho_{k}(\theta^{\prime}),&\mbox{if}\theta^{\prime}\geq\theta_{k};\\ 2q\rho_{k}(\theta^{\prime}),&\mbox{if }\theta^{\prime}<\theta_{k}.\end{cases}$
  2. (II)
    If $Z_{k}(\theta_{k})=-1$, then
    $\rho_{k+1}(\theta^{\prime})=\begin{cases}2q\rho_{k}(\theta^{\prime}),&\mbox{if}\theta^{\prime}\geq\theta_{k};\\ 2p\rho_{k}(\theta^{\prime}),&\mbox{if }\theta^{\prime}<\theta_{k}.\end{cases}$
The next query point $\theta_{k+1}$ is determined as the median of $\rho_{k+1}$. The probability density in the correct side would be increased while the probability density in the wrong side would be reduced. Define $\tau^{\rightarrow}(\delta,a)$ as the number of horizontal moves to reach $a$-neighborhood of $\theta^{*}$ with confidence level $1-\delta$ (which will be specified later). That is, parameter $\theta$ moves to the local $a$-neighborhood of $\theta^{*}$ with probability at least $21-\delta/$ 2 when the number of steps exceeds $\tau^{\rightarrow}(\delta/2,a)$. The horizontal move stops when the total number of movement reaches $\tau^{\rightarrow}(\delta/2,\varepsilon)$ since the precision level $\varepsilon$ is achieved.
In summary, there are two scenarios:
  1. (I)
    If the horizontal move is less than $\tau^{\rightarrow}(\delta/2,a)$ (indicating it may not yet reach the local neighborhood of $\theta^{*}$ within a distance of $a$), the vertical test stops when it reaches $\tau_{1}^{\uparrow}(\theta)$. Then it moves to the next query point $\theta$;
  2. (II)
    If the horizontal move is more than $\tau^{\rightarrow}(\delta/2,a)$ (signifying entry into the local neighborhood of $\theta^{*}$ within a distance of $a$):
    1. (a)
      If the vertical move reaches $\tau^{\uparrow}(\delta/2,\varepsilon)$ first, the algorithm terminates, having identified a $\theta$ within the $\varepsilon$ range of $\theta^{*}$;
    2. (b)
      Alternatively, if the vertical movement reaches $\tau_{1}^{\uparrow}(\theta)$ first, the vertical test stops and then transitions to the next query point $\theta$.
The rationale behind distinguishing between local and non-local scenarios lies in the observation that two distinct situations can result in a significant number of vertical moves: when both compared models are performing well (i.e., $\theta_{1}$ and $\theta_{2}$ are close to $\theta^{*}$), or when both models are performing poorly (i.e., $\theta_{1}$ and $\theta_{2}$ are far from $\theta^{*}$). In the latter scenario, it is imperative that the algorithm does not terminate prematurely, as $\theta$ has not yet approached $\theta^{*}$ sufficiently. Therefore, the stopping criterion $\tau^{\uparrow}(\delta,\varepsilon)$ is only applied when $\theta$ transitions into the local neighborhood of $\theta^{*}$.
Algorithm 2 is described as follows. We initialize the distribution as $\mu_{0}$. Without any historical data, $\mu_{0}$ can be initialized as the uniform distribution over the confidence region. Alternatively, $\mu_{0}$ can be approximated by the empirical distribution of the estimator from the bootstrap method applying to the first phase. At step $k$, the algorithm tests $\theta_{k}$ by collecting $Z_{k,s}(\theta_{k})$. If $\theta$ is located at the local neighborhood of $\theta^{*}$ ($k\geq\tau^{\rightarrow}(\delta/2,a)$) and the current step label $s$ is larger than $\tau^{\uparrow}(\delta/2,\varepsilon)$, then the algorithm stops and $\theta=\theta_{k}$. Otherwise, the vertical move stops when $s$ is larger than $\tau_{1}^{\uparrow}(\theta_{k})$. Once the test on $\theta_{k}$ stops, we update prior distribution $\rho_{k}$ and select the next $\theta_{k+1}$ as the median of the updated distribution.
Algorithm 2 APB (Alignment through Probabilistic Bisection)
1:input: confidence level $\delta$, precision $\varepsilon$; initial distribution $\mu_{0}$; local parameter $a>0$; $k=0$; $\eta$;
2:repeat
3:     pick $\theta_{k}$ as the median of $\mu_{k}$
4:     $s=0$
5:     repeat
6:         test $\theta_{k}$ by collecting $Z_{k,s}(\theta_{k})$
7:         $s=s+1$
8:         if $k\geq\tau^{\rightarrow}(\delta/2,a)$ and step $s$ satisfies stopping criteria $\tau^{\uparrow}(\delta/2,\varepsilon)$ then
9:              Algorithm terminates and $\theta=\theta_{k}$
10:         end if
11:     until vertical step $s$ satisfies stopping criteria $\tau_{1}^{\uparrow}(\theta_{k})$
12:     $k=k+1$; update prior distribution $\rho_{k}$
13:until horizontal step $k$ satisfies stopping criteria $\tau^{\rightarrow}(\delta,\varepsilon)$
3.3 Complexity of Horizontal Moves
Let $T^{\uparrow}(\theta)$ denote the number of comparisons that are collected at $\theta$ and $T^{\rightarrow}$ denote the number of horizontal moves; let $T_{k}^{+}$ denote the total number of comparisons that are collected after $k$ horizontal-moves, where “$+$” denotes both horizontal- and vertical-moves. Note that $T^{\uparrow}(\theta)$, $T^{\rightarrow}$, and $T_{k}^{+}$ are all random variables. In what follows, we discuss the complexity of vertical and horizontal moves, and characterize its sample complexity.
We first characterize the complexity of horizontal moves. Fix some $a\in(0,\theta^{*}+\beta_{\Theta})$. Define three regions: $A=[-\beta_{\Theta},\theta^{*}-a]$, $B=(\theta^{*}-a,\theta^{*}]$, $C=(\theta^{*},\beta_{\Theta}]$. Define $T(a)=\inf\{k^{\prime}\geq 0:\theta_{k}\in B\cup C,\forall k\geq k^{\prime}\}$ to be the time required until the sequence of medians never reenters $A$. Let $\nu_{k}$ denote the (random) measure corresponding to the conditional posterior distribution, conditional on lying to the left of $\theta^{*}$, so that for any measurable $D$, $\nu_{k}(D)=\mu_{k}(D\cap[-\beta_{\Theta},\theta^{*}])/\mu_{k}([-\beta_{\Theta},\theta^{*}])$. Next, fix $\Delta\in(0,1/2)$ and define the stopping time $\tau(a)$ to be the first time that the conditional mass in $B$ lies above $1-\Delta$ and the median $\theta_{k}$ lies to the left of $\theta^{*}$, i.e.,
$\tau(a)=\inf\{k\geq 0:\nu_{k}(B)>1-\Delta,\theta_{k}<\theta^{*}\}.$
In line with the approach taken by Frazier et al. (2019) in establishing a connection between $T(a)$ and $\tau(a)$, we present a similar result in Lemma 3.5.
Lemma 3.5
It holds that
$T(a)\leq_{s.t.}\tau(a)+R,$ (3.1)
where $R$ is a random variable that is independent of $\tau(a)$ and $a$.
From (3.1), to bound $T(\cdot)$, we characterize the distribution of $\tau(\cdot)$. Algorithm 2 starts with an initial distribution $\mu_{0}$. Without any prior information, $\mu_{0}$ can be set as a uniform distribution; otherwise we can use the prior information to update $\mu_{0}$. We first make the following mild assumption regarding the initial prior distribution.
{assumption}
[Prior distribution] The initial prior distribution satisfies that $\mu_{0}(B)\geq a^{\varsigma}$ for some $\varsigma\leq 1$.
The uniform distribution satisfies Assumption 3.5 since $\mu_{0}(B)=\mu_{0}((\theta^{*}-a,\theta^{*}])=a$. With additional information, the prior distribution would be more centered around the true parameter $\theta^{*}$. Thus, $\varsigma$ is assumed to be less than or equal to 1. Under this assumption of prior distribution, we analyze the distribution of $\tau(\cdot)$ in Lemma 3.6.
Lemma 3.6
Under Assumption 3.5, by setting $r=r_{2}\alpha/(4\varsigma)$,
$\mathbb{P}(\tau(\omega e^{-rk})>k/2)\leq e^{-r_{2}\alpha k/4}\cdot\frac{\theta^{*}}{\omega^{\varsigma}}+\beta e^{-r_{1}k/2},$
for some $r_{1},r_{2},\alpha>0$ and any $\omega>0$.
Lemma 3.6 has shown the light tail distribution of $\tau(\cdot)$. For small $\omega$, note that $\omega^{\varsigma}$ decreases in $\varsigma$. That is, when the initial prior distribution is more centered around the true parameter, $\varsigma$ is smaller and thus the right-hand-side tail distribution is lighter.
We next characterize the complexity of horizontal moves.
Theorem 3.7 (Sample complexity of horizontal moves)
Under Assumption 3.5, by setting
$\tau^{\rightarrow}(\delta,\varepsilon)=\max\left\{\frac{4\varsigma(1+\varsigma)\log(\frac{8\theta^{*}}{\delta\varepsilon^{\varsigma}})}{r_{2}\alpha},\frac{2\log(8\beta_{2}/\delta)}{r_{1}},\frac{\log(8\beta_{2}/\delta)}{r_{R}}\right\}.$
we have that
$\mathbb{P}(|\theta_{k}-\theta^{*}|\leq\varepsilon)>1-\delta,\quad\forall k\geq\tau^{\rightarrow}(\delta,\varepsilon)\,.$
Theorem 3.7 states that after $\tau^{\rightarrow}(\delta,\varepsilon)$ horizontal moves, $\theta_{k}$ reaches $\varepsilon$-neighborhood of the true parameter with probability at least $1-\delta$. The complexity of this process depends on $\log(1/\delta)$, $\log(1/\varepsilon)$, and $\varsigma(1+\varsigma)$. When the initial prior distribution is more centered around the true parameter $\theta^{*}$, fewer horizontal moves are needed.
Define the local region of $\theta^{*}$ as $[\theta^{*}-a,\theta^{*}+a]$. We divide the horizontal moves to two phases: Phase I includes all steps until the sequence of medians never reenters $[-\beta_{\Theta},\theta^{*}-a]\cup[\theta^{*}+a,\beta_{\Theta}]$; Phase II includes all steps after Phase I. Similar to the definition of $T(a)$, we define $T^{\prime}(a)$ as the stopping time that the medians never reenters $A^{\prime}=[\theta^{*}+a,\beta_{\Theta}]$. Let $\psi(a)$ be the time required until the sequence of medians never reenters $A\cup A^{\prime}$, i.e., $\psi(a)=\max\{T(a),T^{\prime}(a)\}$. Proposition 3.8 provides an upper bound for the expectation of moves outside of $[\theta^{*}-a,\theta^{*}+a]$.
Proposition 3.8
The expectation of $\psi(a)$ is bounded by
$\mathbb{E}[\psi(a)]\leq\frac{4\theta^{*}}{r_{2}\alpha a}+\frac{4\beta_{1}}{r_{1}}+\frac{2\beta_{2}}{r_{R}}.$
3.4 Complexity of Vertical Moves
The vertical move has two stopping criteria, depending on whether $\theta_{k}$ has moved to the local neighborhood of $\theta^{*}$. If $\theta_{k}$ is still outside of the local area, the algorithm would finish the power-one test; otherwise, the algorithm may terminate earlier before stopping the power-one test.
Lemma 3.9 first characterizes the expected number of queries needed for testing $\theta$.
Lemma 3.9 (Power-one test)
For any $\theta\neq\theta^{*}$,
$\mathbb{E}[\tau_{1}^{\uparrow}(\theta)]\leq c_{1}|2\tilde{p}(\theta)-1|^{-2}\ln(|2\tilde{p}(\theta)-1|^{-1})+c_{2}\,,$
for some $c_{1},c_{2}>0$.
Lemma 3.9 implies that the expected number of vertical moves increases in the order of $O((2\tilde{p}(\theta)-1)^{-2})$ when $\tilde{p}(\theta)$ approaches 1/2. When $\tilde{p}(\theta)$ is bounded away from $1/2$, the expected vertical move is also bounded. Lemma 2.6 shows that $\tilde{p}(\theta)$ is strictly larger than 1/2 for any $\theta\neq\theta^{*}$, then according to Heine–Borel theorem, there exists $\delta^{\prime}>0$ such that $\tilde{p}(\theta)>1/2+\delta^{\prime}$ for all $\theta\in[-\beta_{\Theta},\theta^{*}-a]\cup[\theta^{*}+a,-\beta_{\Theta}]$ (see the proof of Lemma 8.9 in Appendix 8.1). For the ease of notation, we define the bound of vertical moves
$\phi(\tilde{p})=c_{1}|2\tilde{p}-1|^{-2}\ln(|2\tilde{p}-1|^{-1})+c_{2}.$
The selection accuracy bounded away from $1/2$ implies that $\max_{a\leq\|\theta-\theta^{*}\|\leq\beta_{\Theta}}\phi(\tilde{p}(\theta))$ is bounded. Thus, we can infer the expected number of steps—comprising both horizontal and vertical movements—outside of $[\theta^{*}-a,\theta^{*}+a]$ as follows.
Proposition 3.10 (Sample complexity of Phase I)
The expected steps outside of $[\theta^{*}-a,\theta^{*}+a]$ is bounded by
$\mathbb{E}\left[\sum_{k=1}^{\psi(a)}T^{\uparrow}(\theta_{k})\boldsymbol{1}(\|\theta_{k}-\theta^{*}\|\geq a)\right]\leq\left(\frac{4\theta^{*}}{r_{2}\alpha a}+\frac{4\beta_{1}}{r_{1}}+\frac{2\beta_{2}}{r_{R}}\right)\max_{a\leq\|\theta-\theta^{*}\|\leq\beta_{\Theta}}\phi(\tilde{p}(\theta))\,.$
Proposition 3.10 highlights that the sample complexity of Phase I hinges on the choice of $a$ and the span of $\theta$ from the prior distribution. If $\beta_{\Theta}$ is overly expansive, there is a high probability that both selected parameters $\theta_{1}$ and $\theta_{2}$ are distant from the true parameter. Consequently, the accuracy of human selection may diminish considerably, resembling random selection, as illustrated in Example 3.2. Thus, by narrowing the range of $\theta$ based on supplementary information, the sample complexity of Phase I can be notably diminished.
Once Phase I ends, $\theta_{k}$ moves to the $a$-neighborhood of $\theta^{*}$. In scenarios where both choices are close to $\theta^{*}$, discerning the superior parameter becomes more challenging. Depending on whether the selection accuracy converges towards $1/2$ as the detected parameter approaches the true parameter, we classify the problem into two cases: 1) The selection accuracy is bounded away from 1/2 for all $\theta\neq\theta^{*}$; 2) the section accuracy approaches $1/2$. The second case is harder and requires more delicate analysis. We first characterize the sample complexity for case 1) and then analyze case 2).
For the scenario where $\tilde{p}$ is bounded away from $\underline{p}>1/2$, combining the bound on the vertical moves (Proposition 3.10) with the complexity of the horizontal moves (Theorem 3.7), Theorem 3.11 characterizes the total sample complexity.
Theorem 3.11 (Sample complexity when $\underline{p}$ exists)
If $\tilde{p}(\theta)\geq\underline{p}>1/2$ for all $\theta\in\Theta$, then the expected samples collected before algorithm termination is
$\mathbb{E}[T^{+}_{T^{\rightarrow}}]\leq\tau^{\rightarrow}(\delta,\varepsilon)\cdot\phi(\underline{p})=O(\log(1/(\delta\varepsilon))).$
Next, we address the more challenging scenario in which the selection accuracy tends towards $1/2$ as the tested parameter approaches the true parameter.
3.4.1 When the selection accuracy approaches $1/2$.
As the comparison accuracy approaches $1/2$, the vertical move (power-one test) requires collecting more samples. The rate at which the vertical sample complexity increases depends on the rate at which the utility changes as $\theta$ converges to $\theta^{*}$. To account for this effect, we introduce the following assumption.
{assumption}
For any $\theta$ such that $\|\theta-\theta^{*}\|\leq a$, there exists $\lambda_{\Delta}>0$ and $0<\kappa\leq 1$ such that
$|u(c_{\Delta}^{+}(\theta))-u(c_{\Delta}^{-}(\theta))|\geq\lambda_{\Delta}\|\theta-\theta^{*}\|^{\kappa}\,.$
The parameter $\kappa$ characterizes the speed of convergence, with a smaller value indicating a faster convergence rate. It is natural to assume $\kappa\leq 1$ for several reasons. Firstly, human decision-making should not be inferior to learning from a random sample; otherwise, collecting random samples would suffice. We demonstrate this by examining the selection accuracy of learning from a single sample. Suppose the sample $(x,y)$ is generated from $y=\theta^{*}x+\epsilon$ where $\epsilon\sim\mathcal{N}(0,\sigma^{2})$. When comparing $\theta+\Delta$ and $\theta-\Delta$ (without loss of generality, assuming $\theta<\theta^{*}$), $\theta+\Delta$ is more preferable than $\theta-\Delta$ if the residual is smaller. That is, the probability of $\theta+\Delta$ being better is
$\mathbb{P}(\theta+\Delta\succeq\theta-\Delta)$ $=\mathbb{P}\left((\theta^{*}x+\epsilon-(\theta+\Delta)x)^{2}\leq(\theta^{*}x+\epsilon-(\theta-\Delta)x)^{2}\right)$
$=\mathbb{P}(\epsilon\geq(\theta-\theta^{*})|x|)=\frac{1}{2}+\Omega((\theta^{*}-\theta)|x|).$
If we use $B$ i.i.d. samples, then the probability of $\theta+\Delta$ being better is $\frac{1}{2}+\sqrt{B}\Omega((\theta^{*}-\theta)|x|)$. The detailed proof is included in Appendix 9.2. In instances where evaluation is conducted solely on a random sample, $\kappa$ is equal to 1. Considering that humans possess additional side information and can make better comparisons, we anticipate $\kappa$ to be, at most, 1.
Lemma 3.12 (How selection accuracy converges)
Under Assumption 3.4.1, it holds that
$\tilde{p}(\theta)-1/2\geq c\|\theta-\theta^{*}\|^{\kappa}\,,\forall\theta\in[\theta^{*}-a,\theta^{*}+a],$
where $c=\min\{\lambda_{\Delta}/(8\gamma\ln(2)),1/(6a)\}$.
Under Assumption 3.4.1, $\tilde{p}(\theta)$ converges to 1/2 with the speed no faster than the linear convergence. Therefore, when $\theta$ moves to the $a$-neighborhood of $\theta^{*}$, the distance between $\theta$ and $\theta^{*}$ can be bounded by $((\tilde{p}(\theta)-1/2)/c)^{1/\kappa}$. In another word, the closeness between $\tilde{p}(\theta)$ and 1/2 implies the closeness between $\theta$ and $\theta^{*}$. Thus, we can define $\tau^{\uparrow}(\delta,\varepsilon)$ as the threshold of the vertical test such that if the vertical move $T^{\uparrow}(\theta)$ has reached $\tau^{\uparrow}(\delta,\varepsilon)$, then $\theta$ is close enough to $\theta^{*}$.
Theorem 3.13 (Vertical moves in Phase II)
Suppose Assumption 3.4.1 holds. When $T^{\rightarrow}\geq\tau_{1}^{\rightarrow}(a,\delta/2)$ and $T^{\uparrow}(\theta_{T^{\rightarrow}})\geq\tau^{\uparrow}(\delta/2,\varepsilon)$ where
$\tau^{\uparrow}(\delta,\varepsilon)=\max\left\{\tau_{0},\frac{8\log(1/\delta)}{c^{2}\varepsilon^{2\kappa}}\right\}=O(\log(1/\delta)\varepsilon^{-2\kappa})\,,$
and $\tau_{0}=\{\max\ s:\hbar_{s}/s\geq c\varepsilon^{\kappa}/2\}$ where $c$ is defined in Lemma 3.12, then we have
$\mathbb{P}(\|\theta_{T^{\rightarrow}}-\theta^{*}\|\leq\varepsilon)\geq 1-\delta\,.$
When Assumption 3.4.1 holds and $\theta$ moves to the neighborhood of $\theta^{*}$, the sample complexity of the vertical move would not exceed $O(\varepsilon^{-2\kappa})$ where $\varepsilon$ is the precision. The vertical sample complexity increases with the value of $\kappa$. The faster the selection accuracy converges to 1/2, the harder the selection problem, and thus the higher the vertical sample complexity. In the extreme case where $\kappa=0$, $\tilde{p}(\theta)$ is bounded away from 1/2 so $\tau^{\uparrow}(\delta,\varepsilon)$ is bounded. This conclusion is consistent with what we have in Theorem 3.11.
3.5 Sample Complexity
Combining the sample complexity of horizontal moves and vertical moves, we conclude the total sample complexity of APB.
Theorem 3.14 (Sample complexity of APB)
Under Assumption 3.4.1, the expected number of moves is bounded by
$\mathbb{E}\left[\sum_{k=1}^{T^{\rightarrow}}T^{\uparrow}(\theta_{k})\right]\leq H(\delta,\varepsilon;a)=\tilde{O}(\varepsilon^{-2\kappa}),$
where
$H(\delta,\varepsilon;a)=\left(\frac{4\theta^{*}}{r_{2}\alpha a}+\frac{4\beta_{1}}{r_{1}}+\frac{2\beta_{2}}{r_{R}}\right)\max_{a\leq\|\theta-\theta^{*}\|\leq\beta_{\Theta}}\phi(\tilde{p}(\theta))+\max\{\tau^{\rightarrow}(\delta,\varepsilon),\tau^{\rightarrow}(\delta/2,a)\}\tau^{\uparrow}(\delta/2,\varepsilon).$
Theorem 3.14 analyzes the (human) sample complexity in a one-dimensional space, which is in the order of $\tilde{O}(\varepsilon^{-2\kappa})$. The horizontal move is in the logarithm order, so the dominant part comes from the vertical moves in the neighborhood of $\theta^{*}$. If human have the ability to differentiate between any two parameters $\theta_{1}$ and $\theta_{2}$ with probability strictly higher than $1/2+\delta^{\prime}$ as long as $\theta_{1}\neq\theta_{2}$, then the total complexity is in the logarithm order of both confidence $1/\delta$ and precision $1/\varepsilon$.
In a $d$-dimensional space, if naively refining parameters along each dimension to make the parameter within a $\varepsilon$-ball (in $\ell_{p}$-norm), then the total complexity is $dH(\delta/d,\varepsilon/d^{1/p};a)$, which scales as $\tilde{O}(d^{1+\frac{2\kappa}{p}}/\varepsilon^{2\kappa})$. However, in a high-dimensional space, $d$ can be very large while the true parameter may lie in a low-dimensional manifold. However, in high-dimensional spaces, $d$ can be significant even though the true parameter may lie within a low-dimensional manifold. In the following section, we explore leveraging a supervised learning oracle to reduce sample complexity.
4 Alignment: Human-AI Interaction
In the last section, we discussed the framework of purely using human feedback in a one-dimensional space. In practice, a noisy labeled dataset can usually be acquired easily. A noisy labeled dataset can be potentially utilized for learning the low-dimensional representation $\varphi(\cdot)$ and thus reducing the complexity of the refining step. In this section, we propose a two-stage framework where the first stage learns the low-dimensional representation using noisy labels and the second stage refines the model using human feedback.
4.1 Two-stage Framework
Suppose the true parameter lies in a $s$-dimensional space where $s\ll d$. Given confidence $\delta$ and precision $\varepsilon$, we introduce a two-stage framework SFT+HC:
  1. (I)
    In Stage 1, the algorithm feeds $N_{1}^{HAI}(\delta,\varepsilon)$ noisy labeled data $(\boldsymbol{X},Y)$ to the supervised learning oracle (e.g., Lasso), where $\boldsymbol{X}\in\mathbb{R}^{N_{1}^{HAI}\times d}$ is the design matrix and $Y$ is the noisy response. At the end, we learn the $s$-dimensional embedding $\varphi(\cdot)$;
  2. (II)
    In Stage 2, the algorithm asks human to refine the estimator along $s$ dimensions. This stage requires $N_{2}^{HAI}(\delta,\varepsilon)$ data points.
Algorithm 3 SFT+HC (Supervised Fine-Tuning + Human Comparison)
1:input: confidence level $\delta$, precision $\varepsilon$; initial distribution $\mu_{0}$; local parameter $a>0$; $k=0$; $\eta$;
2:Stage 1: Learn the embedding $\varphi(\cdot)$ through SFT oracle trained on data $(\boldsymbol{X},Y)$ with size $N_{1}^{HAI}(\delta,\varepsilon)$;
3:Stage 2: Execute Algorithm 2 (APB) by querying $N_{2}^{HAI}(\delta,\varepsilon)$ samples.
The framework design is grounded in the insight that certain representation learning techniques, like autoencoders or deep learning architectures, are inherently robust to noisy labels (Li et al. 2021, Taghanaki et al. 2021). Stage 1 is dedicated to uncovering underlying patterns and structures within the data. Given the effectiveness of human expertise in model refinement, Stage 2 endeavors to enhance model alignment through targeted sample querying. In what follows, we specifically discuss sparse linear models (Example 2.1), where Stage 1 applies Lasso [^3] OLS and thresholding may serve the same purpose as Lasso. to select important features.
4.2 Illustration on Linear Models
To demonstrate the sample complexity reduction for the two-stage framework compared to purely trained by supervised learning and provide theoretical justifications, we focus on sparse linear models:
$Y_{i}=\left\langle\boldsymbol{\vartheta^{*}},X_{i}\right\rangle+\epsilon_{i}=\left\langle\boldsymbol{\theta}^{*},\varphi(X_{i})\right\rangle+\epsilon_{i},$
where $\varphi(\cdot)$ projects $X_{i}$ to all important features; $S$ is the index set of all important features; the parameter $\boldsymbol{\theta}^{*}$ satisfies that $|\theta^{*}_{i}|\geq\underline{\beta}_{\Theta}>0$ for all $i\in S$. In the initial stage, we employ Lasso for the selection of significant features, followed by human comparisons for alignment. To assess the sample complexity of our two-stage framework, we will sequentially quantify the size of samples needed for two stages. Firstly, we establish the following standard assumptions for the analysis of Lasso.
{assumption}
[sub-Gaussian noise] The observational noise is zero-mean i.i.d. sub-Gaussian with parameter $\sigma$.
{assumption}
For the random design matrix $\boldsymbol{X}\in\mathbb{R}^{n\times d}$ input in the first stage, we assume the following three conditions holds:
  1. (I)
    Mutual incoherence: $\max_{j\in S^{c}}\ \|(\boldsymbol{X}_{S}^{\top}\boldsymbol{X}_{S})^{-1}\boldsymbol{X}_{S}^{\top}x_{j}\|_{1}\leq\alpha_{1}.$
  2. (II)
    Lower eigenvalue: $\lambda_{\min}\left(V_{S}\right)\geq\lambda_{0}>0,$ where $V_{S}=\frac{1}{n}\boldsymbol{X}_{S}^{\top}\boldsymbol{X}_{S}$.
  3. (III)
    $\ell_{\infty}$-curvature condition: $\|Vz\|_{\infty}\geq\alpha_{2}\|z\|_{\infty}\,,\text{ for all }z\in C_{\alpha^{\prime}}(S),$ where $C_{\alpha^{\prime}}(S)=\{\Delta\in\mathbb{R}^{d}|\|\Delta_{S^{c}}\|_{1}\leq\alpha^{\prime}\|\Delta_{S}\|_{1}\}$.
Assumption 4.2 is standard for the analysis of Lasso. Consider a random design matrix $\boldsymbol{X}\in\mathbb{R}^{n\times d}$ with i.i.d. $\mathcal{N}(0,1)$ entries for example, we can easily verify that it satisfies mutual incoherence, lower eigenvalue, and $\ell_{\infty}$-curvature condition. Under Assumption 4.2 and 4.2 and based on Corollary 7.22 in Wainwright (2019), we derive the following theorem that bounds the infinite distance between vector $\hat{\theta}_{S}$ and $\theta^{*}_{S}$.
Theorem 4.1
Under Assumption 4.2 and 4.2, suppose that we solve the Lagrangian Lasso with regularization parameter
$\lambda_{n}=\frac{2C\sigma}{1-\alpha_{1}}\left\{\sqrt{\frac{2\log(d-s)}{n}}+\zeta\right\}$
for $\zeta=\frac{1}{4}\min\left\{\frac{\sqrt{c_{\min}}}{\sigma\underline{\beta}_{\Theta}},\frac{\alpha_{2}(1-\alpha_{1})}{2C\sigma\underline{\beta}_{\Theta}}\right\}$. Then the optimal solution $\hat{\theta}$ is unique with its support contained within $S$, and satisfies the $\ell_{\infty}$-error bound
$\|\hat{\theta}_{S}-\theta^{*}_{S}\|_{\infty}\leq\underbrace{\frac{\sigma}{\sqrt{c_{\min}}}\left\{\sqrt{\frac{2\log s}{n}}+\zeta\right\}+\frac{1}{\alpha_{2}}\lambda_{n}}_{B_{n}}\,,$
all with probability at least $1-4e^{-\frac{n\zeta^{2}}{2}}$. Moreover, there is no false inclusion: The solution has its support set $\hat{S}$ contained with the true support set $S$.
Remark 4.2
Not limited to the linear function, Lasso can also be applied to the nonlinear function class (Plan and Vershynin 2016) or nonparametric variable selection (Li et al. 2023). Moreover, there is a rich literature of representation learning that can be applied in the first stage (Bengio et al. 2013).
Theorem 4.1 implies that as long as $B_{n}\leq\underline{\beta}_{\Theta}$, the variable selection is consistent. Since $\zeta=\frac{1}{4}\min\left\{\frac{\sqrt{c_{\min}}}{\sigma\underline{\beta}_{\Theta}},\frac{\alpha_{2}(1-\alpha_{1})}{2C\sigma\underline{\beta}_{\Theta}}\right\}$, we only need
$\frac{\sigma}{\sqrt{c_{\min}}}\sqrt{\frac{2\log s}{n}}+\frac{1}{\alpha_{2}}\cdot\frac{2C\sigma}{1-\alpha_{1}}\sqrt{\frac{2\log(d-s)}{n}}\leq\frac{1}{2}\underline{\beta}_{\Theta}.$
Thus if there are $n$ noisy samples in which
$n\geq\max\left\{\frac{8\sigma^{2}\log s}{\underline{\beta}_{\Theta}^{2}c_{\min}},\frac{32C^{2}\sigma^{2}\log(d-s)}{(\underline{\beta}_{\Theta}\alpha_{2}(1-\alpha_{1}))^{2}}\right\},$
we are able to select $s$ variables with probability at least $1-4e^{-\frac{n\zeta^{2}}{2}}$. Define
$H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})=\max\left\{\frac{8\sigma^{2}\log s}{\underline{\beta}_{\Theta}^{2}c_{\min}},\frac{32C^{2}\sigma^{2}\log(d-s)}{(\underline{\beta}_{\Theta}\alpha_{2}(1-\alpha_{1}))^{2}},\frac{2\log(4/\delta)}{\zeta^{2}}\right\}=O(\sigma^{2}\log d/\underline{\beta}_{\Theta}^{2}).$
Then, when having more than $H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})$ noisy labeled data at hand, we can have the correct feature selection with probability at least $1-\delta$. Moreover, if we want to do one more step of refinement to restrict the uncertainty set to be within $\mathfrak{r}$-distance where $\mathfrak{r}<\underline{\beta}_{\Theta}$, then we need $H_{0}(\delta,\varepsilon;\mathfrak{r})$ noisy labeled data. Note that $\mathfrak{r}$, which represents the feasible range for refinement, is the parameter that we can choose. The trade-off exists. A smaller value of $\mathfrak{r}$ demands more noisy labels in the initial stage. However, limiting the feasible region to a smaller area could conserve samples required for human comparisons. As previously noted, when both options significantly deviate from the true model, distinguishing the superior one becomes challenging. This challenge implies a greater need for samples to ascertain the correct answer. Consequently, in certain scenarios, opting for $\mathfrak{r}\leq\underline{\beta}_{\Theta}$ to confine the refinement region may be more optimal. Specifically, we select $\mathfrak{r}$ to minimize the total sample complexity:
$\mathfrak{r}^{*}=\min_{0<\mathfrak{r}\leq\underline{\beta}_{\Theta}}H_{0}(\delta,\varepsilon;\mathfrak{r})+H(\delta,\varepsilon;\mathfrak{r}),$ (4.1)
where $H(\delta,\varepsilon;\mathfrak{r})$ is defined in Theorem 3.14. It is important to emphasize that while the first part depends on the noise $\sigma$, the second part does not rely on $\sigma$ but rather on the detection accuracy $\phi(\tilde{p}(\theta))$. Hence, the optimal $\mathfrak{r}^{*}$ is contingent on the function of detection accuracy. For $\theta$ considerably distant from the true parameter $\theta^{*}$, $\tilde{p}(\theta)$ could potentially approach $1/2$, leading to a substantial value for $\phi(\tilde{p}(\theta))$.
4.3 Complexity of SFT+HC
The optimal value of $\mathfrak{r}^{*}$ can be obtained numerically by enumerating and comparing the objectives in Equation (4.1). To establish an upper bound on the sample complexity of the two-stage framework to analytically show the benefit, we set $\mathfrak{r}=\underline{\beta}_{\Theta}$. In this case, we have
$H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})=\tilde{O}(\sigma^{2}/\underline{\beta}_{\Theta}^{2})\quad\text{and}\quad H(\delta,\varepsilon;\underline{\beta}_{\Theta})=\tilde{O}\left(s\bar{\phi}(a,\underline{\beta}_{\Theta})+s^{1+\frac{2\kappa}{p}}/\varepsilon^{2\kappa}\right),$
where $\bar{\phi}(a,D)=\max_{a\leq\|\theta-\theta^{*}\|\leq\underline{\beta}_{\Theta}}\phi(\tilde{p}(\theta))$. Theorem 4.3 is a direct conclusion from Theorem 3.14 and 4.1.
Theorem 4.3 (Computational complexity of SFT+HC)
Under Assumption 4.2 and 4.2, by setting $N_{1}^{HAI}(\delta)=H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})$ and then running Algorithm 2 with parameter range $[-\underline{\beta}_{\Theta},\underline{\beta}_{\Theta}]$, it holds that $\|\hat{\theta}-\theta^{*}\|_{p}\leq\varepsilon$ with probability at least $1-2\delta$. Moreover, under Assumption 3.4.1, the expected sample complexity is
$N_{1}^{HAI}(\delta,\varepsilon)+N_{2}^{HAI}(\delta,\varepsilon)=\tilde{O}(\sigma^{2}/\underline{\beta}_{\Theta}^{2}+s\bar{\phi}(a,\underline{\beta}_{\Theta})+s^{1+\frac{2\kappa}{p}}/\varepsilon^{2\kappa}).$
Theorem 4.3 demonstrates the sample complexity of the two-stage framework SFT+HC. When the distance is measured in two-norm (i.e., $p=2$), the complexity is in the order of $\tilde{O}(\sigma^{2}/\underline{\beta}_{\Theta}^{2}+s\bar{\phi}(a,\underline{\beta}_{\Theta})+s^{1+\kappa}/\varepsilon^{2\kappa})$. Instead, if exclusively utilizing Lasso for the supervised learning oracle to refine the estimator based on $n$ noisy labeled data, subject to certain regularity conditions, the estimation error can be bounded as $\|\hat{\theta}_{n}-\theta^{*}\|_{2}\leq c\sigma\sqrt{s\log d/n}$ with high probability for some constant $c>0$. Equivalently, to ensure that the two-norm distance of the estimation error is within $\varepsilon$, we need to acquire $\tilde{O}(\sigma^{2}s\log d/\varepsilon^{2})$ data points.
Comparing the sample complexity of SFT+HC characterized by Theorem 4.3 for $p=2$, the condition for SFT+HC requiring less sample complexity than SFT is characterized in Proposition 4.4.
Proposition 4.4
For sufficiently small $\varepsilon$, SFT+HC requires less sample complexity than pure SFT for solving the $(\varepsilon,\delta)$-alignment problem when (ignoring the logarithm term)
$\frac{\sigma^{2}}{s^{\kappa}}\gtrapprox\varepsilon^{2-2\kappa}.$ (4.2)
We call the ratio $\sigma^{2}/s^{\kappa}$ as label noise-to-comparison-accuracy ratio (LNCA ratio) and condition (4.2) as LNCA condition. This condition involves two important parameters: the observational noise $\sigma$ and the convergence rate of the accuracy $\kappa$. When the comparison accuracy is bounded away from 1/2, we have $\kappa=0$ and the condition reduces to $\sigma\gtrapprox\varepsilon$, which holds naturally true. In the worst case where $\kappa=1$, the condition becomes $\sigma^{2}\gtrapprox s$. When sample observational noise is higher, the benefit of the extra human comparison step becomes more apparent.
5 Practical Implementations
In the preceding sections, we established a framework that assumes humans can compare any two parametric models. However, such a comparison could be possibly challenging when humans only have access to the models themselves. In practice, it is often more feasible to compare models based on their performance with sampled data, allowing humans to discern which one aligns better with the true underlying model. To address this specific challenge, this section delves into the practical aspects of comparing two models using selected samples.
5.1 Active Learning: How to Select Samples for Comparison
Active learning focuses on the strategic selection of samples to be labeled or compared in our context, with the aim of constructing prediction models in a resource-efficient manner. The key idea is to assess the significance of a sample to determine the value of acquiring its labels. Algorithm 2 actively selects a pair of two models to ask for comparison. In practical scenarios, it often involves soliciting human judgment to compare the responses predicted by two different models for a given sample data, rather than directly comparing their model parameters.
We start with an example. The clinician needs to evaluate the Atherosclerotic Cardiovascular Disease (ASCVD) risk of patients $f_{\boldsymbol{\theta}}$, which is a linear function of contextual information, denoted as $x$, such as demographics (age, gender, racial/ethnic group, and education level), baseline conditions (body mass index, history of ASCVD events, and smoking status), clinical variables (glycosylated hemoglobin (A1c) and systolic blood pressure (sbp)), anti-hypertensive agenets, anti-hyperglycemic agents.
Given sample of patient context $x\in\mathcal{X}$, the clinician evaluates the ASCVD risk according to their expertise. The distance between model $f_{\boldsymbol{\theta}}$ and $f_{\boldsymbol{\theta}^{*}}$ evaluated on sample $x$ can be defined as the absolute value of the difference:
$d_{x}(\boldsymbol{\theta},\boldsymbol{\theta}^{*})=|\varphi(x)^{\top}\boldsymbol{\theta}-\varphi(x)^{\top}\boldsymbol{\theta}^{*}|.$
Then the utility of $f_{\boldsymbol{\theta}}$ on sample $x$ is $u_{x}(\boldsymbol{\theta})=-d(\boldsymbol{\theta},\boldsymbol{\theta}^{*}|x)=-|\varphi(x)^{\top}\boldsymbol{\theta}-\varphi(x)^{\top}\boldsymbol{\theta}^{*}|$. When presented with risk levels predicted by $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$, the clinician would typically opt for the one that is closer to their assessment.
Let us first consider a simple scenario. When $\mathcal{X}$ contains adequate variety of samples such that for any base vector $e_{i}$, there exists sample $x$ such that $\varphi(x)=e_{i}$, then by comparing $f_{\boldsymbol{\theta}_{1}}$ with $f_{\boldsymbol{\theta}_{2}}$ on such sample $x$, we have
$d_{x}(\boldsymbol{\theta}_{1},\boldsymbol{\theta}^{*})=|\theta_{1i}-\theta_{i}^{*}|\quad\text{and}\quad d_{x}(\boldsymbol{\theta}_{2},\boldsymbol{\theta}^{*})=|\theta_{2i}-\theta_{i}^{*}|.$
It implies that the comparison of $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$ when testing on $x$ would be solely on the $i^{th}$ dimension. Thus, the alignment process for the parameter in different dimensions can be decoupled, and SFT+HC can be directly applied where the Stage 2 alignment is proceeded along each dimension.
However, in practice, the dataset $\mathcal{X}$ may be limited, potentially leading to situations where there are no samples $x\in\mathcal{X}$ that satisfy the condition $\varphi(x)=e_{i}$. In such cases, it becomes impractical to refine the model parameter along each dimension separately. The question then arises: how can we efficiently select samples to refine the model parameter? To address this challenge, we propose a procedure for constructing a new basis using the available samples in $\mathcal{X}$. This procedure aims to optimize the alignment process by leveraging the existing data to guide the selection of samples for parameter enhancement.
5.2 New Basis Construction through Gram Schmidt Process
To facilitate the independent alignment of each coordinate, we aim to identify a new set of orthogonal basis, which can be constructed using $\Psi:=\{\varphi(x):x\in\mathcal{X}\}$. Let $\boldsymbol{\alpha}_{1},\cdots,\boldsymbol{\alpha}_{s}$ represent these new basis, and our objective is to learn $\boldsymbol{\theta}^{*}=\sum_{i=1}^{s}\omega_{i}^{*}\boldsymbol{\alpha}_{i}$, where $\omega_{i}^{*}$ are the corresponding coefficients. We describe the process of constructing the orthogonal basis $\{\boldsymbol{\alpha}_{1},\cdots,\boldsymbol{\alpha}_{s}\}$ from the sample space.
First, we pick a set of samples that spans the space of $\Psi$, denoted as $\varphi(x_{1}),\cdots,\varphi(x_{s})$. For simplicity, we define $\boldsymbol{z}_{i}=\varphi(x_{i})$. Then we use the Gram Schmidt process to construct a set of orthogonal basis based on $\boldsymbol{z}_{1},\cdots,\boldsymbol{z}_{s}$:
$\boldsymbol{\alpha}_{k}=\boldsymbol{z}_{k}-\sum_{j=1}^{k-1}\frac{\boldsymbol{z}_{k}^{\top}\boldsymbol{\alpha}_{j}}{\boldsymbol{\alpha}_{j}^{\top}\boldsymbol{\alpha}_{j}}\boldsymbol{\alpha}_{j},\quad\forall 1\leq k\leq s.$
We sequentially align the coordinates $\omega_{1},\cdots,\omega_{s}$. When evaluating on sample $x_{1}$, the true value is represented by $\varphi(x_{1})^{\top}\boldsymbol{\theta}^{*}=\boldsymbol{z}_{1}^{\top}\boldsymbol{\theta}^{*}=\omega_{1}^{*}\boldsymbol{z}_{1}^{\top}\boldsymbol{z}_{1}$. Utilizing a probabilistic bisection process in 2, we actively select two answers for each query. Through this iterative refinement, the final value terminates at $y(x_{1})$, which is guaranteed to be within $\varepsilon$-accuracy of the true parameter $\omega_{1}^{*}$ (with high probability). In the end, the first coordinate is refined to $\hat{\omega}_{1}=y(x_{1})/(\boldsymbol{z}_{1}^{\top}\boldsymbol{z}_{1})$.
Next, the algorithm applies testing to the sample $\boldsymbol{z}_{2}=\varphi(x_{2})=\frac{\boldsymbol{z}_{2}^{\top}\boldsymbol{\alpha}_{1}}{\boldsymbol{\alpha}_{1}^{\top}\boldsymbol{\alpha}_{1}}\boldsymbol{\alpha}_{1}+\boldsymbol{\alpha}_{2}$. Note that the true value is given by
$\varphi(x_{2})^{\top}\boldsymbol{\theta}^{*}=\boldsymbol{z}_{2}^{\top}\boldsymbol{\theta}^{*}=\boldsymbol{z}_{2}^{\top}\boldsymbol{\alpha}_{1}\omega_{1}^{*}+\boldsymbol{\alpha}_{2}^{\top}\boldsymbol{\alpha}_{2}\omega_{2}^{*},$
where $w_{1}$ is refined already. Once more, employing the probabilistic bisection process, the value of $z_{2}^{\top}\boldsymbol{\theta}^{*}$ is iteratively refined to $y(x_{2})$. Upon termination of the probabilistic bisection process, we obtain the aligned value of the second coordinate as
$\hat{\omega}_{2}=\frac{y(x_{2})-\boldsymbol{z}_{2}^{\top}\boldsymbol{\alpha}_{1}\hat{\omega}_{1}}{\boldsymbol{\alpha}_{2}^{\top}\boldsymbol{\alpha}_{2}}.$
Subsequently, at the $k^{th}$ iteration, we test on the sample $\boldsymbol{z}_{k}=\boldsymbol{\alpha}_{k}+\sum_{j=1}^{k-1}\frac{\boldsymbol{z}_{k}^{\top}\boldsymbol{\alpha}_{j}}{\boldsymbol{\alpha}_{j}^{\top}\boldsymbol{\alpha}_{j}}\boldsymbol{\alpha}_{j}$ where the true value equals
$\boldsymbol{z}_{k}^{\top}\boldsymbol{\theta}^{*}=\sum_{j=1}^{k-1}\omega_{j}^{*}\boldsymbol{z}_{k}^{\top}\boldsymbol{\alpha}_{j}+\omega_{k}^{*}\boldsymbol{\alpha}_{k}^{\top}\boldsymbol{\alpha}_{k}.$
Once the value of $\boldsymbol{z}_{k}^{\top}\boldsymbol{\theta}^{*}$ is refined to $y(x_{k})$ through 2, we are able to obtain the aligned value of the $k^{th}$ coordinate as
$\hat{\omega}_{k}=\frac{y(x_{k})-\sum_{j=1}^{k-1}\hat{\omega}_{j}\boldsymbol{z}_{k}^{\top}\boldsymbol{\alpha}_{j}}{\boldsymbol{\alpha}_{k}^{\top}\boldsymbol{\alpha}_{k}}.$
Through asking human to evaluate $x_{1},\cdots,x_{s}$ and refining $\omega_{1},\cdots,\omega_{s}$ in sequence, the parameter $\boldsymbol{\theta}:=\sum_{i=1}^{s}\hat{\omega}_{i}\boldsymbol{\alpha}_{i}$ is refined within a certain accuracy level. We include the pseudo code in Algorithm 4.
At the end, it is essential to emphasize that the set of orthogonal basis is not unique, as there may exist multiple sets. Consequently, we can conduct multiple rounds of alignment, with the option to switch basis from one round to another.
Algorithm 4 ASS (Active Sample Selection)
1:input: available set of covariate $\mathcal{X}$
2:Pick a set of samples that spans the space of $\Psi$: $\varphi(x_{1}),\cdots,\varphi(x_{s})$
3:for $k=1,\cdots,s$ do
$\boldsymbol{\alpha}_{k}=\varphi(x_{k})-\sum_{j=1}^{k-1}\frac{\varphi(x_{k})^{\top}\boldsymbol{\alpha}_{j}}{\boldsymbol{\alpha}_{j}^{\top}\boldsymbol{\alpha}_{j}}\boldsymbol{\alpha}_{j}$
4:end for
5:for $k=1,\cdots,s$ do
6:     Query sample $x_{k}$ through 2 and get output $y(x_{k})$;
7:     Set $\hat{\omega}_{k}=\frac{y(x_{k})-\sum_{j=1}^{k-1}\hat{\omega}_{j}\varphi(x_{k})^{\top}\boldsymbol{\alpha}_{j}}{\boldsymbol{\alpha}_{k}^{\top}\boldsymbol{\alpha}_{k}}$
8:end for
6 Numerical Experiment
6.1 Testing Sample Complexity
In the first experiment, we compare the performance of SFT+HC and pure SFT under the same sample size.
Experiment setup. Consider a high-dimensional linear regression problem where $d=100$ and $s=10$. The first ten coordinates of the true parameter are independently drawn from the uniform distribution and the rest coordinates are zero. The observational noise is drawn from Gaussian distribution with mean zero and variance $\sigma=\{1,2,5\}$; we choose $\kappa\in[0,0.95]$. The selection precision (along each dimension) is set as $\varepsilon=0.1$ and the human expert level in the choice model is $\gamma=1$. We conducted 50 repetitions for each experiment and calculated the average across all results.
In the first stage the two-stage framework, we use lasso and select the sample size large enough such that the estimator for the non-zero coefficient can be deviated from zero. That is, we guarantee that important features are selected with high probability. Once the algorithm SFT+HC stops, we use the same sample size to train lasso and compute the estimation error. We compare the estimation error of two algorithms under the same sample size.
6.1.1 Experiment results.
Figure 3 shows the ratio of the estimation error of two algorithms (Err(SFT+HC)/Err(SFT)) with different values of $\kappa$ when $\sigma=1$. For smaller value of $\kappa$, human’s selection accuracy is higher when $\theta$ approaches $\theta^{*}$. In particular, when $\kappa=0$, the selection accuracy is bounded away from 50%. Therefore, the task is harder when $\kappa$ increases. It is consistent with the increasing trend of the estimation error when $\kappa$ grows from 0 to 1. In addition, a larger value of $\kappa$ leads to a larger sample size, so the estimation error of lasso decreases. The intersection point is around $\kappa=0.3$. When $\kappa\leq 0.3$, SFT+HC performs better than pure SFT. In particular, when $\kappa=0$, the error of two-stage framework is roughly 50% of that in SFT.
Figure 3: Error ratio of the two-stage framework/pure SFT when $\sigma=1$.
The results for noise levels with $\sigma=2$ and $\sigma=5$ are illustrated in Figures 4. Notably, the trends of the error curves exhibit similarity. The threshold, defined as the intersection point of these curves, is approximately 0.55 when $\sigma=2$. However, when $\sigma=5$, the two curves do not intersect, suggesting that SFT+HC consistently outperforms pure SFT.
A comparison across the three figures reveals that the superiority of SFT+HC becomes more pronounced in the presence of higher observational noise. In conclusion, SFT+HC is more effective under conditions of increased noise and reduced value of $\kappa$.
(a) Error ratio when $\sigma=2$.
(b) Error ratio when $\sigma=5$.
Figure 4: Prediction errors with different noise levels.
6.1.2 The variation of expert level $\gamma$.
We change the value of human expert level $\gamma$ to see how the human comparison accuracy would influence the estimation error. The smaller value of $\gamma$ indicates higher accuracy in the comparison task. In particular, when $\gamma$ approaches 0, human can always select the accurate answer. Figure 5 shows the performance of two algorithms with $\gamma=1.1$ and $\gamma=0.8$, respectively. Comparing two figures, when $\gamma=1.1$, SFT+HC performs better than pure SFT when $\kappa$ is lower than approximately 0.45; when $\gamma=0.8$, SFT+HC performs better than pure SFT when $\kappa$ is lower than approximately 0.65. Therefore, it is consistent with our theory that SFT+HC improves its performance when human has a higher expert level in the selection (i.e., the selection accuracy is higher).
(a) Error ratio when $\gamma=1.1$.
(b) Error ratio when $\gamma=0.8$.
Figure 5: Prediction errors with different expert levels.
6.1.3 The variation of important feature dimensions $s$.
The relative performance of two algorithms also depends on the dimension of the problem. For the experiment setup, we set $\gamma=1$ and $\sigma=2$. Figure 6 illustrates the algorithmic performance with important feature dimension $s=20$ and $s=50$, respectively. The figure reveals that the intersection point shifts to a smaller value as the important feature dimension increases. This observation aligns with our theoretical condition (4.2), indicating that as the important feature dimension grows, the advantage of SFT+HC diminishes.
(a) Error ratio when $s=20$.
(b) Error ratio when $s=50$.
Figure 6: Prediction errors with different important feature dimensions.
6.2 Testing LNCA Ratio
The LNCA ratio, $\sigma^{2}/s^{\kappa}$, determines the benefits of SFT+HC over pure SFT. The LNCA ratio varies between different tasks. We are interested in estimating this value for a specific task: to determine the percentage of total area enclosed by a given (red) polygon, as shown in Figure 7, for example.
We divided participants into two distinct groups. The first group was presented with the polygon figure and asked to provide their estimated percentage of the total enclosed area. The second group was shown the same figure but presented with two options for the area percentage. Their task was to select the option they believed was closer to the actual percentage of the enclosed area. For instance, in Figure 7(b), the true percentage is 50%. Human is provided with two options, 71% and 55%, so the correct choice is 55%.
This dual-group approach aimed to explore the efficacy of human estimation when directly asked to estimate the area versus when presented with comparative options. In the latter case, we varied the discrepancy between the two choices to assess how the ease or difficulty of comparison affected the accuracy of their selections. A larger disparity between the choices facilitated easier decision-making, while closer options posed a more challenging comparative task. Within Figure 7, three testing samples are depicted. Notably, sample (a) presents the most challenging comparison among these three instances due to the minimal disparity observed between them.
Fifty different tasks are launched in Amazon Mechanical Turk [^4] https://www.mturk.com/, as shown in Figure 8, and we collected 1300 data points in total. We use the following ways to estimate $\kappa$ and $\sigma$.
(a) 21% vs. 5%. True: 10%.
(b) 71% vs. 55%. True: 50%.
(c) 29% vs. 13%. True: 26%.
Figure 7: Experiment design.
(a) The first group: Select a better answer.
(b) The second group: Estimate the percentage.
Figure 8: Comparison vs. Estimation.
Estimation process for $\kappa$. Given figures with true parameter values $\theta^{*}$ representing the percentage of total area enclosed by a polygon, we aim to estimate the accuracy convergence rate $\kappa$ between two options using a human choice model. We collect multiple samples for each figure and then use these estimates to derive the relationship between the difference in utilities and the accuracy of human choices.
More specifically, among two choices, assume the smaller one is $c^{-}$ and the larger one is $c^{+}$, where the middle point is $\theta=(c^{-}+c^{+})/2$. According to the choice model, human makes the correct choice with accuracy $p(\theta)=\frac{1}{1+\exp(-|u(c^{-})-u(c^{+})|/\gamma)}$. Note that we try to estimate $\kappa$ from Assumption 3.4.1 that
$|u(c^{+})-u(c^{-})|/\gamma\approx\lambda_{\Delta}/\gamma\cdot\|\theta-\theta^{*}\|^{\kappa}\,.$
That is, the accuracy is $p(\theta)=\frac{1}{1+\exp(-\lambda_{\Delta}/\gamma\cdot|\theta-\theta^{*}|^{\kappa})}$. Let $\tilde{\lambda}=-\lambda_{\Delta}/\gamma$. For each data point with tested point $\theta_{i}$ and true parameter $\theta_{i}^{*}$, we define $y_{i}=1$ (the correct choice) with probability $p_{i}(\theta)$ and $y_{i}=0$ (the wrong choice) with probability $1-p_{i}(\theta)$. By maximizing the log-likelihood, we solve the optimization problem
$\max_{\tilde{\lambda},\kappa}\quad\sum_{i=1}^{K}y_{i}\log p_{i}+(1-y_{i})\log(1-p_{i})$
which is equivalent to
$\max\quad\sum_{i=1}^{K}-y_{i}\log(1+\exp(-\tilde{\lambda}|\theta_{i}-\theta^{*}|^{\kappa}))-(1-y_{i})\log(1+\exp(\tilde{\lambda}|\theta_{i}-\theta_{i}^{*}|^{\kappa}))\,.$
From the data collected from Amazon Mechanical Turk, the estimated $\kappa$ is 0.095.
Estimation process for $\sigma$. The other group is asked to directly estimate the percentage of total area. For the figure with true parameter values $\theta_{i}^{*}$, human provides responses by $y_{i}=\theta_{i}^{*}+\epsilon_{i}$. We estimate the variance of the noise by $\hat{\sigma}^{2}=\sum_{i=1}^{N}\frac{(y_{i}-\theta_{i}^{*})^{2}}{N-1}=29.04$.
In the context of a one-dimensional problem, as evidenced by numerical experiments detailed in the preceding section, SFT+FTPB demonstrates a substantial performance advantage over SFT when $\kappa\approx 0.1$ and $\sigma\approx 30$.
7 Concluding Remarks
This work is motivated by the ever-growing observations of human-AI interaction in LLM. We presented a theoretical two-stage framework for explaining the significance of using human comparisons for the model improvement. While the noisy-labeled data is fed into the SFT procedure to learn the low-dimensional representation in the first stage, we sequentially inquire human evaluators to make pair-wise comparisons in the second stage. To address the challenge of efficient acquisition of valuable information from these human comparisons, we introduced a probabilistic bisection method, which factors in the uncertainty arising from the accuracy of human comparisons. The newly introduced concept, label-noise-to-comparison-accuracy ratio, quantifies the relative scale between label noise and human comparison accuracy. We demonstrate the significant advantages of our proposed two-stage framework over pure supervised learning approach under certain conditions.
Our work represents an early attempt to organically integrate machine learning with human knowledge. We consider this paper as a prompt for an open discussion. Some potential extensions to our work are worth investigating. Firstly, we extended the probabilistic bisection algorithm from one dimension to multiple dimensions by refining along each dimension. This expansion raises the question of whether a more efficient bisection algorithm exists in multi-dimensional spaces. As discussed in Frazier et al. (2019), the development of a PBA tailored for multi-dimensional problems remains an open problem. Secondly, although our paper focuses on a prediction problem, the framework has the potential to be extended to the action-based learning, where the goal is to select optimal actions for the decision-making problem. Thirdly, exploring the optimal sample selection to maximize the learning algorithm’s efficiency is a compelling area of inquiry. While we have outlined guidelines for practical sample selection, the quest for the most efficient proposed procedure remains open.

### v2

@@
 Junyu Cao and Mohsen Bayati
+[…]
 […]
@@
 […]
-A Probabilistic Approach for Alignment with Human Comparisons
+A Probabilistic Approach for Model Alignment with Human Comparisons
 […]
-A Probabilistic Approach for Alignment with Human Comparisons
+A Probabilistic Approach for Model Alignment with Human Comparisons
 […]
@@
 A growing trend involves integrating human knowledge into learning frameworks, leveraging subtle human feedback to refine AI models.
-Despite these advances, no comprehensive theoretical framework describing the specific conditions under which human comparisons improve the traditional supervised fine-tuning process has been developed.
-To bridge this gap, this paper studies the effective use of human comparisons to address limitations arising from noisy data and high-dimensional models.
-We propose a two-stage “Supervised Fine Tuning+Human Comparison” (SFT+HC) framework connecting machine learning with human feedback through a probabilistic bisection approach.
-The two-stage framework first learns low-dimensional representations from noisy-labeled data via an SFT procedure, and then uses human comparisons to improve the model alignment.
-To examine the efficacy of the alignment phase, we introduce a novel concept termed the “label-noise-to-comparison-accuracy” (LNCA) ratio.
-This paper theoretically identifies the conditions under which the “SFT+HC” framework outperforms pure SFT approach, leveraging this ratio to highlight the advantage of incorporating human evaluators in reducing sample complexity.
-We validate that the proposed conditions for the LNCA ratio are met in a case study conducted via an Amazon Mechanical Turk experiment.
+While these approaches have shown promising results in practice, the theoretical understanding of when and why such approaches are effective remains limited.
+This work takes steps toward developing a theoretical framework for analyzing the conditions under which human comparisons can enhance the traditional supervised learning process.
+Specifically, this paper studies the effective use of noisy-labeled data and human comparison data to address challenges arising from noisy environment and high-dimensional models.
+We propose a two-stage “Supervised Learning+Learning from Human Feedback” (SL+LHF) framework that connects machine learning with human feedback through a probabilistic bisection approach.
+The two-stage framework first learns low-dimensional representations from noisy-labeled data via an SL procedure and then uses human comparisons to improve the model alignment.
+To examine the efficacy of the alignment phase, we introduce a concept, termed the “label-noise-to-comparison-accuracy” (LNCA) ratio.
+This paper identifies from a theoretical perspective the conditions under which the “SL+LHF” framework outperforms the pure SL approach; we then leverage this LNCA ratio to highlight the advantage of incorporating human evaluators in reducing sample complexity.
+We validate that the LNCA ratio meets the proposed conditions for its use through a case study conducted via Amazon Mechanical Turk (MTurk).
 […]
-model alignment, supervised fine-tuning, probabilistic bisection, human-AI interaction.
+model alignment, supervised learning, probabilistic bisection, human-AI interaction.
 1 Introduction
@@
 Machine learning has emerged as the standard technique for numerous tasks, including natural language processing and computer vision, among others.
-Its great importance has also been demonstrated across various domains of operations research.
-Traditional methodologies have predominantly relied on supervised fine-tuning (SFT) [^1] Supervised fine-tuning is usually used in Large Language Models (LLM).
-We use supervised fine-tuning and supervised learning interchangeably in this paper.
-techniques by training on noisy-labeled data.
-A common practice in this context for decision-makers is to directly adopt the output of the best predictor identified by SFT, where the best predictor can be defined by the minimizer of the empirical risk, for example.
-However, some issues may arise in this approach.
-One significant challenge emerges from the prevalence of large noise, which can substantially hinder model performance.
-On the other hand, to mitigate the model misspecification issue, the growing trend toward increasingly complex models may lead to over-parameterization.
-Over-parameterization presents additional obstacles to the design of efficient and accurate AI systems.
-In instances where the model exhibits high complexity, such as when the feature dimension is comparable to or exceeds the number of data points, multiple distinct estimators may exhibit seemingly equivalent empirical performance on the training dataset but perform poorly on unseen data, a scenario exacerbated by the inherent under-specification of models (D’Amour et al.
+Various streams of operations research also demonstrate its importance.
+To model increasingly complex environments, machine learning models have grown in complexity, often leading to over-parameterization.
+While this increased complexity allows models to capture intricate patterns, it has revealed fundamental limitations in the standard training and testing pipeline.
+Specifically, the traditional supervised learning (SL) approach of training on noisy-labeled data faces significant challenges due to model misspecification (D’Amour et al.
 2022).
-Thus, decision-makers must consider strategic data utilization and refine models to achieve a better performance under different environments.
-Recently, an ever-growing number of papers have incorporated human-knowledge into the learning framework (Christiano et al.
-2017, Li et al.
-2016).
-Fine-tuning with human feedback has shown a promising direction.
-For instance, in the design of chatbots, the performance can be adaptively and subsequently improved through the interactions with humans (Ziegler et al.
+This issue becomes particularly acute in high-complexity scenarios, such as when the feature dimension is comparable to or exceeds the number of training data points.
+In such high-dimensional settings, fundamentally different models may achieve similar performance metrics during testing, despite having different underlying structures and behaviors.
+For instance, as demonstrated by Overman et al.
+(2024), the trained model that minimizes the empirical risk may violate fundamental principles, such as the natural expectation of negative demand elasticity—indicating that the model is not aligned with its expected behavior.
+Consequently, decision makers must consider strategic approaches to achieve better model alignment across different environments.
+One promising approach has been the incorporation of human feedback into the learning framework (Li et al.
+2016, Christiano et al.
+2017, Alur et al.
+2024).
+These studies demonstrate that using human feedback for alignment—refining model parameters that are based on human-labeled data—can effectively address the limitations of traditional training approaches.
+For instance, in the design of chatbots, performance can be adaptively improved through the chatbots’ interactions with humans (Ziegler et al.
 2019, Ouyang et al.
@@
 Ouyang et al.
-(2022) focuses on fine-tuning approaches to aligning language models, where InstructGPT is fine-tuned with human feedback.
-In many cases, human evaluators possess domain expertise that goes beyond what can be encoded into a dataset or estimation method.
-Their deep understanding of the task domain enables them to make informed comparisons that consider context, relevance, and real-world applicability (Deng et al.
+(2022) exemplify this approach in their work on fine-tuning language models, where InstructGPT is fine-tuned with human feedback.
+The effectiveness of human feedback stems from evaluators’ domain expertise that extends beyond what can be encoded into a dataset or estimation method.
+Humans’ deep understanding of the task domain enables them to make informed comparisons that consider context, relevance, and real-world applicability (Deng et al.
 2020).
-“Thinking is difficult, that’s why most people judge” – Carl Jung.
-Coming up with an answer is much more difficult than selecting a better answer from pair-wise comparisons.
-Zhou and Xu (2020) use pair-wise comparisons for story generation; Xu et al.
-(2023) collect human comparisons over two sampled LLM for the algorithm fine-tuning process.
-While prior research has showcased remarkable results when integrating human feedback across various applications, it has yet to establish a comprehensive theoretical framework detailing the specific conditions under which human comparison significantly augments the conventional SFT framework.
-To build on this research trajectory and fill this gap, our paper provides a theoretical framework for the strategic utilization of human comparisons.
-The primary focus is on addressing the inherent limitations posed by noisy data and the complexities of high-dimensional models.
-By actively and effectively leveraging human comparisons, we aim to navigate the intricacies of model alignment with the minimal labeling effort.
-To connect the machine learning framework with human feedback, we propose a two-stage tractable framework through a probabilistic method.
-In the first stage, noisy labeled data is fed into the SFT oracle to obtain low-dimensional representations.
-More specifically, the algorithm can use lasso to select important features from high dimensions or use matrix completion to learn low-rank representations.
-The second stage employs human comparisons to improve model alignment within this low-dimensional space.
-Our framework, which utilizes the random utility model to describe human’s comparison behavior, can incorporate the possibility of false selection in human comparisons.
-Human may provide incorrect responses, though this noise tends to be substantially lower than that present in labeled data.
-To accommodate this aspect, we develop a probabilistic bisection approach for the alignment phase.
-We call this two-stage process as “SFT+HC” framework.
-The goal of our paper is to refine the model with a certain precision level and a confidence level using the minimum possible data.
-The rationale behind this framework design is rooted in the observation that some representation learning techniques, such as autoencoders or deep learning architectures (Zhou and Paffenroth 2017, Akrami et al.
-2019), are inherently robust to noisy labels (Li et al.
+Among various forms of human feedback, comparisons between alternatives have emerged as a particularly effective approach due to their cognitive simplicity.
+As Carl Jung observes, “Thinking is difficult, therefore let the herd pronounce judgment”—similarly, humans find it easier to select between alternatives than to generate absolute evaluations.
+Recent work has demonstrated the practical value of such comparative feedback: Zhou and Xu (2020) use pair-wise comparisons for story generation, while Xu et al.
+(2023) leverage human comparisons between large language models (LLMs) for fine-tuning.
+Although these applications show promising results, we lack a theoretical framework to understand when and why human comparisons enhance traditional supervised learning.
+This gap motivates our paper’s central question:
+How can we effectively combine supervised learning and human feedback to achieve model alignment?
+To make progress towards answering this question, we consider a two-stage probabilistic model that helps us analyze and understand the interplay between supervised learning and human feedback.
+In the first stage, noisy labeled data are fed into the SL oracle to obtain low-dimensional representations (or embeddings).
+For example, selecting a small set of features through Lasso is a way to represent high-dimensional feature spaces; similarly, low-rank matrix estimation is an example of representation of high-dimensional matrices.
+The second stage uses human comparisons to improve model alignment within this low-dimensional space.
+Our framework uses the random utility model to describe humans’ comparison behavior.
+Practically, humans may provide incorrect responses, so we develop a probabilistic bisection approach for the alignment phase.
+We call this two-stage process the Supervised Learning+Learning from Human Feedback (SL+LHF) model.
+The rationale behind this model is rooted in compelling evidence that representation learning techniques, such as deep learning architectures (Zhou and Paffenroth 2017, Akrami et al.
+2019), can effectively learn low-dimensional features even from highly noisy data (Li et al.
 2021, Taghanaki et al.
-2021).
-These methods aim to capture underlying patterns and structures in the data.
-Recently, Tsou et al.
-(2023) show that linear probes trained on the final embedding of pre-trained GPT-2 are surprisingly robust, offering competitive performance and, in some cases, even surpassing the results of full-model fine-tuning, where all weights are updated.
-In the second stage, we adopt human comparisons instead of human estimations because comparison is a much easier task.
-Our approach involves sequential design of comparisons and active data acquisition, with the objective of maximizing information and consequently minimizing sample complexity.
-To describe the effectiveness of the refinement phase through human comparison, we introduce a new concept called “label-noise-to-comparison-accuracy” (LNCA) ratio.
-We theoretically characterize the condition for the “SFT+HC” framework performing better than the pure SFT algorithm based on this LNCA ratio.
-In the large-noise scenario, human evaluators with high selection accuracy can significantly reduce the sample complexity.
+2021, Mousavi-Hosseini et al.
+2024).
+For instance, Tsou et al.
+(2023) recently showed that linear probes trained on the final embedding of pre-trained GPT-2 are surprisingly robust, sometimes even outperforming full-model fine-tuning.
+However, while supervised learning excels at learning these representations, the challenge lies in effectively fine-tuning the final layer—the stage where the model’s predictions are aligned with desired behaviors.
+For this alignment phase, we leverage human expertise through comparisons rather than direct labels, building on the cognitive advantage of comparative judgments discussed above.
+Our approach uses a sequential design of comparisons and active data acquisition to maximize information gain and minimize sample complexity, particularly effective when label noise is high.
+To describe the effectiveness of the refinement phase involving human comparison, we introduce a new concept called the “label-noise-to-comparison-accuracy” (LNCA) ratio.
+We offer a theoretical characterization of the condition in which the SL+LHF framework performs better than the pure SL algorithm based on this LNCA ratio.
+In a large-noise scenario, human evaluators with high selection accuracy can significantly reduce the sample complexity.
 1.1 Our Contributions
-Our contributions are summarized as follows:
-  1. (I)
-    Modeling: In a large-noise environment, machine learning models may exhibit fragility, due to the sensitivity to variations in input data and the dynamic nature of their surroundings.
-To address this challenge, we introduce a novel approach that leverages human comparisons to refine these models.
-To the best of our knowledge, our study represents the first theoretical attempt aimed at explaining the substantial enhancement of model performance under specific conditions through human comparisons.
-We present a novel two-stage framework based on a utility model for human comparisons.
-The model captures a natural phenomena that making distinctions between two options becomes increasingly challenging as they are more similar.
-Consequently, the accuracy of these comparisons may decrease as the model refines.
-  2. (II)
-    Theoretical Contributions:
-    1. (a)
-      Probabilistic bisection method: In the second phase, the decision-maker (DM) sequentially selects two options for human evaluators to compare.
-The fundamental question at this stage pertains to the efficient acquisition of valuable information from these human comparisons.
-To address this challenge, we introduce a probabilistic bisection method, which factors in the uncertainty arising from the accuracy of human comparisons.
-This method involves two types of actions: the ‘vertical move’ involves human evaluations of the same comparison.
-Once the algorithm identifies the superior choice with high probability, it proceeds with the ‘horizontal move,’ where the system selects the next pair for comparison.
-Our objective is to determine identify correct model parameters with a precision of $\epsilon$ and a confidence level of $1-\delta$.
-We establish the sample complexity required to achieve these precision and confidence levels.
-    2. (b)
-      Label-noise-to-comparison-accuracy ratio: We introduce a novel concept known as the ‘noise-to-human-knowledge ratio,’ which quantifies the relative scale between label noise and human-made comparison accuracy.
-In essence, when label noise is substantial, but humans can make selections with high accuracy, our two-stage framework outperforms pure Sequential Fine-Tuning (SFT).
-Specifically, when human selection accuracy deviates from 50%, the sample complexity for the refinement stage is $O(s\log(1/(\delta\varepsilon)))$.
-In cases where selection accuracy approaches 50% as the two options become indistinguishable from the true model, the sample complexity increases.
-We characterize the LNCA ratio conditions under which our SFT+HC framework outperforms pure SFT.
-  3. (III)
-    Numerical Performance: We evaluate the LNCA ratio through an empirical study on Amazon Mechanical Turk with a stylized task.
-The experiment result validates the condition under which “SFT+HC” performs better than pure SFT.
-Subsequently, we evaluated the performance of our proposed SFT+HC framework and compared it with the pure SFT approach across varying LNCA ratios.
-Our findings align with our theoretical results, revealing that the superiority of the SFT+HC framework becomes more pronounced under conditions of higher observational noise, higher accuracy of human comparisons, and lower dimensions of important features.
-1.2 Literature Review
+This paper makes contributions in four areas: theoretical modeling, characterization of sample complexity, practical implementation, and empirical validation.
+Modeling.
+We develop a theoretical framework that explains when and why human comparisons can substantially enhance performance of SL in high-noise environments.
+Our model captures the interaction between SL and human feedback through a utility-based comparative judgment approach.
+Theoretical contributions.
+Our main theoretical results consist of two components.
+First, we develop a probabilistic bisection method—an efficient sequential decision-making process that alternates between “vertical moves” (multiple evaluations of the same comparison) and “horizontal moves” (selection of new comparison pairs) to acquire human comparisons.
+Second, focusing on settings where the SL stage utilizes Lasso for learning low-dimensional representations, we introduce the label noise-to-comparison-accuracy ratio (LNCA) that quantifies the trade-off between label noise and human comparison accuracy.
+Our goal is to refine the model within an $\varepsilon$-distance with confidence at least $1-\delta$, which $\varepsilon$ represents the learning precision and $1-\delta$ denotes the confidence level.
+When human selection accuracy uniformly exceeds random guessing, we prove the refinement stage has sample complexity $O(s\log(1/(\delta\varepsilon)))$, where $s$ is the number of non-zero coefficients in Lasso.
+We characterize conditions under which our SL+LHF framework outperforms pure SL.
+Towards practical implementations.
+While our theoretical analysis examines human comparisons of model coefficients for analytical tractability, direct coefficient comparison is impractical in real applications.
+To bridge this gap, we develop an Active Sample Selection (ASS) framework that translates our theoretical insights into implementable procedures by enabling humans to compare meaningful data points rather than abstract model parameters.
+This framework provides a systematic approach for selecting the most informative samples for human comparison, applicable across a broad set of data distributions.
+Empirical analysis.
+Through Amazon Mechanical Turk (MTurk) experiments to acquire real human feedback on a stylized task, we validate our theoretical findings about the LNCA ratio and demonstrate the superior performance of SL+LHF over pure SL under conditions of higher observational noise, higher human comparison accuracy, and lower-dimensional feature spaces.
+The remainder of this paper is organized as follows: Section 2 reviews related literature.
+Section 3 presents our model setup, followed by Section 4, which introduces human comparison modeling and characterizes sample complexity in one dimension.
+Section 5 develops our two-stage framework integrating SL and human comparisons, with particular focus on sample complexity when using Lasso in the first stage.
+Section 6 details practical implementations for active sample selection, while Section 7 presents numerical experiments and managerial insights.
+Section 8 concludes.
+All proofs and parameter calibrations appear in the appendices.
+2 Literature Review
 Alignment and Learning from Human Feedback.
-The alignment challenge seeks to synchronize human values with machine learning systems and direct learning algorithms towards the objectives and interests of humans.
-Machine learning models frequently demonstrate unforeseen deficiencies in their performance when they are implemented in practical contexts, which was identified as underspecification (D’Amour et al.
+The alignment challenge seeks to synchronize human values with machine learning systems and direct learning algorithms toward the objectives and interests of humans.
+Machine learning models frequently demonstrate unforeseen deficiencies in their performance when they are implemented in practical contexts; this phenomenon has been identified as underspecification (D’Amour et al.
 2022).
-Predictors that perform similarly in the training dataset can exhibit significant variations when deployed in various domains.
+Predictors that perform similarly in the training dataset can exhibit significant variation when they are deployed in other domains.
 The presence of ambiguity in a model might result in instability and suboptimal performance in real-world scenarios.
@@
 2017, Ibarz et al.
-2018), fine-tuning language models to summarize text fine-tuning (Stiennon et al.
-2020), optimize dialogue (Jaques et al.
+2018) and fine-tuning language models to summarize text (Stiennon et al.
+2020) and to optimize dialogue (Jaques et al.
 2019).
-It has been demonstrated that using reinforcement learning from human feedback (RLHF) significantly boosts performance (Bai et al.
+Studies have demonstrated that using reinforcement learning from human feedback (RLHF) significantly boosts performance (Bai et al.
 2022, Glaese et al.
@@
 Ouyang et al.
-(2022) use reinforcement learning from human feedback (Stiennon et al.
+(2022) use RLHF (Stiennon et al.
 2020) to fine-tune GPT-3 to follow a broad class of written instructions.
 In particular, Zhou and Xu (2020) use pair-wise comparisons for story generation.
-They find that when comparing two natural language generation (NLG) models, instead of asking human annotator to assign scores separately for samples generated by different models, which resembles the case in the ADEM model (Lowe et al.
-2017), it is much easier for human annotators to directly compare one sample generated by the first model against another sample from the second model pairwisely and compute the win/loss rate.
+They find that, when comparing two natural language generation (NLG) models, asking a human annotator to assign scores separately for samples generated by different models, which resembles the case in the ADEM model (Lowe et al.
+2017), is more complicated.
+The much easier approach is for human annotators to directly compare one sample generated by the first model against another sample from the second model in a pair-wise fashion and then to compute the win/loss rate.
 Zhu et al.
-(2023) provide reinforcement learning framework with human feedback, specifically, for function approximation for $K$-wise comparisons with policy learning as the target.
+(2023) provide a reinforcement learning framework that includes human feedback – specifically, for function approximation in $K$-wise comparisons, with policy learning as the target.
 Xu et al.
 (2023) propose a fine-tuning algorithm by inducing a complementary distribution over the two sampled LLM responses to model human comparison processes, in accordance with the Bradley-Terry model for pairwise comparisons.
-Different than those above-mentioned work, we provide a theoretical two-stage framework for integrating supervised learning with actively acquired human comparison labels, and characterize the conditions where the two-stage framework outperforms pure SFT algorithms.
+In contrast to all these works, we provide a theoretical, two-stage framework for integrating supervised learning with actively acquired human comparison labels, and we characterize the conditions where the two-stage framework outperforms pure SL algorithms.
+Transfer Learning and Task Alignment.
+Model alignment generally refers to the process of ensuring that machine learning models are tailored to meet specific tasks and objectives.
+A related paradigm is transfer learning, where a model trained on one task is adapted for use in another task through additional fine-tuning (Pan and Yang 2009, Xu et al.
+2021, Bastani et al.
+2022, Du et al.
+2024).
+While our work shares some conceptual similarities with transfer learning, we focus specifically on the impact of noise in both supervised learning and human feedback phases.
+Our framework’s applicability to transfer learning scenarios depends crucially on the relationship between the initial and target tasks’ representations.
+When the target task shares the same low-dimensional representation as the initial task, requiring only updates to the final layer or coefficients, our theoretical results apply directly—the human comparison phase can effectively refine these parameters regardless of whether the comparison objective exactly matches the initial task.
+However, when the target task requires different representations, transfer learning typically demands an additional supervised learning phase with new labeled data to first refine these representations before any human-guided fine-tuning can begin.
+In such cases, our framework’s human comparison stage remains relevant for the final alignment when label noise is high, but it must be preceded by this representation learning phase.
 Active Learning.
 The core concept behind active learning in machine learning theory is the ability of a model to interactively query a human or an oracle to obtain labels for new data points.
-Instead of passively learning from a fixed dataset, an active learning algorithm intelligently selects the most informative instances to query for labels.
-We refer readers to Settles (2009) for a comprehensive review of the active learning algorithm.
-The literature has concentrated on label complexity to demonstrate the benefit of active learning, which refers to the number of labels requested to attain a specific accuracy.
+Instead of passively learning from a fixed dataset, an active learning algorithm intelligently selects the most informative learning instances to query for labels.
+We refer readers to Settles (2009) for a comprehensive review of the active learning algorithm literature.
+The literature has concentrated on studying label complexity to demonstrate the benefit of active learning, which refers to the number of labels requested to attain a specific accuracy.
 For example, Cohn et al.
-(1994) demonstrates that active learning can have a substantially lower label complexity than supervised learning in the noiseless binary classification issue.
+(1994) demonstrate that active learning can have a substantially lower label complexity than supervised learning in the noiseless binary classification issue.
 In the presence of noise, other papers focus on the classification problems (Hanneke 2007, Balcan et al.
-2006, Hanneke 2011) and regression problems (Sugiyama and Nakajima 2009, Beygelzimer et al.
-2009), etc.
-Though our study has a similar goal of using the fewest possible samples to achieve the target accuracy, we concentrate on the human-AI collaboration framework and try to reduce the sample complexity of human comparisons.
-Data-driven Decision Making and Human in the Loop.
-A large body of recent literature on operations research or operations management studied how to effectively utilize data to improve their decisions, including pricing, assortment, recommendation, etc.
-Motivated by the observation that human may sometimes have “private” information which an algorithm does not have access to that can improve performance, Balakrishnan et al.
-(2022) examined new theory that describes algorithm overriding behavior.
-They designed and experimentally tested feature transparency as an implementable approach that can help people better identify and use their private information.
+2006, Hanneke 2011) and the regression problems (Sugiyama and Nakajima 2009, Beygelzimer et al.
+2009).
+Although our study has a similar goal of using the fewest possible samples to achieve the target accuracy, we concentrate on the human-AI collaboration framework and try to reduce the sample complexity of human comparisons.
+Data-Driven Decision Making and Humans in the Loop.
+A large body of recent literature on operations research or operations management has studied how to effectively use data to improve decisions (Mišić and Perakis 2020), including pricing, assortment, and recommendation.
+Motivated by the observation that humans may sometimes have “private” information to which an algorithm does not have access but that can improve performance, Balakrishnan et al.
+(2022) examine new theory that describes algorithm overriding behavior.
+They designed an experiment to test feature transparency as an implementable approach that can help people better identify and use their private information.
 Ibrahim et al.
-(2021) studied how to address the shared information problem between humans and algorithms in a setting where the algorithm makes the final decision using the human’s prediction as input.
+(2021) study how to address the shared information problem between humans and algorithms in a setting where the algorithm makes the final decision using the human’s prediction as input.
 Dietvorst et al.
-(2018, 2015) study algorithm aversion phenomenon where people often fail to use algorithms after learning that they are imperfect, though evidence-based algorithms consistently outperform human forecasters.
-Other papers study Human-AI collaboration on a large variety of topics (Bastani et al.
+(2018, 2015) study the algorithm aversion phenomenon, in which people fail to use algorithms after learning that they are imperfect, even though evidence-based algorithms consistently outperform human forecasters.
+Other papers study human-artificial intelligence (AI) collaboration on a large variety of topics (Bastani et al.
 2021, Wu et al.
 2022, Deng et al.
-2020), including ride-hailing (Benjaafar et al.
+2020, Chen et al.
+2022, Caro and de Tejada Cuenca 2023, Ye et al.
+2024), including ride-hailing (Benjaafar et al.
 2022) and health care (Reverberi et al.
-2022), etc.
+2022).
 Probabilistic Bisection.
-The other related literature, which is independent of the above human-AI ML topics, is probabilistic bisection algorithm (Horstein 1963).
-Algorithms utilizing the PBA framework have been employed in several contexts, notably including the task of target localization (Tsiligkaridis et al.
+Another stream of literature related to our work, which is independent of the previously identified human-AI ML topics, is on probabilistic bisection algorithms (PBAs) (Horstein 1963).
+Algorithms using the PBA framework have been used in several contexts – notably, including the tasks of target localization (Tsiligkaridis et al.
 2014), scanning electron microscopy (Sznitman et al.
 2013), topology estimation (Castro and Nowak 2008), edge detection (Golubev and Levit 2003), and value function approximation for optimal stopping problems (Rodriguez and Ludkovski 2015).
-From the computational perspective, Frazier et al.
-(2019) has shown that probabilistic bisection converges almost as quickly as stochastic approximation.
-In this work, we design the model alignment algorithm by utilizing the probabilistic bisection approach, where the randomness is involved due to the inherent variability in the human judgement process.
-1.3 Organization of the Paper
-The remainder of this paper is organized as follows.
-Section 2 introduces the model setup.
-In Section 3, we introduce the model of human comparisons and characterize the sample complexity in a one-dimensional space.
-In Section 4, we present a two-stage framework which organically integrate SFT and human comparison.
-We specifically discuss the sample complexity when applying lasso in the first stage.
-Section 6 shows numerical experiments and managerial insights.
-Section 7 concludes the paper.
-Additional proofs and parameter calibrations are provided in the appendices.
+From a computational perspective, Frazier et al.
+(2019) have shown that probabilistic bisection converges almost as quickly as stochastic approximation.
+In this work, we design the model alignment algorithm by using the PBA approach, where randomness is involved because of the inherent variability in the human judgment process.
 3 Model
 In this section, we introduce our model setup.
-We first introduce the supervised learning oracle that will be potentially used during the initial learning stage, and then present the fine-tuning procedure through human comparisons during the refinement stage.
+We first introduce the supervised learning oracle that potentially is used during the initial learning stage, and then we present the fine-tuning procedure involving human feedback during the refinement stage.
+The two alternative ways of collecting human feedback are either to collect human labels directly or to ask humans to compare two answers.
+The former approach, which continues the label collecting procedure following the initial learning stage, is called pure supervised learning.
+As previously noted, we call the latter approach the Supervised Learning+Learning from Human Feedback framework.
 3.1 Supervised Learning and Underspecifications
-Given data $\{(X_{i},Y_{i})\}_{i\leq n}$ where $X_{i}\in\mathcal{X}$ and $Y_{i}\in\mathcal{Y}\subseteq\mathbb{R}$, the goal is to learn the true function that is parameterized by $\boldsymbol{\vartheta}^{*}\subseteq\mathbb{R}^{d}$, $h_{\boldsymbol{\vartheta}^{*}}\in\mathcal{H}:\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathcal{H}$ is the function class.
+Given the data $\{(X_{i},Y_{i})\}_{i\leq n}$, where $X_{i}\in\mathcal{X}$ and $Y_{i}\in\mathcal{Y}\subseteq\mathbb{R}$, the goal is to learn the true function that is parameterized by $\boldsymbol{\vartheta}^{*}\subseteq\mathbb{R}^{d}$, $h_{\boldsymbol{\vartheta}^{*}}\in\mathcal{H}:\mathcal{X}\rightarrow\mathcal{Y}$, where $\mathcal{H}$ is the function class.
 Each data point is generated from a stochastic response:
 $Y_{i}=h_{\boldsymbol{\vartheta}^{*}}(X_{i})+\epsilon_{i}\,,$
-where $\epsilon_{i}$ is the noise with mean 0 and variance $\sigma^{2}$, where $\sigma$ could potentially be very large.
-We specifically consider a high-dimensional problem that can be represented by low-dimensional manifolds; that is,
-$h_{\boldsymbol{\vartheta}^{*}}(x)=f_{\boldsymbol{\theta}^{*},\varphi(\cdot)}(x)\,,$
-where $f(\cdot)$ is assumed to be a generalized linear function in $\boldsymbol{\theta}^{*}$ and $\varphi(x)$.
-Here $\boldsymbol{\theta}^{*}\in\Theta\in\mathbb{R}^{s}$ is the unknown parameter [^2] We denote the parameter vector using the bold symbol., $\varphi(x)\in\mathbb{R}^{s}$ is the low-dimensional representation of feature $x$.
-It is assumed that $s\ll d$.
-For simplicity, we will remove the subscript $\varphi(\cdot)$ from $f$ when it is clear from the context.
-We first provide the following two examples as an illustration.
+where $\epsilon_{i}$ is the noise with mean 0 and variance $\sigma^{2}$, and where $\sigma$ could potentially be very large.
+We specifically consider a high-dimensional problem that can be represented by the low-dimensional generalized linear model; that is,
+$h_{\boldsymbol{\vartheta}^{*}}(x)=f(\left\langle\boldsymbol{\theta}^{*},\varphi(x)\right\rangle)\,,$
+where $f(\cdot)$ is a link function.
+Here, $\boldsymbol{\theta}^{*}\in\Theta\in\mathbb{R}^{s}$ is the unknown parameter [^1] We denote the parameter vector using the bold symbol.; and $\varphi(x)\in\mathbb{R}^{s}$ is the kernel function, which is the low-dimensional representation of feature $x$.
+Our goal is to learn both the parameter $\boldsymbol{\theta}^{*}$ and representation $\varphi(x)$.
+We assume that $s\ll d$.
+We provide the following two examples as illustrations.
 Example 3.1 (Sparse Linear Models)
-The parameter $\boldsymbol{\vartheta}^{*}$ has nonzero entries in the first $s$ dimensions and has zero entries in the rest of dimensions.
-In the linear model, $f_{\boldsymbol{\vartheta}^{*}}(x)=\sum_{i=1}^{s}\vartheta^{*}_{i}x_{i}$.
-In this case, $\boldsymbol{\theta}^{*}=(\vartheta^{*}_{1},\cdots,\vartheta^{*}_{s})$, $\varphi(x)=(x_{1},\cdots,x_{s})$, and $f(\cdot)$ is a linear function.
+Suppose $x$, $\boldsymbol{\vartheta}^{*}\in\mathbb{R}^{d}$, and $h_{\boldsymbol{\vartheta}^{*}}(x)=\left\langle\boldsymbol{\vartheta}^{*},x\right\rangle$.
+The parameter $\boldsymbol{\vartheta}^{*}$ has nonzero entries in the first $s$ dimensions and has zero entries in the rest of the dimensions.
+In this case, $h_{\boldsymbol{\vartheta}^{*}}(x)=f(\left\langle\boldsymbol{\theta}^{*},\varphi(x)\right\rangle)$ where $\boldsymbol{\theta}^{*}=(\vartheta^{*}_{1},\cdots,\vartheta^{*}_{s})$, $\varphi(x)=(x_{1},\cdots,x_{s})$, and $f(\cdot)$ is an identity function.
 Example 3.2 (Generalized Low-rank Models)
-In the generalized low-rank models, it is assumed that
-$f_{\boldsymbol{\Theta}^{*}}(X)=\mu(\left\langle\boldsymbol{\Theta}^{*},xz^{\top}\right\rangle)=\mu(x^{\top}\boldsymbol{\Theta}^{*}z)$
-where $\mu(\cdot)$ is the link function, $\boldsymbol{\Theta}^{*}\in\mathbb{R}^{d\times d}$ is a low-rank matrix and $X=xz^{\top}$, i.e., $\boldsymbol{\Theta}^{*}=\sum_{i=1}^{s}\sigma_{i}u_{i}v_{i}^{\top}$ where $u_{i}$ and $v_{i}\in\mathbb{R}^{d}$.
-Thus, $f_{\boldsymbol{\Theta}^{*}}(X)=\sum_{i=1}^{s}\sigma_{i}(x^{\top}u_{i})(v_{i}^{\top}z)$.
-In this case, $\varphi(X)$ is the projection of $X$ onto the low-rank space and $\boldsymbol{\theta}^{*}$ is the $s$-dimensional vector of singular values.
-Suppose $h_{\boldsymbol{\vartheta}}$ can be estimated through a statistical loss function $\ell:\mathcal{Y}\times\mathbb{R}\rightarrow\mathbb{R}$, and the total loss is
+Suppose $x\in\mathbb{R}^{d\times d}$.
+In the generalized low-rank models, we assume that $h_{\boldsymbol{\vartheta}^{*}}(x)=f(\left\langle\boldsymbol{\Theta}^{*},x\right\rangle)$ where $\boldsymbol{\Theta}^{*}\in\mathbb{R}^{d\times d}$ is a low-rank matrix (i.e., $\boldsymbol{\Theta}^{*}=\sum_{i=1}^{s}\sigma_{i}u_{i}v_{i}^{\top}$ where $u_{i}$ and $v_{i}\in\mathbb{R}^{d})$; the trace inner product is defined as $\left\langle\boldsymbol{\Theta}^{*},x\right\rangle=tr(\boldsymbol{\Theta}^{*}x^{\top})$.
+Thus,
+$h_{\boldsymbol{\vartheta}^{*}}(x)=f(\sum_{i=1}^{s}\sigma_{i}\cdot tr(u_{i}v_{i}^{\top}x^{\top}))=f(\left\langle\boldsymbol{\theta}^{*},\varphi(x)\right\rangle),$
+where $\varphi(X)=[tr(u_{i}v_{i}^{\top}x^{\top})]_{i=1}^{s}$ is the projection of $x$ onto the low-rank space, and $\boldsymbol{\theta}^{*}=(\sigma_{1},\cdots,\sigma_{s})$ is the $s$-dimensional vector of singular values.
+Suppose that $h_{\boldsymbol{\vartheta}}$ can be estimated through a statistical loss function $\ell:\mathcal{Y}\times\mathbb{R}\rightarrow\mathbb{R}$ and that the total loss is
 $L_{n}(h_{\boldsymbol{\vartheta}})=\sum_{i=1}^{n}\ell(Y_{i},h_{\boldsymbol{\vartheta}}(X_{i})).$
-For example, $\ell$ can represent the squared loss, the regularized squared loss, among others.
-To deal with the high-dimensional issue, the algorithm learns the low-dimensional representation $\varphi(x)$ as well as the parameter $\boldsymbol{\theta}^{*}$.
-To learn the sparse linear models for example, one can apply lasso to select important features and learn the model parameters.
+For example, $\ell$ can represent the squared loss and the regularized squared loss, among others.
+To deal with the high-dimensional issue, the algorithm learns the low-dimensional representation $\varphi(x)$, as well as the parameter $\boldsymbol{\theta}^{*}$.
+For example, to learn the sparse linear models, one can apply Lasso to select important features and to learn the model parameters.
 To learn the low-rank models, one can use trace regression to learn the low-dimensional representations.
 Define $h_{\hat{\boldsymbol{\vartheta}}_{n}}$ as the function contained in $\mathcal{H}$ that minimizes empirical risk $L_{n}$.
-In the existing literature, it is very common that the decision-maker directly employs the predictor $h_{\hat{\boldsymbol{\vartheta}}_{n}}$.
+In the existing literature, the decision maker very commonly employs the predictor $h_{\hat{\boldsymbol{\vartheta}}_{n}}$ directly.
 However, some issues may arise.
-In high-dimensional problems where significant noise is present, approaches striving to balance the variance-bias trade-off may occasionally diverge from intuitive expectations.
+In high-dimensional problems, where significant noise is present, approaches striving to balance the variance–bias trade-off may violate fundamental principles, such as the natural expectation of negative demand elasticity in economics.
 We use the following example to illustrate this potential issue.
 Example 3.3
-Consider a linear pricing model with high-dimensional contexts, i.e., $Y_{i}=\vartheta_{1}\cdot p_{i}+\vartheta_{2}^{\top}X_{i}+\epsilon_{i}$ where $Y_{i}$ is the demand, $p_{i}$ is the price, and the vector $X_{i}$ represents features of product $i$, and also assume that $\vartheta_{1}<0$.
-However, in a high-noise regime we may incorrectly estimate a positive value for $\vartheta_{1}$.
-Please see Appendix 9.1 for numerical illustrations.
-In Example 2.3, the positive correlation between price and demand is misinterpreted due to model underspecification.
+Consider a linear pricing model with high-dimensional contexts – that is, $Y_{i}=\vartheta_{1}\cdot p_{i}+\vartheta_{2}^{\top}X_{i}+\epsilon_{i}$, where $Y_{i}$ is the demand, $p_{i}$ is the price, and the vector $X_{i}$ represents features of product $i$.
+Also, assume that $\vartheta_{1}<0$.
+However, in a high-noise regime, we may incorrectly estimate a positive value for $\vartheta_{1}$.
+Please see Appendix 10 for numerical illustrations.
+In Example 3.3, the negative coefficient of the price in the demand function is misinterpreted because of model underspecification.
 The requirement of $\vartheta_{1}<0$ should be specified based on human knowledge.
-In the face of the challenge posed by underspecification within a potentially very noisy environment, we take a critical step towards refinement, which we term “fine-tuning through human comparisons.” Human can do comparisons with a much higher precision.
-For example, based on the human knowledge, it can be immediately corrected that the demand and the price is negatively correlated.
-This particular stage comes into play after the output generated by our machine learning pipelines, represented as $\hat{\varphi}(\cdot)$ and $\hat{\boldsymbol{\theta}}_{0}$ (which denotes the estimator of $\varphi$ and $\boldsymbol{\theta}$), has been produced.
-During this stage, we focus on enhancing the model within a lower-dimensional space.
-We move on to the refinement step after the first stage where the kernel function, which provides a low-dimensional representation, has been effectively trained on the noisy-labeled data.
+In the face of the challenge that underspecification poses in a potentially very noisy environment, we take a critical step toward refinement; we call this step “model alignment through human comparisons.” Humans can do comparisons with much greater precision.
+For example, on the basis of their knowledge, humans can immediately correct the negative correlation between demand and price.
+This particular stage comes into play after our machine learning pipelines generate and produce their output $\hat{\varphi}(\cdot)$ and $\hat{\boldsymbol{\theta}}_{0}$ (which denotes the estimator of $\varphi$ and $\boldsymbol{\theta}$) by training on the noisy-labeled data.
+During the second stage, we refine the model in a lower dimensional space.
+Throughout the paper, we focus our analysis on sparse linear models.
 3.2 Utility Model of Human Comparisons
 In this section, we introduce the human comparison model.
-Given two potential models $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$, the system asks human to compare and select the better prediction model.
-If $f_{\boldsymbol{\theta}_{1}}$ is more preferable than $f_{\boldsymbol{\theta}_{2}}$, we denote it as $f_{\boldsymbol{\theta}_{1}}\succeq f_{\boldsymbol{\theta}_{2}}$.
+Given two potential models, $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$, the system asks humans to compare and select the better prediction model; we call this step the model-level comparison.
+If $f_{\boldsymbol{\theta}_{1}}$ is preferable to $f_{\boldsymbol{\theta}_{2}}$, we denote the result as $f_{\boldsymbol{\theta}_{1}}\succeq f_{\boldsymbol{\theta}_{2}}$.
+Remark 3.4 (Model-level Comparison and Sample-level Comparison)
+In practice, the comparison between two models can be conducted in various ways.
+For example, the models can be evaluated on a single data point (either from the humans’ past experience or their observation), denoted as $\{x,y\}$.
+When comparing two models in this case, the one with a smaller squared loss is preferred; that is:
+$\mathbb{P}(\boldsymbol{\theta}_{1}\succeq\boldsymbol{\theta}_{2})$ $=\mathbb{P}\left((y-f_{\boldsymbol{\theta}_{1}}(x))^{2}\leq(y-f_{\boldsymbol{\theta}_{2}}(x))^{2}\right).$
+Alternatively, models also can be assessed on a batch of samples.
+We call this approach the sample-level comparison.
+Note that a model-level comparison is a broader concept.
+We elaborate on how to translate the model-level comparison to the sample-level comparison in Section 6.
+In this section, our focus is on a framework for undertaking a more general model-level comparison.
 The preference over the model $f_{\boldsymbol{\theta}}$ is captured by a utility function $u(\boldsymbol{\theta})$, where $u(\boldsymbol{\theta})$ depicts the distance between the model parameterized by $\boldsymbol{\theta}$ and the true model.
-That is, we assume $u(\boldsymbol{\theta})=-d(\boldsymbol{\theta},\boldsymbol{\theta}^{*})$ where $d(\cdot,\cdot)$ is a distance function.
-The larger value of utility indicates a better model.
+That is, we assume $u(\boldsymbol{\theta})=-d(\boldsymbol{\theta},\boldsymbol{\theta}^{*})$, where $d(\cdot,\cdot)$ is a distance function.
+The larger utility value indicates the better model.
 The true model parameter $\boldsymbol{\theta}^{*}$ achieves the highest utility.
-For example, it can be the negative two-norm distance, i.e., $u(\boldsymbol{\theta})=-\|\boldsymbol{\theta}-\boldsymbol{\theta}^{*}\|_{2}$.
+For example, it can be the negative two-norm distance – that is, $u(\boldsymbol{\theta})=-\|\boldsymbol{\theta}-\boldsymbol{\theta}^{*}\|_{2}$.
 We formally state the assumption about the utility function as follows:
@@
 The utility function is unimodal in each coordinate.
-That is, for each coordinate $\theta_{i}$, when fixing all other coordinates, $u(\boldsymbol{\theta})$ is monotonically increasing when $\theta_{i}<\theta_{i}^{*}$ and monotonically decreasing when $\theta_{i}>\theta_{i}^{*}$.
+That is, for each coordinate $\theta_{i}$, when fixing all other coordinates, $u(\boldsymbol{\theta})$ is monotonically increasing when $\theta_{i}<\theta_{i}^{*}$ and is monotonically decreasing when $\theta_{i}>\theta_{i}^{*}$.
 Moreover, the utility gets lower when $\boldsymbol{\theta}$ gets far away from the true parameter.
@@
 The comparison between $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$ provides information on which side is closer to the true parameter $\boldsymbol{\theta}^{*}$.
-Define the hyperplane $H(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})=\{\boldsymbol{\theta}:d(\boldsymbol{\theta},\boldsymbol{\theta}_{1})=d(\boldsymbol{\theta},\boldsymbol{\theta}_{2})\}$, and half-space $H^{-}(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})=\{\boldsymbol{\theta}:d(\boldsymbol{\theta},\boldsymbol{\theta}_{1})<d(\boldsymbol{\theta},\boldsymbol{\theta}_{2})\}$ and $H^{+}(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})=\{\boldsymbol{\theta}:d(\boldsymbol{\theta},\boldsymbol{\theta}_{1})>d(\boldsymbol{\theta},\boldsymbol{\theta}_{2})\}$.
+Define the hyperplane $H(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})=\{\boldsymbol{\theta}:d(\boldsymbol{\theta},\boldsymbol{\theta}_{1})=d(\boldsymbol{\theta},\boldsymbol{\theta}_{2})\}$, as well as half-space $H^{-}(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})=\{\boldsymbol{\theta}:d(\boldsymbol{\theta},\boldsymbol{\theta}_{1})<d(\boldsymbol{\theta},\boldsymbol{\theta}_{2})\}$ and half-space $H^{+}(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})=\{\boldsymbol{\theta}:d(\boldsymbol{\theta},\boldsymbol{\theta}_{1})>d(\boldsymbol{\theta},\boldsymbol{\theta}_{2})\}$.
 The condition $u(\boldsymbol{\theta}_{1})\geq u(\boldsymbol{\theta}_{2})$ implies that $\boldsymbol{\theta}^{*}\in H^{-}(\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2})$.
 Assumption 3.4 implies that $H(\theta_{1},\theta_{2})$ contains a unique point when $\theta_{1}$ and $\theta_{2}$ are in a one-dimensional space.
+The overall goal is to find the $\boldsymbol{\theta}$ that achieves near the highest utility or that, equivalently, gets close to the true model parameter $\boldsymbol{\theta}^{*}$ within a certain precision.
+Suppose that the human selection process follows a random utility model (RUM).
+In the process of querying, the human’s opinion regarding $f_{\boldsymbol{\theta}}$ is
+$U(\boldsymbol{\theta})=u(\boldsymbol{\theta})+\xi,$ (3.1)
+where $\xi$ is the unobserved noise.
+We assume that the noise $\xi$ follows a Gumbel distribution with parameter $\gamma$, which represents the expert level of the human.
+We say human evaluators make the correct choice if they select $\boldsymbol{\theta}_{1}$ ($\boldsymbol{\theta}_{2}$) when $u(\boldsymbol{\theta}_{1})$ ($u(\boldsymbol{\theta}_{2})$) is higher than $u(\boldsymbol{\theta}_{2})$ ($u(\boldsymbol{\theta}_{1})$).
+A smaller value of $\gamma$ corresponds to a higher certainty of the selection’s correctness.
+When $\gamma$ approaches 0, humans can select the better one with almost 100% accuracy.
 Remark 3.6
-In practice, the comparison between two models can be conducted in various ways.
-For example, it can be evaluated on a single point with high confidence (either from human’s past experience or their observation), denoted as $\{x,y\}$.
-When comparing two models, the one with smaller square loss is more preferred, i.e.,
-$\mathbb{P}(\boldsymbol{\theta}_{1}\succeq\boldsymbol{\theta}_{2})$ $=\mathbb{P}\left((y-f_{\boldsymbol{\theta}_{1}}(x))^{2}\leq(y-f_{\boldsymbol{\theta}_{2}}(x))^{2}\right).$
-We will elaborate this process in Section 5.
-The overall goal is to find $\boldsymbol{\theta}$ that achieves near the highest utility, or equivalently, gets close to the true model parameter $\boldsymbol{\theta}^{*}$ within a certain precision.
-Suppose that human’s selection process follows a random utility model (RUM).
-In the process of querying, human’s opinion regarding $f_{\boldsymbol{\theta}}$ is
-$U(\boldsymbol{\theta})=u(\boldsymbol{\theta})+\beta_{h}+\xi,$
-where $\beta_{h}$ represents the human bias, and $\xi$ is the unobserved noise.
-The human bias $\beta_{h}$ can vary between individuals.
-We assume the noise $\xi$ follows Gumbel distribution with parameter $\gamma$, which represents the expert level of the human.
-We say the human makes the correct choice if he selects $\boldsymbol{\theta}_{1}$ ($\boldsymbol{\theta}_{2}$) when $u(\boldsymbol{\theta}_{1})$ ($u(\boldsymbol{\theta}_{2})$) is higher than $u(\boldsymbol{\theta}_{2})$ ($u(\boldsymbol{\theta}_{1})$).
-A smaller value of $\gamma$ corresponds to a higher certainty of the selection correctness.
-When $\gamma$ approaches 0, human can always select the better one with almost 100% accuracy.
+Our model can also incorporate the human bias in the following way: $U(\boldsymbol{\theta})=u(\boldsymbol{\theta})+\beta_{h}+\xi$, where the human bias $\beta_{h}$ can vary between individuals.
+For simplicity, we focus on the utility model (3.1).
 Lemma 3.7 (Precision)
-For any two models $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$, the probability that human will make the right selection is
+For any two models $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$, the probability that a human will make the right selection is
 $\frac{1}{1+\exp(-|u(\boldsymbol{\theta}_{2})-u(\boldsymbol{\theta}_{1})|/\gamma)}\,,$
-which is strictly larger than 1/2.
-Lemma 2.6 implies that under RUM, the selection accuracy is always higher than 50%, even with the nonzero human bias.
+which is strictly greater than 1/2.
+Lemma 3.7 implies that under a RUM, the selection accuracy is always higher than 50%, even with the nonzero human bias.
 Intuitively, the differentiation would become harder when $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$ both get closer to the true parameter.
-However, how fast it converges to 1/2 when $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$ approach $\boldsymbol{\theta}^{*}$ depend on the distance function.
-Under this probabilistic selection model, we will design the algorithm and investigate the sample complexity.
+However, how fast it converges to 1/2 when $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$ approach $\boldsymbol{\theta}^{*}$ depends on the distance function.
+Under this probabilistic selection model, we design the algorithm and investigate the sample complexity.
 The parameter $\boldsymbol{\theta}$ may lie in a multi-dimensional space.
-To illustrate our algorithm, we first describe the one-dimensional human comparison procedure, and then extend to the multi-dimensional scenario.
+To illustrate our algorithm, we first describe the one-dimensional human comparison procedure, and then we extend to the multi-dimensional scenario.
 4 One-dimensional Human Comparison
-Once the low-dimensional representation $\hat{\varphi}$ has been established in the initial stage, as detailed in Section 4, we introduce a framework for model alignment through human comparisons.
-The parameter $\boldsymbol{\theta}$ may lie in a multi-dimensional space.
-To illustrate our algorithm, we first describe the one-dimensional human comparison procedure, and then extend to the multi-dimensional scenario.
-The primary challenge in designing the framework lies in strategically selecting pairwise comparisons to minimize sample complexity, which bears resemblance to the principles of active learning.
-In this section, we first discuss the deterministic selection (with 100% accuracy) and then study the probabilistic selection.
-For the probabilistic selection, we propose an algorithm named Alignment through Probabilistic Bisection (APB) and characterize its sample complexity.
+The low-dimensional representation $\hat{\varphi}$ is established in the initial stage, as we discuss in Section 5.
+Following this stage, our framework calls for model alignment through human comparisons.
+At this point, the parameter $\boldsymbol{\theta}$ may lie in a multi-dimensional space.
+To illustrate our algorithm, we first describe the one-dimensional human comparison procedure and then extend it to the multi-dimensional scenario.
+The primary challenge in designing the framework lies in strategically selecting pairwise comparisons to minimize the number of comparisons (or sample complexity), which resembles the principles of active learning.
+First, we consider a special case in which humans can select the better model with 100% accuracy (deterministic selection).
+Second, we relax this to a probabilistic selection.
+For the probabilistic selection, we propose an algorithm we call Model Alignment through Probabilistic Bisection (MAPB) and characterize its sample complexity.
 4.1 Deterministic Bisection
@@
 In the one-dimensional case, suppose $\theta^{*}\in\Theta\subseteq\mathbb{R}$ and $|\theta|\leq\beta_{\Theta}$ for all $\theta\in\Theta$.
-The overall goal is to use the least samples to find some $\theta$ such that $d(\theta,\theta^{*})\leq\varepsilon$.
-When human is able to select a better model with 100% accuracy, the deterministic bisection algorithm can be utilized to approach the true parameter.
-The algorithm operates by maintaining a search interval interval $[\theta^{-},\theta^{+}]$ that contains the true parameter $\theta^{*}$.
-The interval is initialized by $[-\beta_{\Theta},\beta_{\Theta}]$, and at round $k$ the center of this interval with respect to the distance $d(\cdot,\cdot)$ is used as query point.
-Specifically, defining the query point $\theta_{k}$ to be the single point in the set $H(\theta^{-},\theta^{+})$, we ask the human to compare utility of two points, denoted by $c_{\Delta}^{-}(\theta_{k})$ and $c_{\Delta}^{+}(\theta_{k})$, that have equal distance $\Delta/2$ from $\theta_{k}$ in opposite directions, where $\Delta$ is called comparison granularity and is a tuning parameter.
-For example, for the Euclidean distance, at each round $\theta_{k}=(\theta^{-}+\theta^{+})/2$, and $c_{\Delta}^{-}(\theta_{k})=\theta_{k}-\Delta/2$ and $c_{\Delta}^{+}(\theta_{k})=\theta_{k}+\Delta/2$.
-If the human selects $c_{\Delta}^{-}(\theta_{k})$ then we eliminate the upper half of the search interval by updating $\theta^{+}$ to be equal to $\theta_{k}$; otherwise, we eliminate the bottom half of the search interval by setting $\theta^{-}$ to be equal to $\theta_{k}$.
-After this bisection step, we continue the process by selecting $\theta_{k+1}$ to be $H(\theta^{-},\theta^{+})$ again, and repeat the above till the length of the search interval is below the precision $\varepsilon$.
-The pseudo code for this is presented in Algorithm 1.
+The overall goal is to use the least number of samples to find a $\theta$ such that $d(\theta,\theta^{*})\leq\varepsilon$.
+When a human is able to select a better model with 100% accuracy, the deterministic bisection algorithm can be used to approach the true parameter.
+The algorithm operates by maintaining a search interval $[\theta^{-},\theta^{+}]$ that contains the true parameter $\theta^{*}$.
+The interval is initialized by $[-\beta_{\Theta},\beta_{\Theta}]$, and at round $k$, the center of this interval with respect to the distance $d(\cdot,\cdot)$ is used as the query point.
+Specifically, after defining the query point $\theta_{k}$ to be the single point in the set $H(\theta^{-},\theta^{+})$, we ask the human to compare the utility of two points, denoted by $c_{\Delta}^{-}(\theta_{k})$ and $c_{\Delta}^{+}(\theta_{k})$.
+These two points have equal distance $\Delta/2$ from $\theta_{k}$ in opposite directions, where $\Delta$ is called comparison granularity and is a tuning parameter.
+For example, for the Euclidean distance, at each round, $\theta_{k}=(\theta^{-}+\theta^{+})/2$; also, $c_{\Delta}^{-}(\theta_{k})=\theta_{k}-\Delta/2$, and $c_{\Delta}^{+}(\theta_{k})=\theta_{k}+\Delta/2$.
+If the human selects $c_{\Delta}^{-}(\theta_{k})$, then we eliminate the upper half of the search interval by updating $\theta^{+}$ to be equal to $\theta_{k}$; otherwise, we eliminate the bottom half of the search interval by setting $\theta^{-}$ to be equal to $\theta_{k}$.
+After this bisection step, we continue the process by selecting $\theta_{k+1}$ to be $H(\theta^{-},\theta^{+})$ again, and repeat the above step until the length of the search interval falls below the precision $\varepsilon$.
+The pseudo-code is presented in Algorithm 1.
 $\theta_{k}$$\theta^{*}$$\theta^{-}$$\theta^{+}$$c_{\Delta}^{+}(\theta_{k})$$c_{\Delta}^{-}(\theta_{k})$keepeliminate
 Figure 1: An illustrative example.
-Among two choices $c_{\Delta}^{-}(\theta_{k})$ and $c_{\Delta}^{+}(\theta_{k})$, $c_{\Delta}^{-}(\theta_{k})$ would be selected since it is closer to the true parameter.
+Among two choices $c_{\Delta}^{-}(\theta_{k})$ and $c_{\Delta}^{+}(\theta_{k})$, $c_{\Delta}^{-}(\theta_{k})$ would be selected because it is closer to the true parameter.
 In this case, the interval $(\theta_{k},\theta^{+}]$ is eliminated.
@@
 9:until $d(\theta^{+},\theta^{-})\leq\varepsilon$.
+We can show the following result for the performance of Algorithm DB.
 Proposition 4.1
-Suppose human makes the selection with 100% accuracy $(\gamma\to 0)$.
+Suppose the human makes the selection with 100% accuracy $(\gamma\to 0)$.
 Fix $\varepsilon>0$.
 After $O(\log_{2}\left(\frac{\beta_{\Theta}}{\varepsilon}\right))$ rounds, Algorithm 1 reaches a point $\theta_{k}$ such that $d(\theta_{k},\theta^{*})\leq\varepsilon$.
-However, the deterministic model of human selection described above exhibits two significant limitations: firstly, in practical scenarios, humans may not always be able to identify the correct choice, implying that $\gamma>0$; secondly, the accuracy of human selections is influenced by the comparison granularity $\Delta$.
-To address these concerns, we introduce a more general probabilistic model in the following subsection.
+Note that the comparison granularity has no impact on Algorithm 1 and its computational complexity because the deterministic model assumes 100% accuracy for any $\Delta$.
+However, this deterministic model of human selection as described exhibits two significant limitations: First, in practical scenarios, humans may not always be able to identify the correct choice, implying that $\gamma>0$; second, the accuracy of human selections is influenced by the comparison granularity $\Delta$.
+To address these limitations, we introduce a more general probabilistic model in the following section.
 4.2 Probabilistic Bisection
-First, we enrich the model to capture the probabilistic nature of human comparisons by noting that, in the RUM, when $\gamma$ is strictly positive, human’s selection is random with certain distribution that favors the correct answer, parameterized by $\gamma$.
-Precisely, Lemma 2.6 indicates that human will select $c_{\Delta}^{-}(\theta)$ with probability
-$p^{-}_{\gamma}(\theta):=\left[1+\exp\left\{\frac{u(c_{\Delta}^{+}(\theta))-u(c_{\Delta}^{-}(\theta))}{\gamma}\right\}\right]^{-1}\,,$
-and selects $c_{\Delta}^{+}(\theta)$ with probability $p^{+}_{\gamma}(\theta)$ equal to $1-p^{-}_{\gamma}(\theta)$.
-We note that, in the above equation, the selection probability depends on the comparison granularity.
-This dependence, naturally captures difficulty of the selection for the human.
-Specifically, when $\Delta$ is too small, $c_{\Delta}^{-}(\theta)$ and $c_{\Delta}^{+}(\theta)$ converge to each other, hence $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$ converge to $1/2$, representing the difficulty of distinguishing between too highly similar choices.
-On the other hand, when $\Delta$ is too large, one expects the selection to become difficult as well as both options are far away from $\theta^{*}$.
-This can be captured for example when the distance function converges to a finite value for far away points from $\theta^{*}$ as in Example 2.4.
+First, we enrich the model to capture the probabilistic nature of human comparisons by noting that, in the RUM, when $\gamma$ is strictly positive, the human’s selection is random, with a certain distribution that favors the correct answer, parameterized by $\gamma$.
+Precisely, as shown in Lemma 3.7, the human selects $c_{\Delta}^{-}(\theta)$ with probability
+$p^{-}_{\gamma}(\theta):=\left[1+\exp\left\{\frac{u\Big{(}c_{\Delta}^{+}(\theta)\Big{)}-u\Big{(}c_{\Delta}^{-}(\theta)\Big{)}}{\gamma}\right\}\right]^{-1}\,,$
+and the human selects $c_{\Delta}^{+}(\theta)$ with probability $p^{+}_{\gamma}(\theta)$, which is equal to $1-p^{-}_{\gamma}(\theta)$.
+We note that, in this equation, the selection probability depends on the comparison granularity.
+This dependence naturally captures the difficulty of the selection for the human.
+On the one hand, when $\Delta$ is too small, $c_{\Delta}^{-}(\theta)$ and $c_{\Delta}^{+}(\theta)$ converge toward each other; hence, $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$ converge to $1/2$, representing the difficulty of distinguishing between two highly similar choices.
+On the other hand, when $\Delta$ is too large, the selection can become difficult as well because both options are far away from $\theta^{*}$.
+To illustrate, this difficulty can be captured when the distance function converges to a finite value for far away points from $\theta^{*}$, as in Example 3.5.
 In this case, $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$ could also converge to 1/2 (see Example 4.2).
-In the remaining of the paper, for brevity, we do not include $\Delta$ in the notations $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$.
+As a result of this phenomenon, and for brevity, we exclude $\Delta$ in the notations $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$ in the remainder of the paper.
 Example 4.2
-Consider $u(\theta)=-\frac{1}{1+\exp(-|\theta|)}$ with the highest utility achieved at $\theta=0$.
-When $\Delta$ is too large, both $u(c_{\Delta}^{+}(\theta))$ and $u(c_{\Delta}^{-}(\theta))$ are close to -1 even $\theta$ is far away from 0, and thus $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$ are both close to 1/2.
+Consider
+$u(\theta)=-\frac{1}{1+\exp(-|\theta|)}\,.$
+Here, the highest utility is achieved at $\theta=0$.
+For any finite $\theta$, when $\Delta$ is too large, both $u(c_{\Delta}^{+}(\theta))$ and $u(c_{\Delta}^{-}(\theta))$ are close to -1, and thus $p^{-}_{\gamma}(\theta)$ and $p^{+}_{\gamma}(\theta)$ are both close to 1/2.
 Definition 4.3 ($(\varepsilon,\delta)$-alignment problem)
-Let $\varepsilon>0$ denote the precision level and $\delta>0$ represent the confidence level.
+Let $\varepsilon>0$ denote the precision level and let $\delta>0$ represent the confidence level.
 The $(\varepsilon,\delta)$-alignment problem aims to identify a $\theta$ such that $\mathbb{P}(d(\theta,\theta^{*})\leq\varepsilon)\geq 1-\delta$.
 Our objective is to solve the $(\varepsilon,\delta)$-alignment problem with small sample complexity.
-The algorithm design is inspired by the probabilistic bisection developed in Frazier et al.
+By selecting appropriate values for these parameters, we can also control other metrics such as risk.
+The algorithm design is inspired by the probabilistic bisection developed by Frazier et al.
 (2019).
-While their work primarily focuses on the analysis of convergence rates, our emphasis is on conducting a sample complexity analysis for an algorithm designed to terminate when both the precision level and confidence level criteria are met.
-If $\theta<\theta^{*}$, then $p_{\gamma}^{-}(\theta)<1/2$ and $p_{\gamma}^{+}(\theta)>1/2$; vice versa.
-We introduce some notation for simplicity.
+Although their work primarily focuses on the analysis of convergence rates, our emphasis is on conducting a sample complexity analysis for an algorithm designed to terminate when both the precision level and confidence level criteria are met.
+If $\theta<\theta^{*}$, then $p_{\gamma}^{-}(\theta)<1/2$ and $p_{\gamma}^{+}(\theta)>1/2$; and vice versa.
+We introduce several notations to enhance simplicity.
 Let $Y(\theta)$ denote the selection: $Y(\theta)=c_{\Delta}^{-}(\theta)$ with probability $p_{\gamma}^{-}(\theta)$ and $Y(\theta)=c_{\Delta}^{+}(\theta)$ with probability $p_{\gamma}^{+}(\theta)$.
-Define $\tilde{Z}(\theta)=2\cdot\boldsymbol{1}(Y(\theta)=c_{\Delta}^{+}(\theta))-1$, so that $\tilde{Z}(\theta)=1$ if $Y(\theta)=c_{\Delta}^{+}(\theta)$ (so that we believe $\theta^{*}$ more likely to be to the right of $\theta$) and $\tilde{Z}(\theta)=-1$ if $Y(\theta)=c_{\Delta}^{-}(\theta)$ (so that we believe $\theta^{*}$ more likely to be to the left of $\theta$).
+Define $\tilde{Z}(\theta)=2\cdot\boldsymbol{1}(Y(\theta)=c_{\Delta}^{+}(\theta))-1$, so that $\tilde{Z}(\theta)=1$ if $Y(\theta)=c_{\Delta}^{+}(\theta)$ (when we believe $\theta^{*}$ is more likely to be to the right of $\theta$) and $\tilde{Z}(\theta)=-1$ if $Y(\theta)=c_{\Delta}^{-}(\theta)$ (when we believe $\theta^{*}$ is more likely to be to the left of $\theta$).
 Note that we have $\tilde{Z}(\theta)=1$ with probability $p_{\gamma}^{+}(\theta)$.
 Also, define
-$\tilde{p}(\theta)=\begin{cases}\mathbb{P}(\tilde{Z}(\theta)=1),&\mbox{if }\theta\leq\theta^{*}\\ \mathbb{P}(\tilde{Z}(\theta)=-1),&\mbox{if }\theta>\theta^{*}\end{cases}\,,$
-so that $\tilde{p}(\theta)$ is the probability that $\tilde{Z}(\theta)$ correctly indicates the direction of the root $\theta^{*}$ from $\theta$.
+$\tilde{p}(\theta)=\mathbb{P}\big{(}\tilde{Z}(\theta)\text{ correctly identifies the direction}\big{)}=\begin{cases}\mathbb{P}(\tilde{Z}(\theta)=1),&\mbox{if }\theta\leq\theta^{*}\\ \mathbb{P}(\tilde{Z}(\theta)=-1),&\mbox{if }\theta>\theta^{*}\end{cases}\,.$
 Figure 2 shows the structure of introducing our main algorithm.
-The general idea of the Algorithm APB is described as follows.
-There are two types of moves: the horizontal moves and the vertical moves.
-For any fixed $\theta$, we ask human to repeatedly evaluate $c_{\Delta}^{-}(\theta)$ and $c_{\Delta}^{+}(\theta)$.
-We refer this as the vertical moves.
-Once the vertical stopping criteria is satisfied, the algorithm decides the next $\theta$ to evaluate according to the updated distribution.
-We call this step as the horizontal moves.
-The horizontal moves continue until the horizontal stopping criteria is satisfied.
-In the following two sections, we discuss horizontal moves and vertical steps, as well as the stopping criteria, in sequence.
-Then we introduce the two-stage Human-AI collaboration framework for multiple dimensions in Section 4 and practical implementations in Section 5.
-One-dimensional APB (Section 3.2) Horizontal moves (Section 3.2.2); Complexity characterization (Section 3.3) Sample complexity (Theorem 3.5) Two-stage framework for Human-AI collaboration (Section 4.1); complexity characterization (Section 4.3) Practical implementations (Section 5) Vertical moves (Section 3.2.1); Complexity characterization (Section 3.4)
-Figure 2: Roadmap for introducing 2.
+Algorithm MAPB, described as follows, involves two types of moves: the horizontal moves and the vertical moves.
+For any fixed $\theta$, we ask the human to repeatedly evaluate $c_{\Delta}^{-}(\theta)$ and $c_{\Delta}^{+}(\theta)$.
+We refer to this as the vertical moves.
+After the vertical stopping criteria are satisfied, the algorithm decides the next $\theta$ to evaluate according to a certain distribution that will be defined.
+We call this step the horizontal move.
+The horizontal moves continue until the horizontal stopping criteria are satisfied.
+In the following two sections, we discuss the vertical moves and horizontal moves, as well as their stopping criteria, in sequence.
+Then, as shown in Figure 2, we introduce the two-stage Human-AI collaboration framework for multiple dimensions in Section 5 and practical implementations in Section 6.
+One-dimensional MAPB (Section 4.2) Horizontal moves (Section 4.2.2); Complexity characterization (Section 4.3) Sample complexity (Theorem 4.5) Extension to multi-dimensional setting (Section 5); Two-stage framework for human-AI collaboration (Section 5.1); Complexity characterization (Section 5.3) Practical implementations (Section 6) Vertical moves (Section 4.2.1); Complexity characterization (Section 4.4)
+Figure 2: Roadmap for introducing Algorithm 2.
 4.2.1 Vertical moves
 Fix some $\theta$.
-Repeatedly, we ask human evaluations of the query point $\theta$, achieved through comparisons of $c^{+}_{\Delta}(\theta)$ and $c^{-}_{\Delta}(\theta)$.
+We repeatedly ask for human evaluations of the query point $\theta$, achieved through comparisons of $c^{+}_{\Delta}(\theta)$ and $c^{-}_{\Delta}(\theta)$.
 We observe a sequence of signs $\{\tilde{Z}_{i}(\theta)\}_{i}$.
 If $\theta<\theta^{*}$, then the expectation $\mathbb{E}[\tilde{Z}_{i}(\theta)]=2\tilde{p}(\theta)-1>0$; if $\theta>\theta^{*}$, then the expectation $\mathbb{E}[\tilde{Z}_{i}(\theta)]=1-2\tilde{p}(\theta)<0$.
-From the response $\tilde{Z}_{i}(\theta)$, we define the simple random walk $S(\theta)=\{S_{m}(\theta):m\geq 0\}$ with $S_{0}(\theta)=0$ and $S_{m}(\theta)=\sum_{i=1}^{m}\tilde{Z}_{i}(\theta)$ for $m\geq 1$.
+From the response $\tilde{Z}_{i}(\theta)$, we define the simple random walk $\{S_{m}(\theta):m\geq 0\}$ with $S_{0}(\theta)=0$ and $S_{m}(\theta)=\sum_{i=1}^{m}\tilde{Z}_{i}(\theta)$ for $m\geq 1$.
 Sequential test of power one indicates whether the drift $\omega$ of a random walk satisfies the hypothesis $\omega<0$ versus $\omega>0$.
-Define the stopping time $\tau^{\uparrow}_{1}(\theta)=\inf\{m\in\mathbb{N}:|S_{m}(\theta)|\geq\hbar_{m}\}$ where $\{\hbar_{m}\}_{m}$ is a positive sequence.
-Here $\uparrow$” ($\rightarrow$) represents “vertical” (horizontal, respectively).
-The test decides that $\theta<\theta^{*}$ if $S_{\tau_{1}^{\uparrow}(\theta)}(\theta)\geq\hbar_{\tau^{\uparrow}_{1}(\theta)}$, that $\theta>\theta^{*}$ if $S_{\tau_{1}^{\uparrow}(\theta)}(\theta)\leq-\hbar_{\tau^{\uparrow}_{1}(\theta)}$ and the test does not make a decision if $\tau^{\uparrow}_{1}(\theta)=\infty$.
-Assume now that the algorithm queries at some point $\theta_{k}\neq\theta^{*}$ at the $(k+1)^{th}$ iteration.
-We then observe a random walk with $m^{th}$ term $S_{k,m}=S_{k,m}(\theta_{k})=\sum_{i=1}^{m}Z_{k,i}(\theta_{k})$ until we reach the end of the power-one test.
-Define the new signal
+Define the stopping time $\tau^{\uparrow}_{1}(\theta)=\inf\{m\in\mathbb{N}:|S_{m}(\theta)|\geq\hbar_{m}\}$, where $\{\hbar_{m}\}_{m}$ is a positive sequence.
+Here, $\uparrow$” and $\rightarrow$” represent vertical and horizontal, respectively.
+The test decides that $\theta<\theta^{*}$, if $S_{\tau_{1}^{\uparrow}(\theta)}(\theta)\geq\hbar_{\tau^{\uparrow}_{1}(\theta)}$; it decides that $\theta>\theta^{*}$ if $S_{\tau_{1}^{\uparrow}(\theta)}(\theta)\leq-\hbar_{\tau^{\uparrow}_{1}(\theta)}$; and the test does not make a decision if $\tau^{\uparrow}_{1}(\theta)=\infty$.
+Assume now that the algorithm queries a point $\theta_{k}$ (not equal to $\theta^{*}$) at the $(k+1)^{th}$ iteration.
+We then observe a random walk with the $m^{th}$ term $S_{k,m}=S_{k,m}(\theta_{k})=\sum_{i=1}^{m}Z_{k,i}(\theta_{k})$ until we reach the end of the test, which is called the power-one test (Robbins and Siegmund 1974).
+Define the new signal:
 $Z_{k}(\theta_{k})=\begin{cases}+1,&\mbox{if }S_{k,\tau_{1}^{\uparrow}(\theta_{k})}>0\,,\\ -1,&\mbox{if }S_{k,\tau_{1}^{\uparrow}(\theta_{k})}<0\,.\end{cases}$
@@
 Define $\hbar_{m}=(2m(\ln(m+1)-\ln\eta))^{1/2}$.
-On the event $\theta_{k}>\theta^{*}$, $\mathbb{P}(Z_{k}(\theta_{k})=+1|\theta_{k})\leq\eta/2$.
-On the event $\theta_{k}<\theta^{*}$, $\mathbb{P}(Z_{k}(\theta_{k})=-1|\theta_{k})\leq\eta/2$.
+In the event that $\theta_{k}>\theta^{*}$, $\mathbb{P}(Z_{k}(\theta_{k})=+1|\theta_{k})\leq\eta/2$.
+In the event that $\theta_{k}<\theta^{*}$, $\mathbb{P}(Z_{k}(\theta_{k})=-1|\theta_{k})\leq\eta/2$.
 That is, the detection accuracy $p(\theta)$ when the vertical move stops is bounded by $p(\theta)\geq 1-\eta/2$.
-Lemma 3.4 states that the detection accuracy is lower bounded by $p_{c}:=1-\eta/2$ where the accuracy can be chosen through the selection of $\eta$.
-As $\theta$ approaches $\theta^{*}$, the differentiation between choices becomes increasingly challenging because the drift tends to diminish, potentially requiring more steps in the testing process.
-However, it is crucial to remember that our objective is to locate a $\theta$ such that $|\theta-\theta^{*}|\leq\varepsilon$ with a confidence level of $1-\delta$.
+Lemma 4.4 states that the detection accuracy is lower bounded by $p_{c}:=1-\eta/2$, where the accuracy can be chosen through the selection of $\eta$.
+As $\theta$ approaches $\theta^{*}$, differentiating between choices becomes increasingly challenging because the drift tends to diminish, potentially requiring more steps in the testing process.
+However, our objective, you must recall, is to locate a $\theta$ such that $|\theta-\theta^{*}|\leq\varepsilon$, with a confidence level of $1-\delta$.
 In the vicinity of $\theta^{*}$, waiting until $|S_{m}(\theta)|\geq\hbar_{m}$ to terminate the test may not be necessary.
-A high volume of queries at a particular point $\theta_{k}$ suggests close proximity between $\theta_{k}$ and $\theta$, making it advisable to consider terminating the test sooner.
-Based on this insight, we introduce the stopping time $\tau^{\uparrow}(\delta,\varepsilon)$ that is determined by the precision and confidence level.
-This stopping time regulates the maximum number of queries carried out at a single point as $\theta$ approaches the local neighborhood of the true parameter (a concept to be elaborated on in Section 3.4).
-Specifically, as $\theta$ converges towards $\theta^{*}$ within a defined local neighborhood, denoted as $\{\theta^{\prime}:d(\theta,\theta^{\prime})\leq a\}$ for some positive constant $a$, surpassing the threshold $\tau^{\uparrow}(\delta,\varepsilon)$ signifies the difficulty in distinguishing between the parameters due to their proximity.
-The algorithm terminates.
+A high volume of queries at a particular point $\theta_{k}$ suggests close proximity between $\theta_{k}$ and $\theta$, so that considering a earlier termination of the test may be advisable.
+Based on this insight, we introduce a stopping time $\tau^{\uparrow}(\delta,\varepsilon)$, which is determined by the precision and confidence levels.
+This stopping time regulates the maximum number of queries carried out at a single point as $\theta$ approaches the local neighborhood of the true parameter (a concept we elaborate on in Section 4.4).
+Specifically, $\theta$ converges toward $\theta^{*}$ in a defined local neighborhood, denoted as $\{\theta^{\prime}:d(\theta,\theta^{\prime})\leq a\}$ for some positive constant $a$; here, surpassing the threshold $\tau^{\uparrow}(\delta,\varepsilon)$ signifies the difficulty in distinguishing between the parameters because of their proximity, and the algorithm terminates.
 4.2.2 Horizontal moves
@@
 For example, $\rho_{0}$ can be chosen as a uniform distribution.
-Choose some $p\in(1/2,p_{c})$ and let $q=1-p$.
+Choose some $p\in(1/2,p_{c})$, and let $q=1-p$.
 When the algorithm queries at the point $\theta_{k}$ to obtain $Z_{k}$, we update the density $\rho_{k}$:
@@
 The next query point $\theta_{k+1}$ is determined as the median of $\rho_{k+1}$.
-The probability density in the correct side would be increased while the probability density in the wrong side would be reduced.
-Define $\tau^{\rightarrow}(\delta,a)$ as the number of horizontal moves to reach $a$-neighborhood of $\theta^{*}$ with confidence level $1-\delta$ (which will be specified later).
-That is, parameter $\theta$ moves to the local $a$-neighborhood of $\theta^{*}$ with probability at least $21-\delta/$ 2 when the number of steps exceeds $\tau^{\rightarrow}(\delta/2,a)$.
-The horizontal move stops when the total number of movement reaches $\tau^{\rightarrow}(\delta/2,\varepsilon)$ since the precision level $\varepsilon$ is achieved.
+The probability density on the correct side would be increased while the probability density on the wrong side would be reduced.
+Define $\tau^{\rightarrow}(\delta,a)$ as the number of horizontal moves to reach $a$-neighborhood of $\theta^{*}$ with confidence level $1-\delta$ (which will be specifically defined in the paper later).
+That is, parameter $\theta$ moves to the local $a$-neighborhood of $\theta^{*}$ with probability of at least $21-\delta/$ 2 when the number of steps exceeds $\tau^{\rightarrow}(\delta/2,a)$.
+The horizontal move stops when the total number of moves reaches $\tau^{\rightarrow}(\delta/2,\varepsilon)$ because at this point the precision level $\varepsilon$ has been achieved.
 In summary, there are two scenarios:
   1. (I)
-    If the horizontal move is less than $\tau^{\rightarrow}(\delta/2,a)$ (indicating it may not yet reach the local neighborhood of $\theta^{*}$ within a distance of $a$), the vertical test stops when it reaches $\tau_{1}^{\uparrow}(\theta)$.
-Then it moves to the next query point $\theta$;
+    If the horizontal move is less than $\tau^{\rightarrow}(\delta/2,a)$ (indicating it may not yet reach the local neighborhood of $\theta^{*}$ within a distance of $a$), the vertical test stops when the movement reaches $\tau_{1}^{\uparrow}(\theta)$.
+Then it moves to the next query point $\theta$.
   2. (II)
-    If the horizontal move is more than $\tau^{\rightarrow}(\delta/2,a)$ (signifying entry into the local neighborhood of $\theta^{*}$ within a distance of $a$):
+    If the horizontal move is more than $\tau^{\rightarrow}(\delta/2,a)$, signifying entry into the local neighborhood of $\theta^{*}$ within a distance of $a$, there are two possible options:
     1. (a)
-      If the vertical move reaches $\tau^{\uparrow}(\delta/2,\varepsilon)$ first, the algorithm terminates, having identified a $\theta$ within the $\varepsilon$ range of $\theta^{*}$;
+      When the vertical move first reaches $\tau^{\uparrow}(\delta/2,\varepsilon)$, the algorithm terminates, having identified a $\theta$ within the $\varepsilon$ range of $\theta^{*}$.
     2. (b)
-      Alternatively, if the vertical movement reaches $\tau_{1}^{\uparrow}(\theta)$ first, the vertical test stops and then transitions to the next query point $\theta$.
-The rationale behind distinguishing between local and non-local scenarios lies in the observation that two distinct situations can result in a significant number of vertical moves: when both compared models are performing well (i.e., $\theta_{1}$ and $\theta_{2}$ are close to $\theta^{*}$), or when both models are performing poorly (i.e., $\theta_{1}$ and $\theta_{2}$ are far from $\theta^{*}$).
-In the latter scenario, it is imperative that the algorithm does not terminate prematurely, as $\theta$ has not yet approached $\theta^{*}$ sufficiently.
+      When the vertical move first reaches $\tau_{1}^{\uparrow}(\theta)$, the vertical test stops and then transitions to the next query point $\theta$.
+The rationale behind distinguishing between local and non-local scenarios lies in the observation that two distinct situations can result in a significant number of vertical moves: (1) when both compared models are performing well (i.e., $\theta_{1}$ and $\theta_{2}$ are close to $\theta^{*}$), or (2) when both models are performing poorly (i.e., $\theta_{1}$ and $\theta_{2}$ are far from $\theta^{*}$).
+In the latter case, that the algorithm not terminate prematurely is crucial because $\theta$ has not yet gotten sufficiently close to $\theta^{*}$.
 Therefore, the stopping criterion $\tau^{\uparrow}(\delta,\varepsilon)$ is only applied when $\theta$ transitions into the local neighborhood of $\theta^{*}$.
@@
 Without any historical data, $\mu_{0}$ can be initialized as the uniform distribution over the confidence region.
-Alternatively, $\mu_{0}$ can be approximated by the empirical distribution of the estimator from the bootstrap method applying to the first phase.
+Alternatively, $\mu_{0}$ can be approximated by the empirical distribution of the estimator using the bootstrap method and applying it to the first phase.
 At step $k$, the algorithm tests $\theta_{k}$ by collecting $Z_{k,s}(\theta_{k})$.
@@
 Otherwise, the vertical move stops when $s$ is larger than $\tau_{1}^{\uparrow}(\theta_{k})$.
-Once the test on $\theta_{k}$ stops, we update prior distribution $\rho_{k}$ and select the next $\theta_{k+1}$ as the median of the updated distribution.
-Algorithm 2 APB (Alignment through Probabilistic Bisection)
+Once the test on $\theta_{k}$ stops, we update the prior distribution $\rho_{k}$ and select the next $\theta_{k+1}$ as the median of the updated distribution.
+Algorithm 2 MAPB (Model Alignment through Probabilistic Bisection)
 1:input: confidence level $\delta$, precision $\varepsilon$; initial distribution $\mu_{0}$; local parameter $a>0$; $k=0$; $\eta$;
@@
 13:until horizontal step $k$ satisfies stopping criteria $\tau^{\rightarrow}(\delta,\varepsilon)$
+14:pick $\theta_{k}$ as the median of $\mu_{k}$
 4.3 Complexity of Horizontal Moves
-Let $T^{\uparrow}(\theta)$ denote the number of comparisons that are collected at $\theta$ and $T^{\rightarrow}$ denote the number of horizontal moves; let $T_{k}^{+}$ denote the total number of comparisons that are collected after $k$ horizontal-moves, where “$+$” denotes both horizontal- and vertical-moves.
+Let $T^{\uparrow}(\theta)$ denote the number of comparisons that are collected at $\theta$ and let $T^{\rightarrow}$ denote the number of horizontal moves; let $T_{k}^{+}$ denote the total number of comparisons that are collected after $k$ horizontal moves, where “$+$” denotes both horizontal and vertical moves.
 Note that $T^{\uparrow}(\theta)$, $T^{\rightarrow}$, and $T_{k}^{+}$ are all random variables.
-In what follows, we discuss the complexity of vertical and horizontal moves, and characterize its sample complexity.
+In the following paragraphs, we discuss the complexity of vertical and horizontal moves, and we characterize its sample complexity.
 We first characterize the complexity of horizontal moves.
 Fix some $a\in(0,\theta^{*}+\beta_{\Theta})$.
-Define three regions: $A=[-\beta_{\Theta},\theta^{*}-a]$, $B=(\theta^{*}-a,\theta^{*}]$, $C=(\theta^{*},\beta_{\Theta}]$.
+Define three regions: $A=[-\beta_{\Theta},\theta^{*}-a]$, $B=(\theta^{*}-a,\theta^{*}]$, and $C=(\theta^{*},\beta_{\Theta}]$.
 Define $T(a)=\inf\{k^{\prime}\geq 0:\theta_{k}\in B\cup C,\forall k\geq k^{\prime}\}$ to be the time required until the sequence of medians never reenters $A$.
 Let $\nu_{k}$ denote the (random) measure corresponding to the conditional posterior distribution, conditional on lying to the left of $\theta^{*}$, so that for any measurable $D$, $\nu_{k}(D)=\mu_{k}(D\cap[-\beta_{\Theta},\theta^{*}])/\mu_{k}([-\beta_{\Theta},\theta^{*}])$.
-Next, fix $\Delta\in(0,1/2)$ and define the stopping time $\tau(a)$ to be the first time that the conditional mass in $B$ lies above $1-\Delta$ and the median $\theta_{k}$ lies to the left of $\theta^{*}$, i.e.,
+Next, fix $\Delta\in(0,1/2)$ and define the stopping time $\tau(a)$ to be the first time that the conditional mass in $B$ lies above $1-\Delta$ and the median $\theta_{k}$ lies to the left of $\theta^{*}$; that is,
 $\tau(a)=\inf\{k\geq 0:\nu_{k}(B)>1-\Delta,\theta_{k}<\theta^{*}\}.$
@@
 Algorithm 2 starts with an initial distribution $\mu_{0}$.
-Without any prior information, $\mu_{0}$ can be set as a uniform distribution; otherwise we can use the prior information to update $\mu_{0}$.
-We first make the following mild assumption regarding the initial prior distribution.
+Without any prior information, $\mu_{0}$ can be set as a uniform distribution; otherwise, we can use the prior information to update $\mu_{0}$.
+First, we make the following mild assumption regarding the initial prior distribution.
 {assumption}
 [Prior distribution] The initial prior distribution satisfies that $\mu_{0}(B)\geq a^{\varsigma}$ for some $\varsigma\leq 1$.
-The uniform distribution satisfies Assumption 3.5 since $\mu_{0}(B)=\mu_{0}((\theta^{*}-a,\theta^{*}])=a$.
+The uniform distribution satisfies Assumption 4.5 because $\mu_{0}(B)=\mu_{0}((\theta^{*}-a,\theta^{*}])=a$.
 With additional information, the prior distribution would be more centered around the true parameter $\theta^{*}$.
@@
 for some $r_{1},r_{2},\alpha>0$ and any $\omega>0$.
-Lemma 3.6 has shown the light tail distribution of $\tau(\cdot)$.
+Lemma 4.6 shows the light tail distribution of $\tau(\cdot)$.
 For small $\omega$, note that $\omega^{\varsigma}$ decreases in $\varsigma$.
-That is, when the initial prior distribution is more centered around the true parameter, $\varsigma$ is smaller and thus the right-hand-side tail distribution is lighter.
+That is, when the initial prior distribution is centered more closely around the true parameter, $\varsigma$ is smaller; thus, the right-hand-side tail distribution is lighter.
 We next characterize the complexity of horizontal moves.
@@
 $\mathbb{P}(|\theta_{k}-\theta^{*}|\leq\varepsilon)>1-\delta,\quad\forall k\geq\tau^{\rightarrow}(\delta,\varepsilon)\,.$
-Theorem 3.7 states that after $\tau^{\rightarrow}(\delta,\varepsilon)$ horizontal moves, $\theta_{k}$ reaches $\varepsilon$-neighborhood of the true parameter with probability at least $1-\delta$.
+Theorem 4.7 states that after $\tau^{\rightarrow}(\delta,\varepsilon)$ horizontal moves, $\theta_{k}$ reaches the $\varepsilon$-neighborhood of the true parameter with probability at least $1-\delta$.
 The complexity of this process depends on $\log(1/\delta)$, $\log(1/\varepsilon)$, and $\varsigma(1+\varsigma)$.
-When the initial prior distribution is more centered around the true parameter $\theta^{*}$, fewer horizontal moves are needed.
+When the initial prior distribution is more closely centered around the true parameter $\theta^{*}$, fewer horizontal moves are needed.
 Define the local region of $\theta^{*}$ as $[\theta^{*}-a,\theta^{*}+a]$.
-We divide the horizontal moves to two phases: Phase I includes all steps until the sequence of medians never reenters $[-\beta_{\Theta},\theta^{*}-a]\cup[\theta^{*}+a,\beta_{\Theta}]$; Phase II includes all steps after Phase I.
-Similar to the definition of $T(a)$, we define $T^{\prime}(a)$ as the stopping time that the medians never reenters $A^{\prime}=[\theta^{*}+a,\beta_{\Theta}]$.
-Let $\psi(a)$ be the time required until the sequence of medians never reenters $A\cup A^{\prime}$, i.e., $\psi(a)=\max\{T(a),T^{\prime}(a)\}$.
+We divide the horizontal moves into two phases: Phase I includes all steps until the sequence of the medians never reenters $[-\beta_{\Theta},\theta^{*}-a]\cup[\theta^{*}+a,\beta_{\Theta}]$; and Phase II includes all steps after Phase I.
+Similar to the definition of $T(a)$, we define $T^{\prime}(a)$ as the stopping time, which is when the medians never reenter $A^{\prime}=[\theta^{*}+a,\beta_{\Theta}]$.
+Let $\psi(a)$ be the time required until the sequence of the medians never reenters $A\cup A^{\prime}$ – that is, $\psi(a)=\max\{T(a),T^{\prime}(a)\}$.
 Proposition 4.8 provides an upper bound for the expectation of moves outside of $[\theta^{*}-a,\theta^{*}+a]$.
@@
 The expectation of $\psi(a)$ is bounded by
-$\mathbb{E}[\psi(a)]\leq\frac{4\theta^{*}}{r_{2}\alpha a}+\frac{4\beta_{1}}{r_{1}}+\frac{2\beta_{2}}{r_{R}}.$
+$\mathbb{E}[\psi(a)]\leq\frac{4\theta^{*}}{r_{2}\alpha a^{\varsigma}}+\frac{4\beta_{1}}{r_{1}}+\frac{2\beta_{2}}{r_{R}}.$
 4.4 Complexity of Vertical Moves
-The vertical move has two stopping criteria, depending on whether $\theta_{k}$ has moved to the local neighborhood of $\theta^{*}$.
-If $\theta_{k}$ is still outside of the local area, the algorithm would finish the power-one test; otherwise, the algorithm may terminate earlier before stopping the power-one test.
-Lemma 3.9 first characterizes the expected number of queries needed for testing $\theta$.
+The vertical move has two stopping criteria that depend on whether $\theta_{k}$ has moved to the local neighborhood of $\theta^{*}$.
+If $\theta_{k}$ is still outside of the local area, the algorithm would finish with the power-one test; otherwise, the algorithm may terminate earlier, before stopping the power-one test.
+Lemma 4.9 characterizes the expected number of queries needed for testing $\theta$.
 Lemma 4.9 (Power-one test)
@@
 When $\tilde{p}(\theta)$ is bounded away from $1/2$, the expected vertical move is also bounded.
-Lemma 2.6 shows that $\tilde{p}(\theta)$ is strictly larger than 1/2 for any $\theta\neq\theta^{*}$, then according to Heine–Borel theorem, there exists $\delta^{\prime}>0$ such that $\tilde{p}(\theta)>1/2+\delta^{\prime}$ for all $\theta\in[-\beta_{\Theta},\theta^{*}-a]\cup[\theta^{*}+a,-\beta_{\Theta}]$ (see the proof of Lemma 8.9 in Appendix 8.1).
-For the ease of notation, we define the bound of vertical moves
+As a result of Lemma 3.7, $\tilde{p}(\theta)$ is strictly larger than 1/2 for any $\theta\neq\theta^{*}$; thus, according to Heine–Borel theorem, there exists $\delta^{\prime}>0$ such that $\tilde{p}(\theta)>1/2+\delta^{\prime}$ for all $\theta\in[-\beta_{\Theta},\theta^{*}-a]\cup[\theta^{*}+a,-\beta_{\Theta}]$.
+(See the proof of Lemma 9.9 in Appendix 9.1.) For the ease of notation, we define the bound of vertical moves:
 $\phi(\tilde{p})=c_{1}|2\tilde{p}-1|^{-2}\ln(|2\tilde{p}-1|^{-1})+c_{2}.$
-The selection accuracy bounded away from $1/2$ implies that $\max_{a\leq\|\theta-\theta^{*}\|\leq\beta_{\Theta}}\phi(\tilde{p}(\theta))$ is bounded.
+The selection accuracy’s being bounded away from $1/2$ implies that $\max_{a\leq\|\theta-\theta^{*}\|\leq\beta_{\Theta}}\phi(\tilde{p}(\theta))$ also is bounded.
 Thus, we can infer the expected number of steps—comprising both horizontal and vertical movements—outside of $[\theta^{*}-a,\theta^{*}+a]$ as follows.
@@
 The expected steps outside of $[\theta^{*}-a,\theta^{*}+a]$ is bounded by
-$\mathbb{E}\left[\sum_{k=1}^{\psi(a)}T^{\uparrow}(\theta_{k})\boldsymbol{1}(\|\theta_{k}-\theta^{*}\|\geq a)\right]\leq\left(\frac{4\theta^{*}}{r_{2}\alpha a}+\frac{4\beta_{1}}{r_{1}}+\frac{2\beta_{2}}{r_{R}}\right)\max_{a\leq\|\theta-\theta^{*}\|\leq\beta_{\Theta}}\phi(\tilde{p}(\theta))\,.$
+$\mathbb{E}\left[\sum_{k=1}^{\psi(a)}T^{\uparrow}(\theta_{k})\boldsymbol{1}(\|\theta_{k}-\theta^{*}\|\geq a)\right]\leq\left(\frac{4\theta^{*}}{r_{2}\alpha a^{\varsigma}}+\frac{4\beta_{1}}{r_{1}}+\frac{2\beta_{2}}{r_{R}}\right)\max_{a\leq\|\theta-\theta^{*}\|\leq\beta_{\Theta}}\phi(\tilde{p}(\theta))\,.$
 Proposition 4.10 highlights that the sample complexity of Phase I hinges on the choice of $a$ and the span of $\theta$ from the prior distribution.
 If $\beta_{\Theta}$ is overly expansive, there is a high probability that both selected parameters $\theta_{1}$ and $\theta_{2}$ are distant from the true parameter.
-Consequently, the accuracy of human selection may diminish considerably, resembling random selection, as illustrated in Example 3.2.
+Consequently, the accuracy of human selection may diminish considerably – and may even resemble random selection, as illustrated in Example 4.2.
 Thus, by narrowing the range of $\theta$ based on supplementary information, the sample complexity of Phase I can be notably diminished.
 Once Phase I ends, $\theta_{k}$ moves to the $a$-neighborhood of $\theta^{*}$.
-In scenarios where both choices are close to $\theta^{*}$, discerning the superior parameter becomes more challenging.
-Depending on whether the selection accuracy converges towards $1/2$ as the detected parameter approaches the true parameter, we classify the problem into two cases: 1) The selection accuracy is bounded away from 1/2 for all $\theta\neq\theta^{*}$; 2) the section accuracy approaches $1/2$.
+In scenarios where both options are close to $\theta^{*}$, discerning the superior parameter becomes more challenging.
+Depending on whether the selection accuracy converges toward $1/2$ as the detected parameter approaches the true parameter, we classify the problem into two cases: 1) The selection accuracy is bounded away from 1/2 for all $\theta\neq\theta^{*}$; and 2) the section accuracy approaches $1/2$.
 The second case is harder and requires more delicate analysis.
-We first characterize the sample complexity for case 1) and then analyze case 2).
+We first characterize the sample complexity for Case 1 and then analyze Case 2.
 For the scenario where $\tilde{p}$ is bounded away from $\underline{p}>1/2$, combining the bound on the vertical moves (Proposition 4.10) with the complexity of the horizontal moves (Theorem 4.7), Theorem 4.11 characterizes the total sample complexity.
 Theorem 4.11 (Sample complexity when $\underline{p}$ exists)
-If $\tilde{p}(\theta)\geq\underline{p}>1/2$ for all $\theta\in\Theta$, then the expected samples collected before algorithm termination is
+If $\tilde{p}(\theta)\geq\underline{p}>1/2$ for all $\theta\in\Theta$, then the expected number of samples collected before algorithm termination is
 $\mathbb{E}[T^{+}_{T^{\rightarrow}}]\leq\tau^{\rightarrow}(\delta,\varepsilon)\cdot\phi(\underline{p})=O(\log(1/(\delta\varepsilon))).$
-Next, we address the more challenging scenario in which the selection accuracy tends towards $1/2$ as the tested parameter approaches the true parameter.
+Next, we address the more challenging scenario, in which the selection accuracy tends toward $1/2$ as the tested parameter approaches the true parameter.
 4.4.1 When the selection accuracy approaches $1/2$.
-As the comparison accuracy approaches $1/2$, the vertical move (power-one test) requires collecting more samples.
+As the comparison accuracy approaches $1/2$, the vertical move (based on the power-one test) requires the collection of more samples.
 The rate at which the vertical sample complexity increases depends on the rate at which the utility changes as $\theta$ converges to $\theta^{*}$.
@@
 The parameter $\kappa$ characterizes the speed of convergence, with a smaller value indicating a faster convergence rate.
-It is natural to assume $\kappa\leq 1$ for several reasons.
-Firstly, human decision-making should not be inferior to learning from a random sample; otherwise, collecting random samples would suffice.
+Assuming that $\kappa\leq 1$ is natural for several reasons.
+First, human decision making should not be inferior to learning from a random sample; otherwise, collecting random samples would suffice.
 We demonstrate this by examining the selection accuracy of learning from a single sample.
 Suppose the sample $(x,y)$ is generated from $y=\theta^{*}x+\epsilon$ where $\epsilon\sim\mathcal{N}(0,\sigma^{2})$.
-When comparing $\theta+\Delta$ and $\theta-\Delta$ (without loss of generality, assuming $\theta<\theta^{*}$), $\theta+\Delta$ is more preferable than $\theta-\Delta$ if the residual is smaller.
-That is, the probability of $\theta+\Delta$ being better is
+When comparing $\theta+\Delta$ and $\theta-\Delta$ (without loss of generality, assuming $\theta<\theta^{*}$), $\theta+\Delta$ is preferable to $\theta-\Delta$ if the residual is smaller.
+That is, the probability that $\theta+\Delta$ is better is
 $\mathbb{P}(\theta+\Delta\succeq\theta-\Delta)$ $=\mathbb{P}\left((\theta^{*}x+\epsilon-(\theta+\Delta)x)^{2}\leq(\theta^{*}x+\epsilon-(\theta-\Delta)x)^{2}\right)$
 $=\mathbb{P}(\epsilon\geq(\theta-\theta^{*})|x|)=\frac{1}{2}+\Omega((\theta^{*}-\theta)|x|).$
-If we use $B$ i.i.d.
-samples, then the probability of $\theta+\Delta$ being better is $\frac{1}{2}+\sqrt{B}\Omega((\theta^{*}-\theta)|x|)$.
-The detailed proof is included in Appendix 9.2.
 In instances where evaluation is conducted solely on a random sample, $\kappa$ is equal to 1.
-Considering that humans possess additional side information and can make better comparisons, we anticipate $\kappa$ to be, at most, 1.
+Considering that humans possess or have access to additional side information and thus can make better comparisons, we anticipate that $\kappa$ is, at most, 1.
 Lemma 4.12 (How selection accuracy converges)
@@
 $\tilde{p}(\theta)-1/2\geq c\|\theta-\theta^{*}\|^{\kappa}\,,\forall\theta\in[\theta^{*}-a,\theta^{*}+a],$
-where $c=\min\{\lambda_{\Delta}/(8\gamma\ln(2)),1/(6a)\}$.
-Under Assumption 3.4.1, $\tilde{p}(\theta)$ converges to 1/2 with the speed no faster than the linear convergence.
+where $c=\min\{\lambda_{\Delta}/(8\gamma\ln(2)),1/(6a^{\kappa})\}$.
+Under Assumption 4.4.1, $\tilde{p}(\theta)$ converges to 1/2 at a speed no faster than the linear convergence.
 Therefore, when $\theta$ moves to the $a$-neighborhood of $\theta^{*}$, the distance between $\theta$ and $\theta^{*}$ can be bounded by $((\tilde{p}(\theta)-1/2)/c)^{1/\kappa}$.
-In another word, the closeness between $\tilde{p}(\theta)$ and 1/2 implies the closeness between $\theta$ and $\theta^{*}$.
-Thus, we can define $\tau^{\uparrow}(\delta,\varepsilon)$ as the threshold of the vertical test such that if the vertical move $T^{\uparrow}(\theta)$ has reached $\tau^{\uparrow}(\delta,\varepsilon)$, then $\theta$ is close enough to $\theta^{*}$.
+In other words, the closeness between $\tilde{p}(\theta)$ and 1/2 implies the closeness between $\theta$ and $\theta^{*}$.
+Thus, we can define $\tau^{\uparrow}(\delta,\varepsilon)$ as the threshold of the vertical test, such that if the vertical move $T^{\uparrow}(\theta)$ has reached $\tau^{\uparrow}(\delta,\varepsilon)$, then $\theta$ is close enough to $\theta^{*}$.
 Theorem 4.13 (Vertical moves in Phase II)
-Suppose Assumption 3.4.1 holds. When $T^{\rightarrow}\geq\tau_{1}^{\rightarrow}(a,\delta/2)$ and $T^{\uparrow}(\theta_{T^{\rightarrow}})\geq\tau^{\uparrow}(\delta/2,\varepsilon)$ where
+Suppose that Assumption 4.4.1 holds. When $T^{\rightarrow}\geq\tau^{\rightarrow}(\delta/2,a)$ and $T^{\uparrow}(\theta_{T^{\rightarrow}})\geq\tau^{\uparrow}(\delta/2,\varepsilon)$, where
 $\tau^{\uparrow}(\delta,\varepsilon)=\max\left\{\tau_{0},\frac{8\log(1/\delta)}{c^{2}\varepsilon^{2\kappa}}\right\}=O(\log(1/\delta)\varepsilon^{-2\kappa})\,,$
-and $\tau_{0}=\{\max\ s:\hbar_{s}/s\geq c\varepsilon^{\kappa}/2\}$ where $c$ is defined in Lemma 3.12, then we have
+and $\tau_{0}=\max\{s:\hbar_{s}/s\geq c\varepsilon^{\kappa}/2\}$, where $c$ is defined in Lemma 4.12, then we have
 $\mathbb{P}(\|\theta_{T^{\rightarrow}}-\theta^{*}\|\leq\varepsilon)\geq 1-\delta\,.$
-When Assumption 3.4.1 holds and $\theta$ moves to the neighborhood of $\theta^{*}$, the sample complexity of the vertical move would not exceed $O(\varepsilon^{-2\kappa})$ where $\varepsilon$ is the precision.
-The vertical sample complexity increases with the value of $\kappa$.
+When Assumption 4.4.1 holds and $\theta$ moves to the neighborhood of $\theta^{*}$, the sample complexity of the vertical move would not exceed $O(\varepsilon^{-2\kappa})$, where $\varepsilon$ is the precision.
+The vertical sample complexity increases as the value of $\kappa$ increases.
 The faster the selection accuracy converges to 1/2, the harder the selection problem, and thus the higher the vertical sample complexity.
-In the extreme case where $\kappa=0$, $\tilde{p}(\theta)$ is bounded away from 1/2 so $\tau^{\uparrow}(\delta,\varepsilon)$ is bounded.
+In the extreme case, where $\kappa=0$, $\tilde{p}(\theta)$ is bounded away from 1/2 so $\tau^{\uparrow}(\delta,\varepsilon)$ is bounded.
 This conclusion is consistent with what we have in Theorem 4.11.
 4.5 Sample Complexity
-Combining the sample complexity of horizontal moves and vertical moves, we conclude the total sample complexity of APB.
-Theorem 3.14 (Sample complexity of APB)
+Combining the sample complexity of horizontal moves and vertical moves, we conclude the total sample complexity of MAPB.
+Theorem 4.14 (Sample complexity of MAPB)
 Under Assumption 4.4.1, the expected number of moves is bounded by
@@
 where
-$H(\delta,\varepsilon;a)=\left(\frac{4\theta^{*}}{r_{2}\alpha a}+\frac{4\beta_{1}}{r_{1}}+\frac{2\beta_{2}}{r_{R}}\right)\max_{a\leq\|\theta-\theta^{*}\|\leq\beta_{\Theta}}\phi(\tilde{p}(\theta))+\max\{\tau^{\rightarrow}(\delta,\varepsilon),\tau^{\rightarrow}(\delta/2,a)\}\tau^{\uparrow}(\delta/2,\varepsilon).$
+$H(\delta,\varepsilon;a)=\left(\frac{4\theta^{*}}{r_{2}\alpha a^{\varsigma}}+\frac{4\beta_{1}}{r_{1}}+\frac{2\beta_{2}}{r_{R}}\right)\max_{a\leq\|\theta-\theta^{*}\|\leq\beta_{\Theta}}\phi(\tilde{p}(\theta))+\tau^{\rightarrow}(\delta/2,\varepsilon)\tau^{\uparrow}(\delta/2,\varepsilon).$
 Theorem 4.14 analyzes the (human) sample complexity in a one-dimensional space, which is in the order of $\tilde{O}(\varepsilon^{-2\kappa})$.
 The horizontal move is in the logarithm order, so the dominant part comes from the vertical moves in the neighborhood of $\theta^{*}$.
-If human have the ability to differentiate between any two parameters $\theta_{1}$ and $\theta_{2}$ with probability strictly higher than $1/2+\delta^{\prime}$ as long as $\theta_{1}\neq\theta_{2}$, then the total complexity is in the logarithm order of both confidence $1/\delta$ and precision $1/\varepsilon$.
-In a $d$-dimensional space, if naively refining parameters along each dimension to make the parameter within a $\varepsilon$-ball (in $\ell_{p}$-norm), then the total complexity is $dH(\delta/d,\varepsilon/d^{1/p};a)$, which scales as $\tilde{O}(d^{1+\frac{2\kappa}{p}}/\varepsilon^{2\kappa})$.
-However, in a high-dimensional space, $d$ can be very large while the true parameter may lie in a low-dimensional manifold.
-However, in high-dimensional spaces, $d$ can be significant even though the true parameter may lie within a low-dimensional manifold.
+If humans have the ability to differentiate between any two parameters $\theta_{1}$ and $\theta_{2}$, with the probability strictly higher than $1/2+\delta^{\prime}$ as long as $\theta_{1}\neq\theta_{2}$, then the total complexity is in the logarithm order of both confidence $1/\delta$ and precision $1/\varepsilon$.
+In a $d$-dimensional space, if naively refining parameters along each dimension to make the parameter within an $\varepsilon$-ball (in $\ell_{\varrho}$-norm), then the total complexity is $dH(\delta/d,\varepsilon/d^{1/\varrho};a)$, which scales as $\tilde{O}(d^{1+\frac{2\kappa}{\varrho}}/\varepsilon^{2\kappa})$.
+However, in high-dimensional spaces, $d$ can be significant even though the true parameter may lie within a low-dimensional space.
 In the following section, we explore leveraging a supervised learning oracle to reduce sample complexity.
-4 Alignment: Human-AI Interaction
-In the last section, we discussed the framework of purely using human feedback in a one-dimensional space.
-In practice, a noisy labeled dataset can usually be acquired easily.
-A noisy labeled dataset can be potentially utilized for learning the low-dimensional representation $\varphi(\cdot)$ and thus reducing the complexity of the refining step.
+5 Model Alignment: Human-AI Interaction
+In the previous section, we discussed the framework of using human feedback in a one-dimensional space.
+In practice, a noisy labeled dataset usually can be acquired easily.
+A noisy labeled dataset potentially can be used for learning the low-dimensional representation $\varphi(\cdot)$ and thus for reducing the complexity of the refining step.
 In this section, we propose a two-stage framework where the first stage learns the low-dimensional representation using noisy labels and the second stage refines the model using human feedback.
 5.1 Two-stage Framework
-Suppose the true parameter lies in a $s$-dimensional space where $s\ll d$. Given confidence $\delta$ and precision $\varepsilon$, we introduce a two-stage framework SFT+HC:
+Suppose the true parameter lies in an $s$-dimensional space where $s\ll d$. Given confidence $\delta$ and precision $\varepsilon$, we introduce a two-stage framework SL+LHF:
   1. (I)
     In Stage 1, the algorithm feeds $N_{1}^{HAI}(\delta,\varepsilon)$ noisy labeled data $(\boldsymbol{X},Y)$ to the supervised learning oracle (e.g., Lasso), where $\boldsymbol{X}\in\mathbb{R}^{N_{1}^{HAI}\times d}$ is the design matrix and $Y$ is the noisy response.
-At the end, we learn the $s$-dimensional embedding $\varphi(\cdot)$;
+At the end, we learn the $s$-dimensional embedding $\varphi(\cdot)$.
   2. (II)
-    In Stage 2, the algorithm asks human to refine the estimator along $s$ dimensions. This stage requires $N_{2}^{HAI}(\delta,\varepsilon)$ data points.
-Algorithm 3 SFT+HC (Supervised Fine-Tuning + Human Comparison)
+    In Stage 2, the algorithm asks a human to refine the estimator along $s$ dimensions. This stage requires $N_{2}^{HAI}(\delta,\varepsilon)$ data points.
+Algorithm 3 SL+LHF (Supervised Fine-Tuning+Human Comparison)
 1:Input: Confidence level $\delta$, precision $\varepsilon$; initial distribution $\mu_{0}$; local parameter $a>0$; $k=0$; $\eta$;
-2:Stage 1: Learn the embedding $\varphi(\cdot)$ through SFT oracle trained on data $(\boldsymbol{X},Y)$ with size $N_{1}^{HAI}(\delta,\varepsilon)$;
-3:Stage 2: Execute Algorithm 2 (APB) by querying $N_{2}^{HAI}(\delta,\varepsilon)$ samples.
+2:Stage 1: Learn the embedding $\varphi(\cdot)$ through SL oracle trained on data $(\boldsymbol{X},Y)$ with size $N_{1}^{HAI}(\delta,\varepsilon)$;
+3:Stage 2: Execute Algorithm 2 (MAPB) by querying $N_{2}^{HAI}(\delta,\varepsilon)$ samples.
 The framework design is grounded in the insight that certain representation learning techniques, like autoencoders or deep learning architectures, are inherently robust to noisy labels (Li et al.
@@
 Stage 1 is dedicated to uncovering underlying patterns and structures within the data.
-Given the effectiveness of human expertise in model refinement, Stage 2 endeavors to enhance model alignment through targeted sample querying.
+Given the effectiveness of human expertise in model refinement, Stage 2 serves to enhance model alignment through targeted sample querying.
 In what follows, we specifically discuss sparse linear models (Example 3.1), where Stage 1 applies Lasso [^2] OLS and thresholding may serve the same purpose as Lasso.
@@
 5.2 Illustration on Linear Models
-To demonstrate the sample complexity reduction for the two-stage framework compared to purely trained by supervised learning and provide theoretical justifications, we focus on sparse linear models:
+To demonstrate the sample complexity reduction for the two-stage framework, compared to a framework trained purely by supervised learning, and to provide theoretical justifications, we focus on sparse linear models:
 $Y_{i}=\left\langle\boldsymbol{\vartheta^{*}},X_{i}\right\rangle+\epsilon_{i}=\left\langle\boldsymbol{\theta}^{*},\varphi(X_{i})\right\rangle+\epsilon_{i},$
-where $\varphi(\cdot)$ projects $X_{i}$ to all important features; $S$ is the index set of all important features; the parameter $\boldsymbol{\theta}^{*}$ satisfies that $|\theta^{*}_{i}|\geq\underline{\beta}_{\Theta}>0$ for all $i\in S$.
-In the initial stage, we employ Lasso for the selection of significant features, followed by human comparisons for alignment.
-To assess the sample complexity of our two-stage framework, we will sequentially quantify the size of samples needed for two stages.
-Firstly, we establish the following standard assumptions for the analysis of Lasso.
+where $\varphi(\cdot)$ projects $X_{i}$ to all important features; $S$ is the index set of all important features; and the parameter $\boldsymbol{\theta}^{*}$ satisfies that $|\theta^{*}_{i}|\geq\underline{\beta}_{\Theta}>0$ for all $i\in S$.
+In the initial stage, we use Lasso for the selection of significant features, followed by human comparisons for model alignment.
+To assess the sample complexity of our two-stage framework, we sequentially quantify the size of samples needed for the two stages.
+First, we establish the following standard assumptions for the analysis of Lasso.
 {assumption}
@@
 {assumption}
-For the random design matrix $\boldsymbol{X}\in\mathbb{R}^{n\times d}$ input in the first stage, we assume the following three conditions holds:
+For the random design matrix $\boldsymbol{X}\in\mathbb{R}^{n\times d}$ input in the first stage, we assume that the following three conditions hold:
   1. (I)
-    Mutual incoherence: $\max_{j\in S^{c}}\ \|(\boldsymbol{X}_{S}^{\top}\boldsymbol{X}_{S})^{-1}\boldsymbol{X}_{S}^{\top}x_{j}\|_{1}\leq\alpha_{1}.$
+    Mutual incoherence: $\max_{j\in S^{c}}\ \|(\boldsymbol{X}_{S}^{\top}\boldsymbol{X}_{S})^{-1}\boldsymbol{X}_{S}^{\top}x_{j}\|_{1}\leq\alpha_{1};$
   2. (II)
-    Lower eigenvalue: $\lambda_{\min}\left(V_{S}\right)\geq\lambda_{0}>0,$ where $V_{S}=\frac{1}{n}\boldsymbol{X}_{S}^{\top}\boldsymbol{X}_{S}$.
+    Lower eigenvalue: $\lambda_{\min}\left(V_{S}\right)\geq c_{\min}>0,$ where $V_{S}=\frac{1}{n}\boldsymbol{X}_{S}^{\top}\boldsymbol{X}_{S}$; and
   3. (III)
@@
 Assumption 5.2 is standard for the analysis of Lasso.
-Consider a random design matrix $\boldsymbol{X}\in\mathbb{R}^{n\times d}$ with i.i.d.
-$\mathcal{N}(0,1)$ entries for example, we can easily verify that it satisfies mutual incoherence, lower eigenvalue, and $\ell_{\infty}$-curvature condition.
-Under Assumption 4.2 and 4.2 and based on Corollary 7.22 in Wainwright (2019), we derive the following theorem that bounds the infinite distance between vector $\hat{\theta}_{S}$ and $\theta^{*}_{S}$.
+For example, consider a random design matrix $\boldsymbol{X}\in\mathbb{R}^{n\times d}$ with i.i.d.
+$\mathcal{N}(0,1)$ entries; we can easily verify that it satisfies mutual incoherence, lower eigenvalue, and $\ell_{\infty}$-curvature condition.
+Under Assumptions 5.2 and 5.2 and based on Corollary 7.22 in Wainwright (2019), we derive the following theorem that bounds the infinite distance between vector $\hat{\theta}_{S}$ and $\theta^{*}_{S}$.
 Theorem 5.1
-Under Assumption 4.2 and 4.2, suppose that we solve the Lagrangian Lasso with regularization parameter
+Under Assumptions 5.2 and 5.2, suppose that we solve the Lagrangian Lasso with regularization parameter
 $\lambda_{n}=\frac{2C\sigma}{1-\alpha_{1}}\left\{\sqrt{\frac{2\log(d-s)}{n}}+\zeta\right\}$
 for $\zeta=\frac{1}{4}\min\left\{\frac{\sqrt{c_{\min}}}{\sigma\underline{\beta}_{\Theta}},\frac{\alpha_{2}(1-\alpha_{1})}{2C\sigma\underline{\beta}_{\Theta}}\right\}$.
-Then the optimal solution $\hat{\theta}$ is unique with its support contained within $S$, and satisfies the $\ell_{\infty}$-error bound
+Then, the optimal solution $\hat{\theta}$ is unique, with its support contained within $S$, and it satisfies the $\ell_{\infty}$-error bound
 $\|\hat{\theta}_{S}-\theta^{*}_{S}\|_{\infty}\leq\underbrace{\frac{\sigma}{\sqrt{c_{\min}}}\left\{\sqrt{\frac{2\log s}{n}}+\zeta\right\}+\frac{1}{\alpha_{2}}\lambda_{n}}_{B_{n}}\,,$
-all with probability at least $1-4e^{-\frac{n\zeta^{2}}{2}}$. Moreover, there is no false inclusion: The solution has its support set $\hat{S}$ contained with the true support set $S$.
+all with probability at least $1-4e^{-\frac{n\zeta^{2}}{2}}$. Moreover, there is no false inclusion: The solution has its support set $\hat{S}$ contained within the true support set $S$.
 Remark 5.2
-Not limited to the linear function, Lasso can also be applied to the nonlinear function class (Plan and Vershynin 2016) or nonparametric variable selection (Li et al.
+Lasso is not limited to the linear function but can also be applied to the nonlinear function class (Plan and Vershynin 2016) or to a nonparametric variable selection (Li et al.
 2023).
-Moreover, there is a rich literature of representation learning that can be applied in the first stage (Bengio et al.
+Moreover, there is a rich stream of literature on representation learning that can be applied in the first stage (Bengio et al.
 2013).
 Theorem 5.1 implies that as long as $B_{n}\leq\underline{\beta}_{\Theta}$, the variable selection is consistent.
-Since $\zeta=\frac{1}{4}\min\left\{\frac{\sqrt{c_{\min}}}{\sigma\underline{\beta}_{\Theta}},\frac{\alpha_{2}(1-\alpha_{1})}{2C\sigma\underline{\beta}_{\Theta}}\right\}$, we only need
+Because $\zeta=\frac{1}{4}\min\left\{\frac{\sqrt{c_{\min}}}{\sigma\underline{\beta}_{\Theta}},\frac{\alpha_{2}(1-\alpha_{1})}{2C\sigma\underline{\beta}_{\Theta}}\right\}$, we only need
 $\frac{\sigma}{\sqrt{c_{\min}}}\sqrt{\frac{2\log s}{n}}+\frac{1}{\alpha_{2}}\cdot\frac{2C\sigma}{1-\alpha_{1}}\sqrt{\frac{2\log(d-s)}{n}}\leq\frac{1}{2}\underline{\beta}_{\Theta}.$
-Thus if there are $n$ noisy samples in which
-$n\geq\max\left\{\frac{8\sigma^{2}\log s}{\underline{\beta}_{\Theta}^{2}c_{\min}},\frac{32C^{2}\sigma^{2}\log(d-s)}{(\underline{\beta}_{\Theta}\alpha_{2}(1-\alpha_{1}))^{2}}\right\},$
+Thus, if there are $n$ noisy samples in which
+$n\geq\max\left\{\frac{32\sigma^{2}\log s}{\underline{\beta}_{\Theta}^{2}c_{\min}},\frac{128C^{2}\sigma^{2}\log(d-s)}{(\underline{\beta}_{\Theta}\alpha_{2}(1-\alpha_{1}))^{2}}\right\},$
 we are able to select $s$ variables with probability at least $1-4e^{-\frac{n\zeta^{2}}{2}}$. Define
-$H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})=\max\left\{\frac{8\sigma^{2}\log s}{\underline{\beta}_{\Theta}^{2}c_{\min}},\frac{32C^{2}\sigma^{2}\log(d-s)}{(\underline{\beta}_{\Theta}\alpha_{2}(1-\alpha_{1}))^{2}},\frac{2\log(4/\delta)}{\zeta^{2}}\right\}=O(\sigma^{2}\log d/\underline{\beta}_{\Theta}^{2}).$
-Then, when having more than $H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})$ noisy labeled data at hand, we can have the correct feature selection with probability at least $1-\delta$.
-Moreover, if we want to do one more step of refinement to restrict the uncertainty set to be within $\mathfrak{r}$-distance where $\mathfrak{r}<\underline{\beta}_{\Theta}$, then we need $H_{0}(\delta,\varepsilon;\mathfrak{r})$ noisy labeled data.
+$H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})=\max\left\{\frac{32\sigma^{2}\log s}{\underline{\beta}_{\Theta}^{2}c_{\min}},\frac{128C^{2}\sigma^{2}\log(d-s)}{(\underline{\beta}_{\Theta}\alpha_{2}(1-\alpha_{1}))^{2}},\frac{2\log(4/\delta)}{\zeta^{2}}\right\}=O(\sigma^{2}\log d/\underline{\beta}_{\Theta}^{2}).$
+Then, when we have more than $H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})$ noisy labeled data at hand, we can have the correct feature selection with probability at least $1-\delta$.
+Moreover, if we want to undertake one more step of refinement to restrict the uncertainty set to be within $\mathfrak{r}$-distance, where $\mathfrak{r}<\underline{\beta}_{\Theta}$, then we need $H_{0}(\delta,\varepsilon;\mathfrak{r})$ noisy labeled data.
 Note that $\mathfrak{r}$, which represents the feasible range for refinement, is the parameter that we can choose.
-The trade-off exists.
-A smaller value of $\mathfrak{r}$ demands more noisy labels in the initial stage.
-However, limiting the feasible region to a smaller area could conserve samples required for human comparisons.
+The trade-off exists: A smaller value of $\mathfrak{r}$ demands more noisy labels in the initial stage.
+However, limiting the feasible region to a smaller area would conserve the samples that are required for human comparisons.
 As previously noted, when both options significantly deviate from the true model, distinguishing the superior one becomes challenging.
-This challenge implies a greater need for samples to ascertain the correct answer.
-Consequently, in certain scenarios, opting for $\mathfrak{r}\leq\underline{\beta}_{\Theta}$ to confine the refinement region may be more optimal.
+This challenge implies a need for more samples to ascertain the correct answer.
+Consequently, in certain scenarios, opting for $\mathfrak{r}\leq\underline{\beta}_{\Theta}$ to confine the refinement region may be optimal.
 Specifically, we select $\mathfrak{r}$ to minimize the total sample complexity:
-$\mathfrak{r}^{*}=\min_{0<\mathfrak{r}\leq\underline{\beta}_{\Theta}}H_{0}(\delta,\varepsilon;\mathfrak{r})+H(\delta,\varepsilon;\mathfrak{r}),$ (4.1)
-where $H(\delta,\varepsilon;\mathfrak{r})$ is defined in Theorem 3.14.
-It is important to emphasize that while the first part depends on the noise $\sigma$, the second part does not rely on $\sigma$ but rather on the detection accuracy $\phi(\tilde{p}(\theta))$.
+$\mathfrak{r}^{*}=\min_{0<\mathfrak{r}\leq\underline{\beta}_{\Theta}}H_{0}(\delta,\varepsilon;\mathfrak{r})+sH(\delta/s,\varepsilon/s^{1/\varrho};\mathfrak{r}),$ (5.1)
+where $H(\delta,\varepsilon;\mathfrak{r})$ is defined in Theorem 4.14 and we measure the distance using $l_{\varrho}$-norm.
+Note that although the first part depends on the noise $\sigma$, the second part relies not on $\sigma$ but on the detection accuracy $\phi(\tilde{p}(\theta))$.
+This point is important to emphasize.
 Hence, the optimal $\mathfrak{r}^{*}$ is contingent on the function of detection accuracy.
 For $\theta$ considerably distant from the true parameter $\theta^{*}$, $\tilde{p}(\theta)$ could potentially approach $1/2$, leading to a substantial value for $\phi(\tilde{p}(\theta))$.
-4.3 Complexity of SFT+HC
-The optimal value of $\mathfrak{r}^{*}$ can be obtained numerically by enumerating and comparing the objectives in Equation (4.1).
-To establish an upper bound on the sample complexity of the two-stage framework to analytically show the benefit, we set $\mathfrak{r}=\underline{\beta}_{\Theta}$.
+5.3 Complexity of SL+LHF
+The optimal value of $\mathfrak{r}^{*}$ can be obtained numerically by enumerating and then comparing the objectives in Equation (5.1).
+To establish an upper bound on the sample complexity for the two-stage framework so that we analytically show its benefit, we set $\mathfrak{r}=\underline{\beta}_{\Theta}$.
 In this case, we have
-$H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})=\tilde{O}(\sigma^{2}/\underline{\beta}_{\Theta}^{2})\quad\text{and}\quad H(\delta,\varepsilon;\underline{\beta}_{\Theta})=\tilde{O}\left(s\bar{\phi}(a,\underline{\beta}_{\Theta})+s^{1+\frac{2\kappa}{p}}/\varepsilon^{2\kappa}\right),$
-where $\bar{\phi}(a,D)=\max_{a\leq\|\theta-\theta^{*}\|\leq\underline{\beta}_{\Theta}}\phi(\tilde{p}(\theta))$. Theorem 4.3 is a direct conclusion from Theorem 3.14 and 4.1.
-Theorem 4.3 (Computational complexity of SFT+HC)
-Under Assumption 4.2 and 4.2, by setting $N_{1}^{HAI}(\delta)=H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})$ and then running Algorithm 2 with parameter range $[-\underline{\beta}_{\Theta},\underline{\beta}_{\Theta}]$, it holds that $\|\hat{\theta}-\theta^{*}\|_{p}\leq\varepsilon$ with probability at least $1-2\delta$.
+$H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})=\tilde{O}(\sigma^{2}/\underline{\beta}_{\Theta}^{2})\quad\text{and}\quad sH(\delta/s,\varepsilon/s^{1/\varrho};\underline{\beta}_{\Theta})=\tilde{O}\left(s\bar{\phi}(a,\underline{\beta}_{\Theta})+s^{1+\frac{2\kappa}{\varrho}}/\varepsilon^{2\kappa}\right),$
+where $\bar{\phi}(a,D)=\max_{a\leq\|\theta-\theta^{*}\|\leq\underline{\beta}_{\Theta}}\phi(\tilde{p}(\theta))$. Theorem 5.3 is a direct conclusion from Theorems 4.14 and 5.1.
+Theorem 5.3 (Computational complexity of SL+LHF)
+Under Assumptions 5.2 and 5.2, set $N_{1}^{HAI}(\delta)=H_{0}(\delta,\varepsilon;\underline{\beta}_{\Theta})$ and then run Algorithm 2 with parameter range $[-\underline{\beta}_{\Theta},\underline{\beta}_{\Theta}]$; then, it holds that $\|\hat{\theta}-\theta^{*}\|_{\varrho}\leq\varepsilon$ with probability at least $1-2\delta$.
 Moreover, under Assumption 4.4.1, the expected sample complexity is
-$N_{1}^{HAI}(\delta,\varepsilon)+N_{2}^{HAI}(\delta,\varepsilon)=\tilde{O}(\sigma^{2}/\underline{\beta}_{\Theta}^{2}+s\bar{\phi}(a,\underline{\beta}_{\Theta})+s^{1+\frac{2\kappa}{p}}/\varepsilon^{2\kappa}).$
-Theorem 4.3 demonstrates the sample complexity of the two-stage framework SFT+HC.
-When the distance is measured in two-norm (i.e., $p=2$), the complexity is in the order of $\tilde{O}(\sigma^{2}/\underline{\beta}_{\Theta}^{2}+s\bar{\phi}(a,\underline{\beta}_{\Theta})+s^{1+\kappa}/\varepsilon^{2\kappa})$.
-Instead, if exclusively utilizing Lasso for the supervised learning oracle to refine the estimator based on $n$ noisy labeled data, subject to certain regularity conditions, the estimation error can be bounded as $\|\hat{\theta}_{n}-\theta^{*}\|_{2}\leq c\sigma\sqrt{s\log d/n}$ with high probability for some constant $c>0$.
-Equivalently, to ensure that the two-norm distance of the estimation error is within $\varepsilon$, we need to acquire $\tilde{O}(\sigma^{2}s\log d/\varepsilon^{2})$ data points.
-Comparing the sample complexity of SFT+HC characterized by Theorem 4.3 for $p=2$, the condition for SFT+HC requiring less sample complexity than SFT is characterized in Proposition 4.4.
+$N_{1}^{HAI}(\delta,\varepsilon)+N_{2}^{HAI}(\delta,\varepsilon)=\tilde{O}(\sigma^{2}/\underline{\beta}_{\Theta}^{2}+s\bar{\phi}(a,\underline{\beta}_{\Theta})+s^{1+\frac{2\kappa}{\varrho}}/\varepsilon^{2\kappa}).$
+Theorem 5.3 demonstrates the sample complexity of the two-stage framework SL+LHF.
+When the distance is measured in two-norm (i.e., $\varrho=2$), the complexity is in the order of $\tilde{O}(\sigma^{2}/\underline{\beta}_{\Theta}^{2}+s\bar{\phi}(a,\underline{\beta}_{\Theta})+s^{1+\kappa}/\varepsilon^{2\kappa})$.
+Instead, if Lasso is used exclusively for the supervised learning oracle to refine the estimator based on $n$ noisy labeled data (subject to certain regularity conditions), the estimation error can be bounded as $\|\hat{\theta}_{n}-\theta^{*}\|_{2}\leq c^{\prime}\sigma\sqrt{s\log d/n}$ with high probability for some constant $c^{\prime}>0$.
+Equivalently, we need to acquire $\tilde{O}(\sigma^{2}s\log d/\varepsilon^{2})$ data points to ensure that the two-norm distance of the estimation error is within $\varepsilon$.
+To compare the sample complexity of SL+LHF as indicated by Theorem 5.3 for $\varrho=2$, we characterize the condition of SL+LHF that requires less sample complexity than SL in Proposition 5.4.
 Proposition 5.4
-For sufficiently small $\varepsilon$, SFT+HC requires less sample complexity than pure SFT for solving the $(\varepsilon,\delta)$-alignment problem when (ignoring the logarithm term)
+For sufficiently small $\varepsilon$, SL+LHF requires less sample complexity than pure SL for solving the $(\varepsilon,\delta)$-alignment problem (ignoring the logarithm term) when:
 $\frac{\sigma^{2}}{s^{\kappa}}\gtrapprox\varepsilon^{2-2\kappa}.$ (5.2)
-We call the ratio $\sigma^{2}/s^{\kappa}$ as label noise-to-comparison-accuracy ratio (LNCA ratio) and condition (4.2) as LNCA condition.
+We call the ratio $\sigma^{2}/s^{\kappa}$ the label noise-to-comparison-accuracy ratio (LNCA ratio), and we call Condition (5.2) the LNCA condition.
 This condition involves two important parameters: the observational noise $\sigma$ and the convergence rate of the accuracy $\kappa$.
-When the comparison accuracy is bounded away from 1/2, we have $\kappa=0$ and the condition reduces to $\sigma\gtrapprox\varepsilon$, which holds naturally true.
-In the worst case where $\kappa=1$, the condition becomes $\sigma^{2}\gtrapprox s$.
-When sample observational noise is higher, the benefit of the extra human comparison step becomes more apparent.
+When the comparison accuracy is bounded away from 1/2, we have $\kappa=0$, and the LNCA condition reduces to $\sigma\gtrapprox\varepsilon$, which naturally holds true.
+In the worst case scenario, where $\kappa=1$, the LNCA condition becomes $\sigma^{2}\gtrapprox s$.
+When the observational noise of the sample is higher, the benefit of the human comparison step becomes more apparent.
+As previously mentioned, our framework can be extended to scenarios where the tasks differ between the initial feature-learning stage and the final implementation stage, provided the low-dimensional representation remains consistent.
+This is because the sample complexity of the second refinement stage in our algorithm remains unaffected as long as the representation stays unchanged.
+Meanwhile, Lasso must relearn the varying parameters.
+Therefore, the sufficient condition outlined in Proposition 5.4 continues to hold.
 6 Practical Implementations
-In the preceding sections, we established a framework that assumes humans can compare any two parametric models.
-However, such a comparison could be possibly challenging when humans only have access to the models themselves.
-In practice, it is often more feasible to compare models based on their performance with sampled data, allowing humans to discern which one aligns better with the true underlying model.
-To address this specific challenge, this section delves into the practical aspects of comparing two models using selected samples.
+In the preceding sections, we established a framework that assumes the opportunity for humans to compare any two parametric models.
+However, such a comparison could be challenging when humans have access only to the models themselves.
+In practice, comparing models based on their performance with sampled data is often more useful because this engagement allows humans to discern which model aligns better with the true underlying model.
+To address this specific challenge, in this section, we present the practical aspects of comparing two models using selected samples.
 6.1 Active Learning: How to Select Samples for Comparison
@@
 The key idea is to assess the significance of a sample to determine the value of acquiring its labels.
-Algorithm 2 actively selects a pair of two models to ask for comparison.
-In practical scenarios, it often involves soliciting human judgment to compare the responses predicted by two different models for a given sample data, rather than directly comparing their model parameters.
+Algorithm 2 actively selects a pair of models for comparison.
+In practical scenarios, this choice often involves soliciting human judgment to compare the responses predicted by two different models for the given sample data, rather than directly comparing their model parameters.
 We start with an example.
-The clinician needs to evaluate the Atherosclerotic Cardiovascular Disease (ASCVD) risk of patients $f_{\boldsymbol{\theta}}$, which is a linear function of contextual information, denoted as $x$, such as demographics (age, gender, racial/ethnic group, and education level), baseline conditions (body mass index, history of ASCVD events, and smoking status), clinical variables (glycosylated hemoglobin (A1c) and systolic blood pressure (sbp)), anti-hypertensive agenets, anti-hyperglycemic agents.
-Given sample of patient context $x\in\mathcal{X}$, the clinician evaluates the ASCVD risk according to their expertise.
-The distance between model $f_{\boldsymbol{\theta}}$ and $f_{\boldsymbol{\theta}^{*}}$ evaluated on sample $x$ can be defined as the absolute value of the difference:
+The clinician needs to evaluate the atherosclerotic cardiovascular disease (ASCVD) risk of patients $f_{\boldsymbol{\theta}}$, which is a linear function of contextual information, denoted as $x$.
+Such information comprises demographics (age, gender, racial/ethnic group, and education level); baseline conditions (body mass index, history of ASCVD events, and smoking status); clinical variables (glycosylated hemoglobin (A1c) and systolic blood pressure (sbp)); anti-hypertensive agents; and anti-hyperglycemic agents.
+Given a sample of patients’ contexts $x\in\mathcal{X}$, clinicians evaluate the ASCVD risk according to their expertise.
+The distance between models $f_{\boldsymbol{\theta}}$ and $f_{\boldsymbol{\theta}^{*}}$, evaluated on sample $x$, can be defined as the absolute value of the difference:
 $d_{x}(\boldsymbol{\theta},\boldsymbol{\theta}^{*})=|\varphi(x)^{\top}\boldsymbol{\theta}-\varphi(x)^{\top}\boldsymbol{\theta}^{*}|.$
-Then the utility of $f_{\boldsymbol{\theta}}$ on sample $x$ is $u_{x}(\boldsymbol{\theta})=-d(\boldsymbol{\theta},\boldsymbol{\theta}^{*}|x)=-|\varphi(x)^{\top}\boldsymbol{\theta}-\varphi(x)^{\top}\boldsymbol{\theta}^{*}|$.
-When presented with risk levels predicted by $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$, the clinician would typically opt for the one that is closer to their assessment.
+Then, the utility of $f_{\boldsymbol{\theta}}$ on sample $x$ is $u_{x}(\boldsymbol{\theta})=-d(\boldsymbol{\theta},\boldsymbol{\theta}^{*}|x)=-|\varphi(x)^{\top}\boldsymbol{\theta}-\varphi(x)^{\top}\boldsymbol{\theta}^{*}|$.
+When presented with risk levels predicted by $f_{\boldsymbol{\theta}_{1}}$ and $f_{\boldsymbol{\theta}_{2}}$, clinicians would typically opt for the model that produces results closest to their assessment.
 Let us first consider a simple scenario.
-When $\mathcal{X}$ contains adequate variety of samples such that for any base vector $e_{i}$, there exists sample $x$ such that $\varphi(x)=e_{i}$, then by comparing $f_{\boldsymbol{\theta}_{1}}$ with $f_{\boldsymbol{\theta}_{2}}$ on such sample $x$, we have
+When $\mathcal{X}$ contains an adequate variety of samples, such that for any base vector $e_{i}$, there exists sample $x$ such that $\varphi(x)=e_{i}$, then by comparing $f_{\boldsymbol{\theta}_{1}}$ with $f_{\boldsymbol{\theta}_{2}}$ on this sample $x$, we have
 $d_{x}(\boldsymbol{\theta}_{1},\boldsymbol{\theta}^{*})=|\theta_{1i}-\theta_{i}^{*}|\quad\text{and}\quad d_{x}(\boldsymbol{\theta}_{2},\boldsymbol{\theta}^{*})=|\theta_{2i}-\theta_{i}^{*}|.$
-It implies that the comparison of $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$ when testing on $x$ would be solely on the $i^{th}$ dimension.
-Thus, the alignment process for the parameter in different dimensions can be decoupled, and SFT+HC can be directly applied where the Stage 2 alignment is proceeded along each dimension.
+It implies that the comparison of $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$, when testing on $x$, would focus only on the $i^{th}$ dimension.
+Thus, the alignment process for the parameter in different dimensions can be decoupled, and SL+LHF can be directly applied, where the Stage 2 alignment proceeds along each dimension.
 However, in practice, the dataset $\mathcal{X}$ may be limited, potentially leading to situations where there are no samples $x\in\mathcal{X}$ that satisfy the condition $\varphi(x)=e_{i}$.
-In such cases, it becomes impractical to refine the model parameter along each dimension separately.
+In such cases, refining the model parameter along each dimension separatelybecomes impractical.
 The question then arises: How can we efficiently select samples to refine the model parameter?
-To address this challenge, we propose a procedure for constructing a new basis using the available samples in $\mathcal{X}$.
+To address this challenge, we propose a procedure for constructing a new basis of comparison using the available samples in $\mathcal{X}$.
 This procedure aims to optimize the alignment process by leveraging the existing data to guide the selection of samples for parameter enhancement.
-5.2 New Basis Construction through Gram Schmidt Process
+6.2 New Basis Construction through the Gram Schmidt Process
 To facilitate the independent alignment of each coordinate, we aim to identify a new set of orthogonal basis, which can be constructed using $\Psi:=\{\varphi(x):x\in\mathcal{X}\}$.
-Let $\boldsymbol{\alpha}_{1},\cdots,\boldsymbol{\alpha}_{s}$ represent these new basis, and our objective is to learn $\boldsymbol{\theta}^{*}=\sum_{i=1}^{s}\omega_{i}^{*}\boldsymbol{\alpha}_{i}$, where $\omega_{i}^{*}$ are the corresponding coefficients.
+Let $\boldsymbol{\alpha}_{1},\cdots,\boldsymbol{\alpha}_{s}$ represent these new bases, and our objective is to learn $\boldsymbol{\theta}^{*}=\sum_{i=1}^{s}\omega_{i}^{*}\boldsymbol{\alpha}_{i}$, where $\omega_{i}^{*}$ are the corresponding coefficients.
 We describe the process of constructing the orthogonal basis $\{\boldsymbol{\alpha}_{1},\cdots,\boldsymbol{\alpha}_{s}\}$ from the sample space.
@@
 For simplicity, we define $\boldsymbol{z}_{i}=\varphi(x_{i})$.
-Then we use the Gram Schmidt process to construct a set of orthogonal basis based on $\boldsymbol{z}_{1},\cdots,\boldsymbol{z}_{s}$:
+Then, we use the Gram Schmidt process to construct a set of orthogonal bases, based on $\boldsymbol{z}_{1},\cdots,\boldsymbol{z}_{s}$:
 $\boldsymbol{\alpha}_{k}=\boldsymbol{z}_{k}-\sum_{j=1}^{k-1}\frac{\boldsymbol{z}_{k}^{\top}\boldsymbol{\alpha}_{j}}{\boldsymbol{\alpha}_{j}^{\top}\boldsymbol{\alpha}_{j}}\boldsymbol{\alpha}_{j},\quad\forall 1\leq k\leq s.$
-We sequentially align the coordinates $\omega_{1},\cdots,\omega_{s}$.
-When evaluating on sample $x_{1}$, the true value is represented by $\varphi(x_{1})^{\top}\boldsymbol{\theta}^{*}=\boldsymbol{z}_{1}^{\top}\boldsymbol{\theta}^{*}=\omega_{1}^{*}\boldsymbol{z}_{1}^{\top}\boldsymbol{z}_{1}$.
-Utilizing a probabilistic bisection process in 2, we actively select two answers for each query.
+We sequentially align the coordinates as $\omega_{1},\cdots,\omega_{s}$.
+When evaluating a model on sample $x_{1}$, the true value is represented by $\varphi(x_{1})^{\top}\boldsymbol{\theta}^{*}=\boldsymbol{z}_{1}^{\top}\boldsymbol{\theta}^{*}=\omega_{1}^{*}\boldsymbol{z}_{1}^{\top}\boldsymbol{z}_{1}$.
+Using a probabilistic bisection process in Algorithm 2, we actively select two answers for each query.
 Through this iterative refinement, the final value terminates at $y(x_{1})$, which is guaranteed to be within $\varepsilon$-accuracy of the true parameter $\omega_{1}^{*}$ (with high probability).
@@
 $\varphi(x_{2})^{\top}\boldsymbol{\theta}^{*}=\boldsymbol{z}_{2}^{\top}\boldsymbol{\theta}^{*}=\boldsymbol{z}_{2}^{\top}\boldsymbol{\alpha}_{1}\omega_{1}^{*}+\boldsymbol{\alpha}_{2}^{\top}\boldsymbol{\alpha}_{2}\omega_{2}^{*},$
-where $w_{1}$ is refined already.
-Once more, employing the probabilistic bisection process, the value of $z_{2}^{\top}\boldsymbol{\theta}^{*}$ is iteratively refined to $y(x_{2})$.
+where $w_{1}$ is already refined.
+Once more, using the probabilistic bisection process, the value of $z_{2}^{\top}\boldsymbol{\theta}^{*}$ is iteratively refined to $y(x_{2})$.
 Upon termination of the probabilistic bisection process, we obtain the aligned value of the second coordinate as
 $\hat{\omega}_{2}=\frac{y(x_{2})-\boldsymbol{z}_{2}^{\top}\boldsymbol{\alpha}_{1}\hat{\omega}_{1}}{\boldsymbol{\alpha}_{2}^{\top}\boldsymbol{\alpha}_{2}}.$
-Subsequently, at the $k^{th}$ iteration, we test on the sample $\boldsymbol{z}_{k}=\boldsymbol{\alpha}_{k}+\sum_{j=1}^{k-1}\frac{\boldsymbol{z}_{k}^{\top}\boldsymbol{\alpha}_{j}}{\boldsymbol{\alpha}_{j}^{\top}\boldsymbol{\alpha}_{j}}\boldsymbol{\alpha}_{j}$ where the true value equals
+Subsequently, at the $k^{th}$ iteration, we test on the sample $\boldsymbol{z}_{k}=\boldsymbol{\alpha}_{k}+\sum_{j=1}^{k-1}\frac{\boldsymbol{z}_{k}^{\top}\boldsymbol{\alpha}_{j}}{\boldsymbol{\alpha}_{j}^{\top}\boldsymbol{\alpha}_{j}}\boldsymbol{\alpha}_{j}$, where the true value equals
 $\boldsymbol{z}_{k}^{\top}\boldsymbol{\theta}^{*}=\sum_{j=1}^{k-1}\omega_{j}^{*}\boldsymbol{z}_{k}^{\top}\boldsymbol{\alpha}_{j}+\omega_{k}^{*}\boldsymbol{\alpha}_{k}^{\top}\boldsymbol{\alpha}_{k}.$
-Once the value of $\boldsymbol{z}_{k}^{\top}\boldsymbol{\theta}^{*}$ is refined to $y(x_{k})$ through 2, we are able to obtain the aligned value of the $k^{th}$ coordinate as
+After the value of $\boldsymbol{z}_{k}^{\top}\boldsymbol{\theta}^{*}$ is refined to $y(x_{k})$ through Algorithm 2, we can obtain the aligned value of the $k^{th}$ coordinate as
 $\hat{\omega}_{k}=\frac{y(x_{k})-\sum_{j=1}^{k-1}\hat{\omega}_{j}\boldsymbol{z}_{k}^{\top}\boldsymbol{\alpha}_{j}}{\boldsymbol{\alpha}_{k}^{\top}\boldsymbol{\alpha}_{k}}.$
-Through asking human to evaluate $x_{1},\cdots,x_{s}$ and refining $\omega_{1},\cdots,\omega_{s}$ in sequence, the parameter $\boldsymbol{\theta}:=\sum_{i=1}^{s}\hat{\omega}_{i}\boldsymbol{\alpha}_{i}$ is refined within a certain accuracy level.
+By asking a human to evaluate $x_{1},\cdots,x_{s}$ and by refining $\omega_{1},\cdots,\omega_{s}$ in sequence, we can refine the parameter $\boldsymbol{\theta}:=\sum_{i=1}^{s}\hat{\omega}_{i}\boldsymbol{\alpha}_{i}$ within a certain accuracy level.
 We include the pseudo code in Algorithm 4.
-At the end, it is essential to emphasize that the set of orthogonal basis is not unique, as there may exist multiple sets.
-Consequently, we can conduct multiple rounds of alignment, with the option to switch basis from one round to another.
+Note that the set of orthogonal bases is not unique; multiple sets may exist, and recognizing this possibility is essential.
+Consequently, we can conduct multiple rounds of alignment, with the option to switch the basis from one round to another.
 Algorithm 4 ASS (Active Sample Selection)
@@
 5:for $k=1,\cdots,s$ do
-6:     Query sample $x_{k}$ through 2 and get output $y(x_{k})$;
+6:     Query sample $x_{k}$ through Algorithm 2 and get output $y(x_{k})$
 7:     Set $\hat{\omega}_{k}=\frac{y(x_{k})-\sum_{j=1}^{k-1}\hat{\omega}_{j}\varphi(x_{k})^{\top}\boldsymbol{\alpha}_{j}}{\boldsymbol{\alpha}_{k}^{\top}\boldsymbol{\alpha}_{k}}$
@@
 7 Numerical Experiment
-6.1 Testing Sample Complexity
-In the first experiment, we compare the performance of SFT+HC and pure SFT under the same sample size.
+7.1 Synthetic Data Experiments
+In the first experiment, we compare the performance of SL+LHF and SL using the same sample size to assess how parameter variations affect the performance of the two algorithms.
 Experiment setup.
 Consider a high-dimensional linear regression problem where $d=100$ and $s=10$.
-The first ten coordinates of the true parameter are independently drawn from the uniform distribution and the rest coordinates are zero.
-The observational noise is drawn from Gaussian distribution with mean zero and variance $\sigma=\{1,2,5\}$; we choose $\kappa\in[0,0.95]$.
-The selection precision (along each dimension) is set as $\varepsilon=0.1$ and the human expert level in the choice model is $\gamma=1$.
-We conducted 50 repetitions for each experiment and calculated the average across all results.
-In the first stage the two-stage framework, we use lasso and select the sample size large enough such that the estimator for the non-zero coefficient can be deviated from zero.
-That is, we guarantee that important features are selected with high probability.
-Once the algorithm SFT+HC stops, we use the same sample size to train lasso and compute the estimation error.
-We compare the estimation error of two algorithms under the same sample size.
+The first ten coordinates of the true parameter are independently drawn from the uniform distribution, and the rest of the coordinates are zero.
+The observational noise is drawn from Gaussian distribution with a mean of zero and variance $\sigma\in\{1,2,5\}$; we choose $\kappa\in[0,0.95]$, where $\kappa$ comes from Assumption 4.4.1.
+Specifically, the difference of the utility in the two answers $c_{\Delta}^{+}(\theta)$ and $c_{\Delta}^{-}(\theta)$ is assumed to be $\|\theta-\theta^{*}\|^{\kappa}$.
+The selection precision (along each dimension) is set at $\varepsilon=0.1$, and the human expertise level in the choice model is $\gamma=1$.
+We conducted 30 repetitions for each experiment.
+In the first stage of the two-stage framework, we selected a dataset with a sample size proportional to $\sigma^{2}\log d$ to input into Lasso.
+According to Theorem 5.1, this selection ensures that important features are identified with high probability.
+We chose the regularization parameter using cross-validation.
+The second stage is then executed, requiring human comparison samples, to reach the desired precision threshold $\varepsilon$.
+Once the SL+LHF algorithm completes, the combined sample size for both stages is recorded as the sample size for SL+LHF.
+Then, for a fair comparison, we provided the same sample size to SL (to train a lasso with the same number of observations) and computed the estimation error.
+We then looked at the ratio of the estimation errors of the two approaches, denoted by Err(SL+LHF)/Err(SL).
 7.1.1 Experiment results.
-Figure 3 shows the ratio of the estimation error of two algorithms (Err(SFT+HC)/Err(SFT)) with different values of $\kappa$ when $\sigma=1$.
-For smaller value of $\kappa$, human’s selection accuracy is higher when $\theta$ approaches $\theta^{*}$.
+Figure 3 shows the median of the ratio Err(SL+LHF)/Err(SL) using different values of $\kappa$ when $\sigma$ takes values $1$, $2$, or $5$.
+For all values of $\sigma$, when $\kappa$ is smaller, the human’s selection accuracy is higher as $\theta$ approaches $\theta^{*}$.
 In particular, when $\kappa=0$, the selection accuracy is bounded away from 50%.
-Therefore, the task is harder when $\kappa$ increases.
-It is consistent with the increasing trend of the estimation error when $\kappa$ grows from 0 to 1.
-In addition, a larger value of $\kappa$ leads to a larger sample size, so the estimation error of lasso decreases.
-The intersection point is around $\kappa=0.3$.
-When $\kappa\leq 0.3$, SFT+HC performs better than pure SFT.
-In particular, when $\kappa=0$, the error of two-stage framework is roughly 50% of that in SFT.
-Figure 3: Error ratio of the two-stage framework/pure SFT when $\sigma=1$.
-The results for noise levels with $\sigma=2$ and $\sigma=5$ are illustrated in Figures 4.
-Notably, the trends of the error curves exhibit similarity.
-The threshold, defined as the intersection point of these curves, is approximately 0.55 when $\sigma=2$.
-However, when $\sigma=5$, the two curves do not intersect, suggesting that SFT+HC consistently outperforms pure SFT.
-A comparison across the three figures reveals that the superiority of SFT+HC becomes more pronounced in the presence of higher observational noise.
-In conclusion, SFT+HC is more effective under conditions of increased noise and reduced value of $\kappa$.
-(a) Error ratio when $\sigma=2$.
-(b) Error ratio when $\sigma=5$.
-Figure 4: Prediction errors with different noise levels.
-6.1.2 The variation of expert level $\gamma$.
-We change the value of human expert level $\gamma$ to see how the human comparison accuracy would influence the estimation error.
+Therefore, the task is harder as $\kappa$ increases.
+It is consistent with the increasing trend in the estimation error ratio as $\kappa$ grows from 0 to 1.
+In addition, a larger value of $\kappa$ leads to a larger sample size, which leads to a decrease in the estimation error of lasso.
+When $\sigma=1$, the intersection point of the curve, with the horizontal line of the ratio equal to $1$ (red dashed line), is around $\kappa=0.2$.
+That is, when $\kappa\leq 0.2$, SL+LHF outperforms pure SL in more than half of the cases.
+In particular, when $\kappa=0$, the median of the error ratio is close to 0.
+Figure 3: Error ratio of the two-stage framework to the pure SL with different values of $\sigma$.
+The results for noise levels with $\sigma=2$ and $\sigma=5$ are also illustrated in Figure 3.
+Notably, the trends in the error curves are similar.
+The threshold, defined as the intersection point with the red dashed line, is approximately 0.4 when $\sigma=2$ and approximately $0.65$ when $\sigma=5$.
+Overall, the three curves show that the superiority of SL+LHF becomes more pronounced in the presence of higher observational noise.
+In conclusion, SL+LHF is more effective under increased noise or reduced $\kappa$.
+7.1.2 The variation of expertise level $\gamma$.
+We vary the parameter $\gamma$ to model different levels of human expertise and to analyze how the accuracy of human comparisons affects the estimation error.
+For the experiment setup, we set $\sigma=2$ and $s=10$.
 The smaller value of $\gamma$ indicates higher accuracy in the comparison task.
-In particular, when $\gamma$ approaches 0, human can always select the accurate answer.
-Figure 5 shows the performance of two algorithms with $\gamma=1.1$ and $\gamma=0.8$, respectively.
-Comparing two figures, when $\gamma=1.1$, SFT+HC performs better than pure SFT when $\kappa$ is lower than approximately 0.45; when $\gamma=0.8$, SFT+HC performs better than pure SFT when $\kappa$ is lower than approximately 0.65.
-Therefore, it is consistent with our theory that SFT+HC improves its performance when human has a higher expert level in the selection (i.e., the selection accuracy is higher).
-(a) Error ratio when $\gamma=1.1$.
-(b) Error ratio when $\gamma=0.8$.
-Figure 5: Prediction errors with different expert levels.
+In particular, as $\gamma$ approaches 0, the human can always select the accurate answer.
+Figure 4 shows the performance of the two algorithms with $\gamma$ equal to $0.8$, $1$, and $1.1$.
+Comparing the three curves, when $\gamma=1.1$, SL+LHF performs better than pure SL if $\kappa$ is lower than approximately 0.3; when $\gamma=0.8$, SL+LHF performs better than pure SL if $\kappa$ is lower than approximately 0.5.
+Therefore, we find support for our theory that the performance of SL+LHF improves when the human has a higher expertise level in the selection (i.e., the selection accuracy is higher).
+Figure 4: Error ratio of the two-stage framework to the pure SL with different values of $\gamma$.
 7.1.3 The variation of important feature dimensions $s$.
@@
 For the experiment setup, we set $\gamma=1$ and $\sigma=2$.
-Figure 6 illustrates the algorithmic performance with important feature dimension $s=20$ and $s=50$, respectively.
-The figure reveals that the intersection point shifts to a smaller value as the important feature dimension increases.
-This observation aligns with our theoretical condition (4.2), indicating that as the important feature dimension grows, the advantage of SFT+HC diminishes.
-(a) Error ratio when $s=20$.
-(b) Error ratio when $s=50$.
-Figure 6: Prediction errors with different important feature dimensions.
-6.2 Testing LNCA Ratio
-The LNCA ratio, $\sigma^{2}/s^{\kappa}$, determines the benefits of SFT+HC over pure SFT.
-The LNCA ratio varies between different tasks.
-We are interested in estimating this value for a specific task: to determine the percentage of total area enclosed by a given (red) polygon, as shown in Figure 7, for example.
+Figure 5 illustrates the algorithmic performance when important feature dimension $s$ takes values $10$, $20$, and $50$.
+The figure reveals that the intersection point shifts to a smaller value as the important feature dimension value increases.
+This observation aligns with our theoretical condition (5.2), indicating that as the value of the important feature dimension grows, the advantage of SL+LHF diminishes.
+This result highlights the importance of low-dimensional representations in the success of this method.
+Figure 5: Error ratio of the two-stage framework to the pure SL with different values of $s$.
+7.2 MTurk Experiment on the LNCA Ratio
+In this section, we first outline the setup of our MTurk experiment [^3] https://www.mturk.com/.
+We then describe the estimation process for key parameters, including the accuracy convergence rate $\kappa$, expertise level $\gamma$, and noise level $\sigma$.
+After these parameters were estimated, we conducted simulations to compare the performance of the two-stage framework with that of the pure SL approach.
+We designed a task in which participants are asked to estimate the number of points within a square.
 We divided participants into two distinct groups.
-The first group was presented with the polygon figure and asked to provide their estimated percentage of the total enclosed area.
-The second group was shown the same figure but presented with two options for the area percentage.
-Their task was to select the option they believed was closer to the actual percentage of the enclosed area.
-For instance, in Figure 7(b), the true percentage is 50%.
-Human is provided with two options, 71% and 55%, so the correct choice is 55%.
-This dual-group approach aimed to explore the efficacy of human estimation when directly asked to estimate the area versus when presented with comparative options.
-In the latter case, we varied the discrepancy between the two choices to assess how the ease or difficulty of comparison affected the accuracy of their selections.
+The first group was presented with one of the squares and two options for the estimated number of points.
+The second group was shown the same square and asked to provide their estimated number of points.
+(See Figures 6 and 7.) The task for the second group was to select the option they believed was closer to the actual number.
+For instance, in Figure 6(b), the true number is 32.
+Human participants were provided with two options: 55 and 35; the correct choice is 35.
+With this dual-group approach, we aimed to explore the efficacy of human estimation when participants were asked to estimate the number of points versus to choose from comparative options.
+In the latter case, we varied the discrepancy between the two choices to assess how the ease or difficulty of comparison affected the accuracy of participants’ selections.
 A larger disparity between the choices facilitated easier decision-making, while closer options posed a more challenging comparative task.
-Within Figure 7, three testing samples are depicted.
-Notably, sample (a) presents the most challenging comparison among these three instances due to the minimal disparity observed between them.
-Fifty different tasks are launched in Amazon Mechanical Turk [^4] https://www.mturk.com/, as shown in Figure 8, and we collected 1300 data points in total.
-We use the following ways to estimate $\kappa$ and $\sigma$.
-(a) 21% vs. 5%. True: 10%.
-(b) 71% vs. 55%. True: 50%.
-(c) 29% vs. 13%. True: 26%.
-Figure 7: Experiment design.
-(a) The first group: Select a better answer.
-(b) The second group: Estimate the percentage.
-Figure 8: Comparison vs. Estimation.
+Figure 6 shows three testing samples.
+Notably, sample (c) presents the most challenging comparison because of the minimal disparity ($|u(c^{+})-u(c^{-})|)$ observed between them.
+[^4] The gap between the first choice (45) and the true value (36) is 9 while the gap between the second choice (25) and the true value is 11.
+Thus, the difference between 9 and 11 is only 2.
+To compare the performance of the two-stage framework and the pure supervised learning, we first estimated $\kappa$ and $\sigma$ from the online experiment in Amazon MTurk.
+Then we ran the simulation to compare two methods under different precision levels.
+We launched 50 different tasks, all structured in the same way as the task shown in Figure 7, and collected 500 data points in total.
+We estimated $\kappa$ and $\sigma$ as described in the following.
+(a) 96 vs. 76. True: 64.
+(b) 55 vs. 35. True: 32.
+(c) 45 vs. 25. True: 36.
+Figure 6: Experiment design for Group 1.
+(a) Estimate the number of points.
+Figure 7: Experiment design for Group 2.
 Estimation process for $\kappa$.
-Given figures with true parameter values $\theta^{*}$ representing the percentage of total area enclosed by a polygon, we aim to estimate the accuracy convergence rate $\kappa$ between two options using a human choice model.
-We collect multiple samples for each figure and then use these estimates to derive the relationship between the difference in utilities and the accuracy of human choices.
-More specifically, among two choices, assume the smaller one is $c^{-}$ and the larger one is $c^{+}$, where the middle point is $\theta=(c^{-}+c^{+})/2$.
-According to the choice model, human makes the correct choice with accuracy $p(\theta)=\frac{1}{1+\exp(-|u(c^{-})-u(c^{+})|/\gamma)}$.
+Given figures with true parameter value $\theta^{*}$ (i.e., the number of points in the figure), we estimate the accuracy convergence rate $\kappa$ between two options using a human choice model.
+We collect multiple samples for each figure and use these estimates to derive the relationship between the difference in utility and the accuracy of human choices.
+Specifically, among two choices, assume the smaller one is $c^{-}$, the larger one is $c^{+}$, and the middle point is $\theta=(c^{-}+c^{+})/2$.
+According to the choice model, the human makes the correct choice with accuracy $p(\theta)=\frac{1}{1+\exp(-|u(c^{-})-u(c^{+})|/\gamma)}$.
 Note that we try to estimate $\kappa$ from Assumption 4.4.1 that
@@
 That is, the accuracy is $p(\theta)=\frac{1}{1+\exp(-\lambda_{\Delta}/\gamma\cdot|\theta-\theta^{*}|^{\kappa})}$.
-Let $\tilde{\lambda}=-\lambda_{\Delta}/\gamma$.
-For each data point with tested point $\theta_{i}$ and true parameter $\theta_{i}^{*}$, we define $y_{i}=1$ (the correct choice) with probability $p_{i}(\theta)$ and $y_{i}=0$ (the wrong choice) with probability $1-p_{i}(\theta)$.
+Let $\tilde{\lambda}=\lambda_{\Delta}/\gamma$.
+For each data point with tested point $\theta_{i}$ and true parameter $\theta_{i}^{*}$, we define $y_{i}=1$ (the correct choice) with probability $p_{i}(\theta)$, and we define $y_{i}=0$ (the wrong choice) with probability $1-p_{i}(\theta)$.
 By maximizing the log-likelihood, we solve the optimization problem
-$\max_{\tilde{\lambda},\kappa}\quad\sum_{i=1}^{K}y_{i}\log p_{i}+(1-y_{i})\log(1-p_{i})$
+$\max_{\tilde{\lambda},\kappa}\quad\sum_{i=1}^{K}y_{i}\log p_{i}+(1-y_{i})\log(1-p_{i}),$
 which is equivalent to
-$\max\quad\sum_{i=1}^{K}-y_{i}\log(1+\exp(-\tilde{\lambda}|\theta_{i}-\theta^{*}|^{\kappa}))-(1-y_{i})\log(1+\exp(\tilde{\lambda}|\theta_{i}-\theta_{i}^{*}|^{\kappa}))\,.$
-From the data collected from Amazon Mechanical Turk, the estimated $\kappa$ is 0.095.
+$\max_{\tilde{\lambda},\kappa}\quad\sum_{i=1}^{K}-y_{i}\log(1+\exp(-\tilde{\lambda}|\theta_{i}-\theta^{*}|^{\kappa}))-(1-y_{i})\log(1+\exp(\tilde{\lambda}|\theta_{i}-\theta_{i}^{*}|^{\kappa}))\,.$
+We solve this optimization problem simply by enumerating different values of $\tilde{\lambda}$ and $\kappa$.
 Estimation process for $\sigma$.
 The other group is asked to directly estimate the percentage of total area.
-For the figure with true parameter values $\theta_{i}^{*}$, human provides responses by $y_{i}=\theta_{i}^{*}+\epsilon_{i}$.
-We estimate the variance of the noise by $\hat{\sigma}^{2}=\sum_{i=1}^{N}\frac{(y_{i}-\theta_{i}^{*})^{2}}{N-1}=29.04$.
-In the context of a one-dimensional problem, as evidenced by numerical experiments detailed in the preceding section, SFT+FTPB demonstrates a substantial performance advantage over SFT when $\kappa\approx 0.1$ and $\sigma\approx 30$.
-7 Concluding Remarks
+For the figure with true parameter values $\theta_{i}^{*}$, the human provides responses by $y_{i}=\theta_{i}^{*}+\epsilon_{i}$.
+We estimate the variance of the noise by $\hat{\sigma}^{2}=\sum_{i=1}^{N}\frac{(y_{i}-\theta_{i}^{*})^{2}}{N-1}$.
+Empirical result.
+We applied a bootstrap resampling method to the data collected from Amazon MTurk, generating 500 resamples to estimate key parameters, including $\kappa$, $\tilde{\lambda}$, and $\sigma$.
+The resulting mean estimates are $\kappa=0.328$, $\tilde{\lambda}=0.132$, and $\sigma^{2}=29.665^{2}$.
+The variances are 0.061, 0.006, and 4.600 for $\kappa$, $\tilde{\lambda}$, and $\sigma$, respectively.
+Figure 8 presents the median empirical error of the two-stage framework across varying precision levels, based on 100 repetitions for each value of $\varepsilon$.
+The results demonstrate that, for any given precision level, the empirical error remains within the targeted threshold.
+In addition, an upward trend in the empirical error is observed as the precision level increases.
+The green plot illustrates the sample size required to achieve a specified precision, showing a rapid decline in sample size as precision increases.
+For a clearer comparison with the pure SL, Figure 9 displays the error ratio between the two-stage framework and the pure SL.
+When $\varepsilon=0.3\%$, the median of the estimation error of the two-stage framework is approximately 16% of the pure SL.
+Increasing the precision to $\varepsilon=4.3\%$, the median of the estimation error of the two-stage framework is approximately 45% of the pure SL.
+This trend of a decreasing error ratio with tighter precision aligns with Proposition 5.4, highlighting the increasing advantage of the two-stage framework as $\varepsilon$ approaches 0.
+Figure 8: Empirical error of the two-stage framework for different values of $\varepsilon$.
+Figure 9: Error ratio of the two-stage framework to the pure SL for different values of $\varepsilon$.
+In conclusion, we acknowledge limitations in our numerical experiments.
+Ideally, our two-stage framework could be used to acquire responses sequentially from freelancers, first asking them which number is closer and and then asking them for an estimate of the number of points.
+Then the two results could be compared.
+However, implementing this comparison would be challenging because of the adaptive nature of our acquisition process, where each subsequent question depends on previously collected answers.
+Moreover, acquiring responses from the Amazon MTurk platform is costly.
+Given the complexity and expense of this procedure, we opted to use simulations to evaluate the performance of our framework.
+8 Discussions and Concluding Remarks
 This work is motivated by the ever-growing observations of human-AI interaction in LLM.
-We presented a theoretical two-stage framework for explaining the significance of using human comparisons for the model improvement.
-While the noisy-labeled data is fed into the SFT procedure to learn the low-dimensional representation in the first stage, we sequentially inquire human evaluators to make pair-wise comparisons in the second stage.
-To address the challenge of efficient acquisition of valuable information from these human comparisons, we introduced a probabilistic bisection method, which factors in the uncertainty arising from the accuracy of human comparisons.
-The newly introduced concept, label-noise-to-comparison-accuracy ratio, quantifies the relative scale between label noise and human comparison accuracy.
-We demonstrate the significant advantages of our proposed two-stage framework over pure supervised learning approach under certain conditions.
-Our work represents an early attempt to organically integrate machine learning with human knowledge.
-We consider this paper as a prompt for an open discussion.
+We presented a theoretical two-stage framework for explaining the significance of using human comparisons for model improvement.
+In the first stage, the noisy-labeled data is fed into the SL procedure to learn the low-dimensional representation; we then ask human evaluators to make pair-wise comparisons in the second stage.
+To address the challenge of efficient acquisition of valuable information from these human comparisons, we introduced a probabilistic bisection method, which factors in the uncertainty that arises from the accuracy of human comparisons.
+The newly introduced concept, the LNCA ratio, quantifies the relative scale between label noise and human comparison accuracy.
+We demonstrate the significant advantages of our proposed two-stage framework over a pure supervised learning approach under certain conditions.
+8.1 Implications for Supervised Fine-Tuning
+Supervised fine-tuning (SFT) is the process of refining a pre-trained model on a labeled dataset with specific input-output pairs tailored to a target task.
+Our proposed framework can fit into SFT in the following way.
+First, the pre-trained model produces a representation, denoted as $\hat{\varphi}(\cdot)$, along with initial parameters $\hat{\boldsymbol{\theta}}_{0}$.
+For a new task, if $\hat{\varphi}(\cdot)$ remains the same, we proceed directly to the second phase, where human feedback is solicited for model alignment through comparison data.
+However, if the representation $\hat{\varphi}(\cdot)$ also requires adaptation, we first update the embeddings by training on a combination of newly collected and existing noisy-labeled data, using techniques such as transfer learning (Pan 2020).
+In this adapted embedding space, human comparisons are then actively acquired in the second stage to enhance model alignment.
+8.2 Future Directions
+Our work represents an early attempt to organically integrate two sources of data: noisy-labeled data and human comparison data.
+We position this paper as a prompt for an open discussion.
 Some potential extensions to our work are worth investigating.
-Firstly, we extended the probabilistic bisection algorithm from one dimension to multiple dimensions by refining along each dimension.
+First, we extended the probabilistic bisection algorithm from one dimension to multiple dimensions by refining along each dimension.
 This expansion raises the question of whether a more efficient bisection algorithm exists in multi-dimensional spaces.
 As discussed in Frazier et al.
-(2019), the development of a PBA tailored for multi-dimensional problems remains an open problem.
-Secondly, although our paper focuses on a prediction problem, the framework has the potential to be extended to the action-based learning, where the goal is to select optimal actions for the decision-making problem.
-Thirdly, exploring the optimal sample selection to maximize the learning algorithm’s efficiency is a compelling area of inquiry.
-While we have outlined guidelines for practical sample selection, the quest for the most efficient proposed procedure remains open.
+(2019), the development of a PBA tailored to multi-dimensional problems remains an open problem.
+Second, although our paper focuses on a prediction problem, the framework has the potential to be extended to action-based learning, where the goal is to select optimal actions for the decision-making problem.
+Third, exploring the optimal sample selection to maximize the learning algorithm’s efficiency is a compelling area of inquiry.
+Although we have offered guidelines for practical sample selection, the quest for the most efficient proposed procedure remains open.